Israel Journal of Mathematics

Cumulative Index: Vols. 1–100

Aaronson, J., An ergodic theorem with large normalising constants, 38 (1981), 182–188 Aaronson, J., Category theorems for some ergodic multiplier properties, 51 (1985),

151–162

Aaronson, J., On the ergodic theory of non-integrable functions and infinite measure

spaces, 27 (1977), 163–173

Aaronson, J., On the pointwise ergodic behaviour of transformations preserving infinite

measures, 32 (1979), 67–82

Aaronson, J., Rational ergodicity and a metric invariant for Markov shifts, 27 (1977),

93–123

Aaronson, J., Rational ergodicity, bounded rational ergodicity and some continuous

measures on the circle, 33 (1979), 181–197

Aaronson, J., The eigenvalues of non-singular transformations, 45 (1983), 297–312 Aaronson, J., Glasner, E. and Misiurewicz, M., Rigid sets in the plane, 68 (1989),

307–326

Aaronson, J. and Keane, M., Isomorphism of random walks, 87 (1994), 37–63 Aaronson, J., Liggett, T. and Picco, P., Equivalence of renewal sequences and

isomorphism of random walks, 87 (1994), 65–76

Aaronson, J., Lin, M. and Weiss, B., Correction to: “Mixing properties of Markov operators and ergodic transformations, and ergodicity of Cartesian products” [Israel J. Math. 33 (1979), 198–224], 37 (1980), 364–364

Aaronson, J., Lin, M. and Weiss, B., Mixing properties of Markov operators and ergodic

transformations, and ergodicity of Cartesian products, 33 (1979), 198–224

Aaronson, J. and Weiss, B., Generic distributional limits for measure preserving

transformations, 47 (1984), 251–259

Abboud, E., Q-admissibility of certain nonsolvable groups, 59 (1987), 272–282 Abels, H., Finiteness properties of certain arithmetic groups in the function field case, 76

(1991), 113–128

Abels, H., Margulis, G. A. and Soifer, G. A., Semigroups containing proximal linear

maps, 91 (1995), 1–30

Abhyankar, S. S., Nice equations for nice groups, 88 (1994), 1–23 Abraham, U. and Shelah, S., Isomorphism types of Aronszajn trees, 50 (1985), 75–113 Abraham, U. and Shore, R. A., Initial segments of the degrees of size ℵ_{1}, 53 (1986),

1–51

Abramenko, P., Finiteness properties of Chevalley groups over 𝔽_{q}[t], 87 (1994), 203–223 Abramovich, S., The increase of sums and products dependent on (y_{1}, …, y_{n}) by

rearrangement of this set, 5 (1967), 177–181

Abrams, G., Non-induced isomorphisms of matrix rings, 99 (1997), 343–347 Acosta, M. D., Every real Banach space can be renormed to satisfy the denseness of

numerical radius attaining operators, 81 (1993), 273–280

Acosta, M. D., Aguirre, F. J. and Payá, R., There is no bilinear Bishop–Phelps

theorem, 93 (1996), 221–227

Acquistapace, P. and Terreni, B., Existence and sharp regularity results for linear

parabolic non-autonomous integro-differential equations, 53 (1986), 257–303

Aczel, P., Representability in some systems of second order arithmetic, 8 (1970), 309–328 Adams, J., Character of the oscillator representation, 98 (1997), 229–252 Adams, S., Indecomposability of treed equivalence relations, 64 (1988), 362–380 Adams, S., Representation varieties of arithmetic groups and polynomial periodicity of

Betti numbers, 88 (1994), 73–124

Adams, S., Very weak Bernoulli for amenable groups, 78 (1992), 145–176 Adams, S. and Lyons, R., Amenability, Kazhdan’s property and percolation for trees,

groups and equivalence relations, 75 (1991), 341–370

Adams, S. see Sarnak, P. Adamski, W., On the existence of invariant probability measures, 65 (1989), 79–95 Adelman, O., Brownian motion never increases: a new proof to a result of Dvoretzky,

Erdős and Kakutani, 50 (1985), 189–192

Adelman, O. and Dvoretzky, A., Plane Brownian motion has strictly n-multiple points,

52 (1985), 361–364

Adin, R. M. and Frumkin, A., The conjugacy character of S_{n} tends to be regular, 59

(1987), 234–240

Adler, R. L., Goodwyn, L. W. and Weiss, B., Equivalence of topological Markov

shifts, 27 (1977), 49–63

Adler, R. L. and Shields, P. C., Skew products of Bernoulli shifts with rotations, 12

(1972), 215–222

Adler, R. L. and Shields, P. C., Skew products of Bernoulli shifts with rotations. II,

19 (1974), 228–236

Adler, R. L. and Weiss, B., The ergodic infinite measure preserving transformation of

Boole, 16 (1973), 263–278

Agmon, S. and Kannai, Y., On the asymptotic behavoir of spectral functions and resolvent

kernels of elliptic operators, 5 (1967), 1–30

Agnarsson, G., Amitsur, S. A. and Robson, J. C., Recognition of matrix rings II, 96

(1996), 1–13

Agranovsky, M. L. and Semenov, A. M., Deformations of balls in Schiffer’s conjecture

for Riemannian symmetric spaces, 95 (1996), 43–59

Aguadé, J., A note on realizing polynomial algebras, 38 (1981), 95–99 Aguadé, J., Computing Lannes T functor, 65 (1989), 303–310 Aguadé, J., Constructing modular classifying spaces, 66 (1989), 23–40 Aguirre, F. J. see Acosta, M. D. Aharoni, I., Every separable metric space is Lipschitz equivalent to a subset of c_{0}^{+}, 19

(1974), 284–291

Aharoni, I., Uniform embeddings of Banach spaces, 27 (1977), 174–179 Aharoni, I. and Lindenstrauss, J., An extension of a result of Ribe, 52 (1985), 59–64 Aharoni, I., Maurey, B. and Mityagin, B. S., Uniform embeddings of metric spaces

and of Banach spaces into Hilbert spaces, 52 (1985), 251–265

Aharoni, R. and Loebl, M., Strongly perfect infinite graphs, 90 (1995), 81–91 Aharonov, D., A remark on an areally mean p-valent function, 6 (1968), 119–120 Aharonov, D., On the Bieberbach conjecture for functions with a small second coefficient,

15 (1973), 137–139

Aharonov, D., Proof of the Bieberbach conjecture for a certain class of univalent functions,

8 (1970), 103–104

Aharonov, D., Schiffer, M. M. and Zalcman, L., Potato kugel, 40 (1981), 331–339 Aiken, J. G., A problem of Dyer, Porcelli, and Rosenfeld, 24 (1976), 191–198 Aizicovici, S., On a semilinear Volterra integrodifferential equation, 36 (1980), 273–284 Akcoglu, M. A. and King, J. L., An example of pointwise non-convergence of iterated

conditional expectation operators, 94 (1996), 179–188

Akcoglu, M. A. and Krengel, U., Two examples of local ergodic divergence, 33 (1979),

225–230

Akcoglu, M. A. and Sucheston, L., An ergodic theorem on Banach lattices, 51 (1985), 208–222 Albertson, M. O. and Berman, D. M., Every planar graph has an acyclic 7-coloring,

28 (1977), 169–174

Alcalde Cuesta, F. and Hector, G., Intégration symplectique des variétés de Poisson

régulières, 90 (1995), 125–165

Alev, J., Polynôme de Hilbert–Samuel dans les algèbres enveloppantes et les algèbres de

groupe, 37 (1980), 231–240

Alfandary, G., On graphs related to conjugacy classes of groups, 86 (1994), 211–220 Alfsen, E. M., M-structure and intersection properties of balls in Banach spaces, 13

(1972), 235–245

Alikakos, N. D. and Hess, P., On stabilization of discrete monotone dynamical systems,

59 (1987), 185–194

Alikakos, N. D. and Rostamian, R., On the uniformization of the solutions of the porous

medium equation in R^{N}, 47 (1984), 270–290

Aljadeff, E., On the surjectivity of some trace maps, 86 (1994), 221–232 Aljadeff, E. and Robinson, D. J. S., Solution of the twisting problem for skew group

algebras, 91 (1995), 409–417

Aljadeff, E. and Rosset, S., Growth and uniqueness of rank, 64 (1988), 251–256 Alltop, W. O., On the Frattini normal embeddability of products of p-groups, 23 (1976),

31–38

Alon, N., Independent sets in regular graphs and sum-free subsets of finite groups, 73

(1991), 247–256

Alon, N., On the number of certain subgraphs contained in graphs with a given number

of edges, 53 (1986), 97–120

Alon, N., On the number of subgraphs of prescribed type of graphs with a given number

of edges, 38 (1981), 116–130

Alon, N., The linear arboricity of graphs, 62 (1988), 311–325 Alon, N., Kim, J-H. and Spencer, J., Nearly perfect matchings in regular simple

hypergraphs, 100 (1997), 171–187

Alon, N. and Milman, V. D., Embedding of l_{∞}^{k} in finite dimensional Banach spaces, 45

(1983), 265–280

Alon, N., Mohar, B. and Sanders, D. P., On acyclic colorings of graphs on surfaces,

94 (1996), 273–283

Alonso, A. I. and Obaya, R., Linear invariant measures for recurrent linear systems, 92

(1995), 185–205

Alspach, B., Heinrich, K. and Rosenfeld, M., Edge partitions of the complete

symmetric directed graph and related designs, 40 (1981), 118–128

Alspach, D. E., A quotient of C(ω^{ω}) which is not isomorphic to a subspace of C(α),

α < ω_{1}, 35 (1980), 49–60

Alspach, D. E., On the complemented subspaces of X_{p}, 74 (1991), 33–45 Alspach, D. E., Quotients of C[0, 1] with separable dual, 29 (1978), 361–384 Alspach, D. E. and Benyamini, Y., A geometrical property of C(K) spaces, 64 (1988),

179–194

Alspach, D. E. and Benyamini, Y., C(K) quotients of separable L_{∞} spaces, 32 (1979),

145–160

Alspach, D. E. and Benyamini, Y., Primariness of spaces of continuous functions on

ordinals, 27 (1977), 64–92

Alspach, D. E. and Carothers, N. L., Constructing unconditional finite dimensional

decompositions, 70 (1990), 236–256

Altshuler, A., A remark on the polytopality of an interesting 3-sphere, 48 (1984), 159–160 Altshuler, A., Lattice characterization of convex 3-polytopes and of polygonizations of

2-manifolds, 8 (1970), 57–64

Altshuler, A., Bokowski, J. and Schuchert, P., Spatial polyhedra without diagonals, 86 (1994), 373–396 Altshuler, A. and Brehm, U., The topological structure of 3-pseudomanifolds, 39 (1981),

63–73

Altshuler, A. and Perles, M. A., Quotient polytopes of cyclic polytopes. Part I:

Structure and characterization, 36 (1980), 97–125

Altshuler, A. and Perles, M. A., Quotient polytopes of cyclic polytopes. Part II:

Stability of the f-vector and of the k-skeleton, 41 (1982), 275–297

Altshuler, A. and Shemer, I., Construction theorems for polytopes, 47 (1984), 99–110 Altshuler, A. see Brehm, U. Altshuler, Z., Characterization of c_{0} and l_{p} among Banach spaces with symmetric basis,

24 (1976), 39–44

Altshuler, Z., Uniform convexity in Lorentz sequence spaces, 20 (1975), 260–274 Altshuler, Z., Casazza, P. G. and Lin, B. L., On symmetric basic sequences in Lorentz

sequence spaces, 15 (1973), 140–155

Amann, H., Dual semigroups and second order linear elliptic boundary value problems, 45

(1983), 225–254

Amberg, B. and Fransman, A., Products of groups and group classes, 87 (1994), 9–18 Amir, D., On isomorphisms of continuous function spaces, 3 (1965), 205–210 Amir, D., On projections and simultaneous extensions, 2 (1964), 245–248 Amir, D. and Arbel, B., On injections and surjections of continuous function spaces, 15

(1973), 301–310

Amir, D. and Milman, V. D., A quantitative finite-dimensional Krivine theorem, 50

(1985), 1–12

Amir, D. and Milman, V. D., Unconditional and symmetric sets in n-dimensional normed

spaces, 37 (1980), 3–20

Amir, D. and Ziegler, Z., On generalized absolutely monotone functions, 7 (1969),

137–146

Amit, R. and Shelah, S., The complete finitely axiomatized theories of order are dense,

23 (1976), 200–208

Amitsur, S. A., A note on PI-rings, 10 (1971), 210–211 Amitsur, S. A., Complexes of rings, 2 (1964), 143–154 Amitsur, S. A., Identities and linear dependence, 22 (1975), 127–137 Amitsur, S. A., Identities in rings with involutions, 7 (1969), 63–68 Amitsur, S. A., Jacob Levitzki, 1904–1956, 19 (1974), 1–3 Amitsur, S. A., On central division algebras, 12 (1972), 408–420 Amitsur, S. A., Polynomial identities, 19 (1974), 183–199 Amitsur, S. A., Rings with involution, 6 (1968), 99–106 Amitsur, S. A., The generic division rings, 17 (1974), 241–247 Amitsur, S. A., The sequence of codimensions of PI-algebras, 47 (1984), 1–22 Amitsur, S. A. and Baum, D., Finite-dimensional subalgebras of division rings, 96 (1996),

15–22

Amitsur, S. A. and Rowen, L. H., Elements of reduced trace 0, 87 (1994), 161–179 Amitsur, S. A., Rowen, L. H. and Tignol, J-P., Division algebras of degree 4 and 8

with involution, 33 (1979), 133–148

Amitsur, S. A. and Small, L. W., GK-dimensions of corners and ideals, 69 (1990),

152–160

Amitsur, S. A. and Small, L. W., Polynomials over division rings, 31 (1978), 353–358 Amitsur, S. A. and Small, L. W., Algebras over infinite fields, revisited, 96 (1996),

23–25

Amitsur, S. A. see Agnarsson, G. Amitsur, S. A. see Tignol, J-P. Amour, L., Almost periodicity for the nonlinear cubic Schrödinger equation, 92 (1995),

333–348 Andersen, N. J. M. and Christensen, J. P. R., Some results on Borel structures with

applications to subseries convergence in abelian topological groups, 15 (1973), 414–420

Anderson, G. D., Vamanamurthy, M. K. and Vuorinen, M., Distortion functions for

plane quasiconformal mappings, 62 (1988), 1–16

Anderson, G. D., Vamanamurthy, M. K. and Vuorinen, M., Functional inequalities

for the Teichmüller ring capacity, 88 (1994), 367–380

Anderson, J. H. and Stampfli, J. G., Commutators and compressions, 10 (1971),

433–441

Anderson, R. M., A non-standard representation for Brownian motion and Itô integration,

25 (1976), 15–46

Andrew, A. D., James’ quasi-reflexive space is not isomorphic to any subspace of its dual,

38 (1981), 276–282

Andrew, A. D., On subsequences of the Haar system in C(Δ), 31 (1978), 85–90 Andrew, A. D. and Demko, S., Banded perturbations of the unit vector basis in some

sequence spaces, 29 (1978), 389–392

Aoki, N., A simple proof of the Bernoullicity of ergodic automorphisms on compact abelian

groups, 38 (1981), 189–198

Apt, K. R., Inductive definitions, models of comprehension and invariant definability, 29

(1978), 221–238

Apter, A. W., On a problem inspired by determinacy, 61 (1988), 256–270 Apter, A. W., On the class of measurable cardinals without the axiom of choice, 79

(1992), 367–379

Apter, A. W., On the least strongly compact cardinal, 35 (1980), 225–233 Apter, A. W., Some results on consecutive large cardinals II: Applications of Radin

forcing, 52 (1985), 273–292

Ara, P. and del Rio, A., A question of Passman on the symmetric ring of quotients, 68

(1989), 348–352

Arad (Ardinast), Z., Products of Fitting classes of finite groups, 12 (1972), 61–69 Arad, Z., Groups with S_{2′}-subgroups and S_{3′}-subgroups, 17 (1974), 76–83 Arazy, J., The isometries of C_{p}, 22 (1975), 247–256 Arazy, J. and Friedman, Y., The isometries of C_{p}^{n,m} into C_{p}, 26 (1977), 151–165 Arbel, B. see Amir, D. Arendt, W., Vector-valued Laplace transforms and Cauchy problems, 59 (1987), 327–352 Argyros, S. A., Weak compactness in L^{1}(λ) and injective Banach spaces, 37 (1980),

21–33

Argyros, S. A., Negrepontis, S. and Zachariades, T., Weakly stable Banach spaces,

57 (1987), 68–88

Argyros, S. A. and Benyamini, Y., Universal WCG Banach spaces and universal Eberlein

compacts, 58 (1987), 305–320

Argyros, S. A. and Deliyanni, I., Non-dentable sets in Banach spaces with separable

dual, 81 (1993), 53–64

Arias, A., l_{p}^{n} superspaces of spans of independent random variables, 63 (1988), 139–148 Arias de Reyna, J., Correction to: “Normed barely Baire spaces” [Israel J. Math. 42

(1982), no. 1-2, 33–36], 50 (1985), 264–264

Arias de Reyna, J., Normed barely Baire spaces, 42 (1982), 33–36 Arias-de-Reyna, J. and Rodríguez-Piazza, L., Finite metric spaces needing high

dimension for Lipschitz embeddings in Banach spaces, 79 (1992), 103–111

Armendariz, E. P. and Park, J. K., The double centralizer theorem for division algebras,

45 (1983), 63–68

Arnold, L., Convergence in probability of random power series and a related problem in

linear topological spaces, 5 (1967), 127–134

Arnoux, P., Ornstein, D. S. and Weiss, B., Cutting and stacking, interval exchanges and geometric models, 50 (1985), 160–168 Artstein, Z., Distributions of random sets and random selections, 46 (1983), 313–324 Artzy, R., Loop laws and collineations of planes, 10 (1971), 106–113 Artzy, R., Non-euclidean incidence planes, 4 (1966), 43–53 Asada, M. and H. Nakamura, On graded quotient modules of mapping class groups of

surfaces, 90 (1995), 93–113

Aschbacher, M., Simple connectivity of p-group complexes, 82 (1993), 1–43 Aschbacher, M. and Segev, Y., Locally connected simplicial maps, 77 (1992), 285–303 Asensio Mayor, J. and Martinez Hernandez, J., Flat envelopes in commutative rings,

62 (1988), 123–128

Ash, A., Farrell cohomology of GL(n, Z), 67 (1989), 327–336 Ash, J. M., Rieders, E. and Kaufman, R. P., The Cantor–Lebesgue property, 84

(1993), 179–191

Asplund, E., A k-extreme point is the limit of k-exposed points, 1 (1963), 161–162 Asplund, E., Averaged norms, 5 (1967), 227–233 Asplund, E., Farthest points in reflexive locally uniformly rotund Banach spaces, 4 (1966),

213–216

Asplund, E., Sets with unique farthest points, 5 (1967), 201–209 Assani, I., Lesigne, E. and Rudolph, D. J., Wiener–Wintner return-times ergodic

theorem, 92 (1995), 375–395

Assouad, P., Remarques sur un article de Israel Aharoni sur les prolongements Lipschitziens

dans c_{0}, 31 (1978), 97–100

Attouch, H. and Damlamian, A., On multivalued evolution equations in Hilbert spaces,

12 (1972), 373–390

Atzmon, A., Non-singly-generated closed ideals in group algebras, 7 (1969), 303–310 Atzmon, A., Spectral synthesis in regular Banach algebras, 8 (1970), 197–212 Aumann, R. J. and Hart, S., Bi-convexity and bi-martingales, 54 (1986), 159–180 Aumann, R. J. and Kurz, M., Power and taxes in a multi-commodity economy, 27 (1977),

185–234

Auslander, J. and Berg, K., A condition for zero entropy, 69 (1990), 59–64 Auslander, J. and Katznelson, Y., Continuous maps of the circle without periodic

points, 32 (1979), 375–381

Avraham, U., Free sets for nowhere-dense set mappings, 39 (1981), 167–176 Avraham, U., Devlin, K. J. and Shelah, S., The consistency with CH of some

consequences of Martin’s axiom plus 2^{ℵ_{0}} > ℵ_{1}, 31 (1978), 19–33

Avraham, U. and Shelah, S., Martin’s axiom does not imply that every two ℵ_{1}-dense

sets of reals are isomorphic, 38 (1981), 161–176

Bae, S. H. see Kim, J. M. Baernstein, A., II and Schober, G., Estimates for inverse coefficients of univalent

functions from integral means, 36 (1980), 75–82

Bagemihl, F., On the sharpness of Meier’s analogue of Fatou’s theorem, 4 (1966), 230–232 Baggett, L., On circle-valued cocycles of an ergodic measure-preserving transformation,

61 (1988), 29–38

Baggett, L. and Merrill, K., Smooth cocycles for an irrational rotation, 79 (1992),

281–288

Bahturin, Y., Fischman, D. and Montgomery, S., On the generalized Lie structure of

associative algebras, 96 (1996), 27–48

Baillon, J.-B. and Haddad, G., Quelques propriétés des opérateurs angle-bornés et

n-cycliquement monotones, 26 (1977), 137–150

Bainov, D., Domshlak, Yu. I. and Simeonov, P. S., On the oscillation properties of first-order impulsive differential equations with a deviating argument, 98 (1997), 167–187

Bainov, D., Minchev, E. and Myshkis, A., Monotone iterative method for semilinear elliptic systems of operator-differential equations, 92 (1995), 285–293 Bak, J., Leviatan, D., Newman, D. J. and Tzimbalario, J., Generalized polynomial

approximation, 15 (1973), 337–349

Bak, J., v. Golitschek, M. and Leviatan, D., The rate of approximation by means of

polynomials with restricted coefficients, 26 (1977), 265–275

Balcar, B. and Simon, P., Baire number of the spaces of uniform ultrafilters, 92 (1995),

263–272

Balder, E. J., On equivalence of strong and weak convergence in L_{1}-spaces under extreme

point conditions, 75 (1991), 21–47

Baldwin, S. and Beaudoin, R. E., Countable dense homogeneous spaces under Martin’s

axiom, 65 (1989), 153–164

Ball, K., Inequalities and sphere-packing in l_{p}, 58 (1987), 243–256 Ballester-Bolinches, A., 𝔥-normalizers and local definitions of saturated formations of

finite groups, 67 (1989), 312–326

Ballester-Bolinches, A., Remarks on formations, 73 (1991), 97–106 Ballester-Bolinches, A., Doerk, K. and Ezquerro, L. M., A question in the theory

of saturated formations of finite soluble groups, 84 (1993), 229–238

Barcelo, H. and Goupil, A., Non-broken circuits of reflection groups and factorization

in D_{n}, 91 (1995), 285–306

Barnette, D. W., 3-trees in polyhedral maps, 79 (1992), 251–256 Barnette, D. W., A construction of 3-connected graphs, 86 (1994), 397–407 Barnette, D. W., A family of neighborly polytopes, 39 (1981), 127–140 Barnette, D. W., A simple 4-dimensional nonfacet, 7 (1969), 16–20 Barnette, D. W., All triangulations of the projective plane are geometrically realizable

in E^{4}, 44 (1983), 75–87

Barnette, D. W., An invertible non-polyhedral diagram, 36 (1980), 89–96 Barnette, D. W., Closed 2-cell embeddings in the projective plane, 79 (1992), 161–172 Barnette, D. W., Decompositions of homology manifolds and their graphs, 41 (1982),

203–212

Barnette, D. W., Generating planar 4-connected graphs, 14 (1973), 1–13 Barnette, D. W., Generating the c^{∗}5-connected graphs, 28 (1977), 151–160 Barnette, D. W., Graph theorems for manifolds, 16 (1973), 62–72 Barnette, D. W., Inequalities for f-vectors of 4-polytopes, 11 (1972), 284–291 Barnette, D. W., Nonfacets for shellable spheres, 35 (1980), 286–288 Barnette, D. W., Polyhedral embeddings in the projective plane, 73 (1991), 91–95 Barnette, D. W., Projections of 3-polytopes, 8 (1970), 304–308 Barnette, D. W., The graphs of polytopes with involutory automorphisms, 9 (1971),

290–298

Barnette, D. W., The minimum number of vertices of a simple polytope, 10 (1971),

121–125

Barnette, D. W., Unique embeddings of simple projective plane polyhedral maps, 67

(1989), 251–256

Barnette, D. W. and Edelson, A. L., All 2-manifolds have finitely many minimal

triangulations, 67 (1989), 123–128

Barnette, D. W. and Edelson, A. L., All orientable 2-manifolds have finitely many

minimal triangulations, 62 (1988), 90–98

Barnette, D. W. and Wegner, G., Hamiltonian circuits in simple 3-polytopes with up

to 26 vertices, 19 (1974), 212–216

Baron, S., On a generalization of a theorem of B. Tyler, 43 (1982), 105–115 Barthélemy, L. and Bénilan, P., Sous-potentiels d’un opérateur nonlinéaire, 61 (1988),

85–112

Bartoszek, W., On convolution powers on semidirect products, 90 (1995), 115–123 Bartoszek, W., On the residuality of mixing by convolutions probabilities, 80 (1992),

183–193

Barwise, J., Some applications of Henkin quantifiers, 25 (1976), 47–63 Barwise, J. and Fisher, E. R., The Shoenfield absoluteness lemma, 8 (1970), 329–339 Barwise, J. and Kunen, K., Hanf numbers for fragments of L_{∞ω}, 10 (1971), 306–320 Bass, H. and Lubotzky, A., Automorphisms of groups and of schemes of finite type, 44

(1983), 1–22

Bastero, J., Embedding unconditional stable Banach spaces into symmetric stable Banach

spaces, 53 (1986), 373–380

Baston, V. and Beck, A., Generalizations in the linear search problem, 90 (1995),

301–323

Bates, S. M., On smooth, nonlinear surjections of Banach spaces, 100 (1997), 209–220 Baum, D. see Amitsur, S. A. Baumgartner, J. E., Partition relations for uncountable ordinals, 21 (1975), 296–307 Baur, W., Elimination of quantifiers for modules, 25 (1976), 64–70 Bayer, M. M., Equidecomposable and weakly neighborly polytopes, 81 (1993), 301–320 Bazzoni, S. and Salce, L., Groups in the class semigroups of valuation domains, 95

(1996), 135–155

Beaudoin, R. E. see Baldwin, S. Beaulieu, R. A., Jensen, A. and Jøndrup, S., Towards a more general notion of

Gelfand–Kirillov dimension, 93 (1996), 73–92

Beck, A., More on the linear search problem, 3 (1965), 61–70 Beck, A., On the linear search problem, 2 (1964), 221–228 Beck, A. and Beck, M., Son of the linear search problem, 48 (1984), 109–122 Beck, A. and Beck, M., The linear search problem rides again, 53 (1986), 365–372 Beck, A. and Newman, D. J., Yet more on the linear search problem, 8 (1970), 419–429 Beck, A. and Warren, P., The return of the linear search problem, 14 (1973), 169–183 Beck, A. see Baston, V. Beck, I., On modules whose endomorphism ring is local, 29 (1978), 393–407 Beck, M. see Beck, A. Becker, H., Determinacy implies that ℵ_{2} is supercompact, 40 (1981), 229–234 Bedi, S. S. and Ram, J., Jacobson radical of skew polynomial rings and skew group rings,

35 (1980), 327–338

Beer, G., Continuity of the straight line path for convex sets, 58 (1987), 205–212 Behrends, E., Normal operators and multipliers on complex Banach spaces and a symmetry

property of L^{1}-predual spaces, 47 (1984), 23–28

Behzad, M. and Radjavi, H., The line analog of Ramsey numbers, 5 (1967), 93–96 Bellenot, S. F., Banach spaces with trivial isometries, 56 (1986), 89–96 Bellenot, S. F., The Banach space T and the fast growing hierarchy from logic, 47

(1984), 305–313

Bellenot, S. F., Uniformly complemented l_{p}^{n}’s in quasi-reflexive Banach spaces, 39 (1981),

234–246

Beller, E., Factorization for non-Nevanlinna classes of analytic functions, 27 (1977),

320–330

Beller, E., Zeros of A^{p} functions and related classes of analytic functions, 22 (1975),

68–80

Bellow, A. and Furstenberg, H., An application of number theory to ergodic theory

and the construction of uniquely ergodic models, 33 (1979), 231–240

Ben-Artzi, M., A limiting absorption principle for Schrödinger operators with spherically

symmetric exploding potentials, 40 (1981), 259–274

Ben-David, S., Correction to: “On Shelah’s compactness of cardinals” [Israel J. Math. 31

(1978), no. 1, 34–56], 31 (1978), 394–394

Ben-David, S., On Shelah’s compactness of cardinals, 31 (1978), 34–56 Ben-David, S. and Shelah, S., Non-special Aronszajn trees on ℵ_{ω+1}, 53 (1986), 93–96 Ben-Israel, A., A modified Newton-Raphson method for the solution of systems of

equations, 3 (1965), 94–98 Ben-Israel, A., Erratum: “Theorems of the alternative for complex linear inequalities”,

7 (1969), 292–292

Ben-Israel, A., On direct sum decompositions of Hestenes algebras, 2 (1964), 50–54 Ben-Israel, A., On iterative methods for solving nonlinear least squares problems over

convex sets, 5 (1967), 211–224

Ben-Israel, A., The geometry of solvability and duality in linear programming, 1 (1963),

181–187

Ben-Israel, A., Theorems of the alternative for complex linear inequalities, 7 (1969),

129–136

Ben-Israel, A. see Zlobec, S. Benda, M. and Ketonen, J., Regularity of ultrafilters, 17 (1974), 231–240 Bender, H., Goldschmidt’s 2-signalizer functor theorem, 22 (1975), 208–213 Bénilan, P. and Diaz, J. I., Comparison of solutions of nonlinear evolution problems with

different nonlinear terms, 42 (1982), 241–257

Bénilan, P. see Barthélemy, L. Benjamini, I., On the support of harmonic measure for the random walk, 100 (1997), 1–6 Benjamini, I. and Peres, Y., On the Hausdorff dimension of fibres, 74 (1991), 267–279 Bennett, G., Dor, L. E., Goodman, V., Johnson, W. B. and Newman, C. M., On

uncomplemented subspaces of L_{p}, 1 < p < 2, 26 (1977), 178–187

Bensoussan, A. and Temam, R., Equations aux dérivées partielles stochastiques non

linéaires (1), 11 (1972), 95–129

Benyamini, Y., An extension theorem for separable Banach spaces, 29 (1978), 24–30 Benyamini, Y., An M-space which is not isomorphic to a C(K) space, 28 (1977), 98–102 Benyamini, Y., Constants of simultaneous extension of continuous functions, 16 (1973),

258–262

Benyamini, Y., Near isometries in the class of L^{1}-preduals, 20 (1975), 275–281 Benyamini, Y., Separable G spaces are isomorphic to C(K) spaces, 14 (1973), 287–293 Benyamini, Y. and Lin, P. K., An operator on L^{p} without best compact approximation,

51 (1985), 298–304

Benyamini, Y. and Lindenstrauss, J., A predual of l_{1} which is not isomorphic to a C(K)

space, 13 (1972), 246–254

Benyamini, Y. and Starbird, T., Embedding weakly compact sets into Hilbert space, 23

(1976), 137–141

Benyamini, Y. see Alspach, D. E. Benyamini, Y. see Argyros, S. A. Benyash-Krivetz, V. V. see Rapinchuk, A. S. Berbee, H., Periodicity and absolute regularity, 55 (1986), 289–304 Bercovici, H. and Foias, C., A real variable restatement of Riemann’s hypothesis, 48

(1984), 57–68

Berele, A., Cocharacters of Z/2Z-graded algebras, 61 (1988), 225–234 Berele, A., Homogeneous polynomial identities, 42 (1982), 258–272 Berele, A., Magnum P.I., 51 (1985), 13–19 Berele, A., On a theorem of Kemer, 51 (1985), 121–124 Berele, A., Giambruno, A. and Regev, A., Involution codimensions and trace codi-

mensions of matrices are asymptotically equal, 96 (1996), 49–62

Berele, A. and Regev, A., On the codimensions of the verbally prime P.I. algebras, 91

(1995), 239–247

Berele, A. and Saltman, D. J., The centers of generic division algebras with involution,

63 (1988), 98–118

Berend, D., A recurrence property of smooth functions, 62 (1988), 32–36 Berend, D. and Bergelson, V., Characterization of joint ergodicity for non-commuting

transformations, 56 (1986), 123–128

Berend, D. and Bergelson, V., Jointly ergodic measure-preserving transformations, 49 (1984), 307–314 Berend, D., Boshernitzan, M. D. and Kolesnik, G., Distribution modulo 1 of some

oscillating sequences, II, 92 (1995), 125–147

Berend, D. and Kolesnik, G., Distribution modulo 1 of some oscillating sequences, 71

(1990), 161–179

Berenstein, C. A. and Gay, R., A local version of the two-circles theorem, 55 (1986),

267–288

Berenstein, C. A. and Pascuas, D., Morera and mean-value type theorems in the

hyperbolic disk, 86 (1994), 61–106

Berg, I. D., Peck, N. T. and Porta, H., There is no continuous projection from the measurable functions on the square onto the measurable functions on the interval, 14 (1973), 199–204

Berg, K. see Auslander, J. Bergelson, V. see Berend, D. Bergen, J., Cohen, M. and Fischman, D., Irreducible actions and faithful actions of

Hopf algebras, 72 (1990), 5–18

Bergen, J. and Grzeszczuk, P., Invariants of Lie superalgebras acting on associative

algebras, 94 (1996), 403–428

Bergen, J. and Montgomery, S., Smash products and outer derivations, 53 (1986),

321–345

Bergen, J., Montgomery, S. and Passman, D. S., Radicals of crossed products of

enveloping algebras, 59 (1987), 167–184

Bergman, G. M., On the growth of algebras with bialgebra action, 96 (1996), 63–96 Bergman, G. M., Some examples in PI ring theory, 18 (1974), 257–277 Berk, R. H., A note on a Monte-Carlo estimator of Darling and Robbins, 7 (1969), 365–368 Berkovich, V. G., On the comparison theorem for étale cohomology of non-Archimedean

analytic spaces, 92 (1995), 45–59

Berkovich, Y. G., On the number of elements of given order in a finite p-group, 73 (1991),

107–112

Berman, D. M. see Albertson, M. O. Berman, J., A limit theorem for the moments of sums of independent random variables,

31 (1978), 383–393

Berman, S. and Moody, R. V., Extensions of Chevalley groups, 22 (1975), 42–51 Berndt, J., Prüfer, F. and Vanhecke, L., Geodesic spheres and two-point homogeneous

spaces, 93 (1996), 373–385

Berrick, A. J., Fibrations with Hopfian properties, 66 (1989), 41–53 Bers, L., A remark on Mumford’s compactness theorem, 12 (1972), 400–407 Bertram, E. A., Large centralizers in finite solvable groups, 47 (1984), 335–344 Bertram, E. A., Lower bounds for the number of conjugacy classes in finite solvable

groups, 75 (1991), 243–255

Bertrand-Mathis, A., Ensembles intersectifs et récurrence de Poincaré, 55 (1986),

184–198

Bessaga, C. and Klee, V., Two topological properties of topological linear spaces, 2

(1964), 211–220

Bessaga, C. and Pełczynski, A., On extreme points in separable conjugate spaces, 4

(1966), 262–264

Betke, U. and Gritzmann, P., A combinatorial condition for the existence of polyhedral

2-manifolds, 42 (1982), 297–299

Bezdek, K., On the illumination of unbounded closed convex sets, 80 (1992), 87–96 Bezdek, K., Kiss, G. and Mollard, M., An illumination problem for zonoids, 81 (1993),

265–272

Bhatt, R. C., A summation formula for Appell’s function F_{2}, 3 (1965), 87–88 Bhattacharjee, M., The probability of generating certain profinite groups by two

elements, 86 (1994), 311–329

Bialostocki, A., Nilpotent injectors in alternating groups, 44 (1983), 335–344 Bialostocki, A., Nilpotent injectors in symmetric groups, 41 (1982), 261–273 Bialostocki, A., On products of two nilpotent subgroups of a finite group, 20 (1975),

178–188

Bican, L. and Rangaswamy, K. M., PB^{∞}-subgroups of Butler groups, 95 (1996), 375–391 Bickel, P. J. and Yahav, J. A., The number of visits of vector walks to bounded regions,

3 (1965), 181–186

Biegert, W., Tauber-Konstanten für verschiedene Tauber-Bedingungen bei den Kreisver-

fahren der Limitierungstheorie, 4 (1966), 97–112

Bienenstock, E., Sets of degrees of computable fields, 27 (1977), 348–356 Bierstone, E. and Milman, P. D., Local analytic invariants and splitting theorems in

differential analysis, 60 (1987), 257–280

Bilu, Y., Effective analysis of integral points on algebraic curves, 90 (1995), 235–252 Binkowska, M. and Kamiński, B., Entropy increase for ℤ^{d}-actions on a Lebesgue space,

88 (1994), 307–318

Birman, J. S. and Wajnryb, B., Markov classes in certain finite quotients of Artin’s braid

group, 56 (1986), 160–178

Birman, J. S. see Wajnryb, B. Blair, C. E., An extension of a theorem of Jeroslow and Kortanek, 17 (1974), 111–115 Blair, D. E. and Chen, B.-Y., On CR-submanifolds of Hermitian manifolds, 34 (1979),

353–363

Blair, D. E., Koufogiorgos, T. and Papantoniou, B. J., Contact metric manifolds

satisfying a nullity condition, 91 (1995), 189–214

Blair, W. D. and Small, L. W., Embeddings in Artinian rings and Sylvester rank

functions, 58 (1987), 10–18

Blanchard, F., Dumont, J. M. and Thomas, A., Generic sequences, transducers and

multiplication of normal numbers, 80 (1992), 257–287

Blanchard, F. and Maass, A., Dynamical properties of expansive one-sided cellular

automata, 99 (1997), 149–174

Blasco, J. L. and Ivorra, C., Injective spaces of real-valued functions with the Baire

property, 91 (1995), 341–348

Blasius, D., On multiplicities for SL(n), 88 (1994), 237–251 Blass, A. and Shelah, S., Ultrafilters with small generating sets, 65 (1989), 259–271 Blatt, J. M. and Gutfreund, H., Minimum nastiness curve fitting, 6 (1968), 80–98 Blei, R. C. and Körner, T. W., Combinatorial dimension and random sets, 47 (1984),

65–74

Blind, G. and Blind, R., Convex polytopes without triangular faces, 71 (1990), 129–134 Blind, R. see Blind, G. Bloch, E. D., A combinatorial Chern–Weil theorem for 2-plane bundles with even Euler

characteristic, 67 (1989), 193–216

Blot, J., Calculus of variations in mean and convex Lagrangians, III, 67 (1989), 337–344 Blum, J. R., A note on mixing transformations, 9 (1971), 464–465 Blum, J. R. and Brennan, M., On the strong law of large numbers for dependent random

variables, 37 (1980), 241–245

Blum, J. R. and Epstein, B., On the Fourier transforms of an interesting class of measures,

10 (1971), 302–305

Blum, J. R. and Epstein, B., On the Fourier–Stieltjes coefficients of Cantor-type

distributions, 17 (1974), 35–45

Blum, J. R. and Reich, J. I., The individual ergodic theorem for p-sequences, 27 (1977),

180–184

Blum, J. R. and Reichaw, M., Correction: “Two integral inequalities”, Israel Journal of

Mathematics, Vol. 9, No. 1, 1971, pp. 20–26, 10 (1971), 259–259

Blum, J. R. and Reichaw, M., Two integral inequalities, 9 (1971), 20–26 Bogdanović, S. and Ćirić, M., Orthogonal sums of semigroups, 90 (1995), 423–428 Bogdanović, S. see Ćirić, M.

Bokor, I., On genus and cancellation in homotopy, 73 (1991), 361–379 Bokowski, J., Ewald, G. and Kleinschmidt, P., On combinatorial and affine automor-

phisms of polytopes, 47 (1984), 123–130

Bokowski, J. and Shemer, I., Neighborly 6-polytopes with 10 vertices, 58 (1987), 103–124 Bokowski, J. and Sturmfels, B., Polytopal and nonpolytopal spheres: an algorithmic

approach, 57 (1987), 257–271

Bokowski, J. see Altshuler, A. Bokut’, L. and Malcolmson, P., Gröbner–Shirshov bases for quantum enveloping

algebras, 96 (1996), 97–113

Bollobás, B. and Brightwell, G., Graphs whose every transitive orientation contains

almost every relation, 59 (1987), 112–128

Bollobás, B. and Erdős, P., Alternating Hamiltonian cycles, 23 (1976), 126–131 Bollobás, B., Gotsman, C. and Shamir, E., Connectivity and dynamics for random

subgraphs of the directed cube, 83 (1993), 321–328

Bollobás, B. and Rosenfeld, M., Sorting in one round, 38 (1981), 154–160 Bonnet, R. and Shelah, S., On HCO spaces. An uncountable compact T_{2} space, different

from ℵ_{1}+1, which is homeomorphic to each of its uncountable closed subspaces, 84 (1993), 289–332

Bonnice, W. E. and Reay, J. R., Interior points of convex hulls, 4 (1966), 243–248 Borel, A. and Schreiber, B. M., p-Adic linear groups with ergodic automorphisms, 24

(1976), 199–205

Borel, J.-P., Suites de longueur minimale associées à un ensemble normal donne, 64

(1988), 229–250

Borges, C. J. R., A remark on a problem of Mrówka, 3 (1965), 232–235 Borwein, J. M., Partially monotone operators and the generic differentiability of convex–

concave and biconvex mappings, 54 (1986), 42–50

Borwein, P. and Erdélyi, T., Notes on lacunary Müntz polynomials, 76 (1991), 183–192 Bose, C., Mixing examples in the class of piecewise monotone and continuous maps of the

unit interval, 83 (1993), 129–152

Boshernitzan, M. D. see Berend, D. Bourgain, J., A characterization of non-Dunford–Pettis operators on L^{1}, 37 (1980), 48–53 Bourgain, J., A new class of L^{1}-spaces, 39 (1981), 113–126 Bourgain, J., A remark on Schrödinger operators, 77 (1992), 1–16 Bourgain, J., A remark on the behaviour of L^{p}-multipliers and the range of operators

acting on L^{p}-spaces, 79 (1992), 193–206

Bourgain, J., A Szemerédi type theorem for sets of positive density in R^{k}, 54 (1986),

307–316

Bourgain, J., Almost sure convergence and bounded entropy, 63 (1988), 79–97 Bourgain, J., An averaging result for l^{1}-sequences and applications to weakly conditionally

compact sets in L_{X}^{1}, 32 (1979), 289–298

Bourgain, J., Construction of sets of positive measure not containing an affine image of a

given infinite structure, 60 (1987), 333–344

Bourgain, J., Dunford–Pettis operators on L^{1} and the Radon–Nikodym property, 37

(1980), 34–47

Bourgain, J., Estimates related to sumfree subsets of sets of integers, 97 (1997), 71–92 Bourgain, J., Hausdorff dimension and distance sets, 87 (1994), 193–201 Bourgain, J., Homogeneous polynomials on the ball and polynomial bases, 68 (1989),

327–347

Bourgain, J., On dentability and the Bishop–Phelps property, 28 (1977), 265–271 Bourgain, J., On Lipschitz embedding of finite metric spaces in Hilbert space, 52 (1985),

46–52

Bourgain, J., On Λ(p)-subsets of squares, 67 (1989), 291–311 Bourgain, J., On the L^{p}-bounds for maximal functions associated to convex bodies in R^{n},

54 (1986), 257–265 Bourgain, J., On the maximal ergodic theorem for certain subsets of the integers, 61

(1988), 39–72

Bourgain, J., On the pointwise ergodic theorem on L^{p} for arithmetic sets, 61 (1988),

73–84

Bourgain, J., On the primarity of H^{∞}-spaces, 45 (1983), 329–336 Bourgain, J., On the similarity problem for polynomially bounded operators on Hilbert

space, 54 (1986), 227–241

Bourgain, J., On the spectral type of Ornstein’s class one transformations, 84 (1993),

53–63

Bourgain, J., Problems of almost everywhere convergence related to harmonic analysis

and number theory, 71 (1990), 97–127

Bourgain, J., Ruzsa’s problem on sets of recurrence, 59 (1987), 150–166 Bourgain, J., Some consequences of Pisier’s approach to interpolation, 77 (1992), 165–185 Bourgain, J., The dimension conjecture for polydisc algebras, 48 (1984), 289–304 Bourgain, J., The metrical interpretation of superreflexivity in Banach spaces, 56 (1986),

222–230

Bourgain, J., Figiel, T. and Milman, V. D., On Hilbertian subsets of finite metric

spaces, 55 (1986), 147–152

Bourgain, J., Kahn, J., Kalai, G., Katznelson, Y. and Linial, N., The influence of

variables in product spaces, 77 (1992), 55–64

Bourgain, J. and Lindenstrauss, J., Distribution of points on spheres and approximation

by zonotopes, 64 (1988), 25–31

Bourgain, J. and Rosenthal, H. P., Martingales valued in certain subspaces of L^{1}, 37

(1980), 54–75

Bourgain, J. and Szarek, S. J., The Banach–Mazur distance to the cube and the

Dvoretzky-Rogers factorization, 62 (1988), 169–180

Bourgain, J. and Tzafriri, L., Embedding l_{p}^{k} in subspaces of L_{p} for p > 2, 72 (1990),

321–340

Bourgain, J. and Tzafriri, L., Invertibility of “large” submatrices with applications to

the geometry of Banach spaces and harmonic analysis, 57 (1987), 137–224

Bouscaren, E., Martin’s conjecture for ω-stable theories, 49 (1984), 15–25 Bousfield, A. K., Homotopy spectral sequences and obstructions, 66 (1989), 54–104 Boutellier, R. and Pfluger, A., Some extremal problems for functions of bounded

boundary rotation, 39 (1981), 46–62

Bowen, R., Bernoulli maps of the interval, 28 (1977), 161–168 Bowen, R., Smooth partitions of Anosov diffeomorphisms are weak Bernoulli, 21 (1975),

95–100

Bowen, R., Weak mixing and unique ergodicity on homogeneous spaces, 23 (1976), 267–273 Bowen, R. and Marcus, B., Unique ergodicity for horocycle foliations, 26 (1977), 43–67 Boyadzhiev, K. N., Sinclair type inequalities for the local spectral radius and related

topics, 57 (1987), 272–284

Boyarsky, A. see Góra, P. Boyarsky, A. see Proppe, H. Boyle, A. K. and Feller, E. H., The endomorphism ring of a Δ-module over a right

Noetherian ring, 45 (1983), 313–328

Boyle, M., Constraints on the degree of a sofic homomorphism and the induced multipli-

cation of measures on unstable sets, 53 (1986), 52–68

Boyle, M. and Handelman, D., Orbit equivalence, flow equivalence and ordered coho-

mology, 95 (1996), 169–210

Bramson, M. and Kalikow, S., Nonuniqueness in g-functions, 84 (1993), 153–160 Brandt, A., Interior estimates for second-order elliptic differential (or finite-difference)

equations via the maximum principle, 7 (1969), 95–121

Brandt, A., Interior Schauder estimates for parabolic differential- (or difference-) equations via the maximum principle, 7 (1969), 254–262 Braun, A., Localization, completion and the AR property in Noetherian P.I. rings, 96

(1996), 115–140

Braun, A., Localization in Noetherian P.I. rings as a finite criterion, 60 (1987), 54–64 Braun, A., On P.I. ring and standard identities, 55 (1986), 345–349 Braun, A., P.I. G-rings and the contractability of primes, 43 (1982), 116–128 Braun, A., P.I. rings and the localization at height 1 prime ideals, 62 (1988), 113–122 Breen, M., A decomposition theorem for m-convex sets, 24 (1976), 211–216 Breen, M., A Helly-number for k-almost-neighborly sets, 11 (1972), 347–348 Breen, M., An example concerning unions of two starshaped sets in the plane, 17 (1974),

347–349

Breen, M., An R^{d} analogue of Valentine’s theorem on 3-convex sets, 24 (1976), 206–210 Breen, M., Essential and inessential points of local nonconvexity, 12 (1972), 347–355 Breen, M., Intersectional properties concerning planar m-convex sets, 36 (1980), 83–88 Breen, M., Primitive Radon partitions for cyclic polytopes, 15 (1973), 156–157 Breen, M., The combinatorial structure of (m, n)-convex sets, 15 (1973), 367–374 Breen, M., Tilings whose members have finitely many neighbors, 52 (1985), 140–146 Breen, M. and Kay, D. C., General decomposition theorems for m-convex sets in the

plane, 24 (1976), 217–233

Brehm, U. and Altshuler, A., On weakly neighborly polyhedral maps of arbitrary genus,

53 (1986), 137–157

Brehm, U. and Schild, G., Realizability of the torus and the projective plane in ℝ^{4}, 91

(1995), 249–251

Brehm, U. see Altshuler, A. Brendle, J. and Judah, H., Perfect sets of random reals, 83 (1993), 153–176 Brennan, M. see Blum, J. R. Brezin, J., Ellis, R. and Shapiro, L., Recognizing G-induced flows, 17 (1974), 56–65 Brezis, H., Equations d’évolution du second ordre associées à des opérateurs monotones,

12 (1972), 51–60

Brezis, H., Intégrales convexes dans les espaces de Sobolev, 13 (1972), 9–23 Brezis, H., Propriétés régularisantes de certains semi-groupes non linéaires, 9 (1971),

513–534

Brezis, H. and Haraux, A., Image d’une somme d’opérateurs monotones et applications,

23 (1976), 165–186

Brezis, H. and Lions, P. L., Produits infinis de resolvantes, 29 (1978), 329–345 Brezis, H. and Pazy, A., Accretive sets and differential equations in Banach spaces, 8

(1970), 367–383

Bridges, W. G., On the replications of certain multiplicative designs, 12 (1972), 369–372 Brightwell, G. see Bollobás, B. Brin, M., On the spectrum of Anosov diffeomorphisms, 36 (1980), 201–204 Briskin, M. and Yomdin, Y., Critical and near-critical values in polynomial control

problems, I: One-dimensional case, 78 (1992), 257–280

Brodsky, S. D. and Howie, J., The universal torsion-free image of a group, 98 (1997),

209–228

Brown, D. J. and Loeb, P. A., The values of nonstandard exchange economies, 25

(1976), 71–86

Brown, E. and Johnson, L., A class of planar four-colorable graphs, 11 (1972), 53–56 Brown, K. S. and Dror, E., The Artin-Rees property and homology, 22 (1975), 93–109 Brown, L. and Ito, T., Some non-quasireflexive spaces having unique isomorphic preduals,

20 (1975), 321–325

Brown, L. and Ito, T., Uniqueness of preduals of certain Banach spaces, 24 (1976),

234–243

Brownell, F. H., The uniqueness of known Dirac atomic eigenfunctions, 9 (1971), 172–179 Brozovic, D. P. and Solomon, R. M., On groups of hyperbolic length, 98 (1997), 61–99 Bruck, R. E., A simple proof of the mean ergodic theorem for nonlinear contractions in

Banach spaces, 32 (1979), 107–116

Bruck, R. E., On the almost-convergence of iterates of a nonexpansive mapping in Hilbert

space and the structure of the weak ω-limit set, 29 (1978), 1–16

Bruck, R. E., On the convex approximation property and the asymptotic behavior of

nonlinear contractions in Banach spaces, 38 (1981), 304–314

Bruckner, A. M., A category analogue of a theorem on metric density, 11 (1972), 249–253 Bruckner, A. M. and Bruckner, J. B., Generalized convex kernels, 2 (1964), 27–32 Bruckner, J. B. see Bruckner, A. M. Brunel, A. and Revuz, D., Sur la théorie du renouvellement pour les groupes non

abéliens, 20 (1975), 46–56

Brunner, G., Eine Verallgemeinerung der CW-Komplexe, 12 (1972), 306–313 Brussel, E. S., Decomposability and embeddability of discretely Henselian division

algebras, 96 (1996), 141–183

Brustein, R., Ne’eman, Y. and Sternberg, S., Duality, crossing and MacLane’s

coherence, 72 (1990), 19–37

Bruzzo, U., Berezin integration on DeWitt supermanifolds, 80 (1992), 161–169 Buchi, J. R. and Danhof, K. J., Definability in normal theories, 14 (1973), 248–256 Buechler, S., Invariants for ω-categorical, ω-stable theories, 52 (1985), 65–81 Buechler, S., Resplendency and recursive definability in ω-stable theories, 49 (1984),

26–33

Bueso, J. L., Jara, P. and Torrecillas, B., Decomposition of injective modules relative

to a torsion theory, 52 (1985), 266–272

Burak, T., On semigroups generated by restrictions of elliptic operators to invariant

subspaces, 12 (1972), 79–93

Burak, T., Regularity properties of solutions of some abstract parabolic equations, 16

(1973), 418–445

Burak, T., Two point problems and analyticity of solutions of abstract parabolic equations,

16 (1973), 404–417

Burke, D. and Michael, E., On a theorem of V. V. Filippov, 11 (1972), 394–397 Burke, M. R. and Just, W., Liftings for Haar measure on {0, 1}^{κ}, 73 (1991), 33–44 Burke, M. R. and Shelah, S., Linear liftings for non-complete probability spaces, 79

(1992), 289–296

Burns, D. and Schuster, S., Embedding (p, p − 1) graphs in their complements, 30

(1978), 313–320

Burshtein, N. see Schönheim, J. Burton, G. R., Some characterisations of the ellipsoid, 28 (1977), 339–349 Burton, R. M., A non-Bernoulli skew product which is loosely Bernoulli, 35 (1980),

339–348

Burton, R. M. and Steif, J. E., New results on measures of maximal entropy, 89 (1995),

275–300

Burton, R. M., Denker, M. and Smorodinsky, M., Finite state bilaterally deterministic

strongly mixing processes, 95 (1996), 115–133

Burton, R. M. and Keane, M., Topological and metric properties of infinite clusters in

stationary two-dimensional site percolation, 76 (1991), 299–316

Buzzi, J., Intrinsic ergodicity of smooth interval maps, 100 (1997), 125–161 Byers, W. see Proppe, H. Caenepeel, S., Computing the Brauer–Long group of a Hopf algebra. I: The cohomological

theory, 72 (1990), 38–83

Caenepeel, S., Militaru, G. and Zhu, S., Crossed modules and Doi–Hopf modules, 100

(1997), 221–247

Camina, A. R., Some conditions which almost characterize Frobenius groups, 31 (1978), 153–160 Campbell, C. M. and Thomas, R. M., On (2, n)-groups related to Fibonacci groups, 58

(1987), 370–380

Campbell, D. M., Eventually areally mean p-valent functions, 16 (1973), 216–236 Camps, R. and Dicks, W., On semilocal rings, 81 (1993), 203–211 Canham, R. J., A theorem on arrangements of lines in the plane, 7 (1969), 393–397 Cannarsa, P. and Vespri, V., Generation of analytic semigroups in the L^{p} topology by

elliptic operators in R^{n}, 61 (1988), 235–255

Capon, M., Primarité de L^{p}(l_{r}), 1 < p, r < ∞, 36 (1980), 346–364 Capon, M., Primarité de L^{p}(L^{r}), 1 < p, r < ∞, 42 (1982), 87–98 Carl, B. and Defant, A., An inequality between the p- and (p, 1)-summing norm of finite

rank operators from C(K)-spaces, 74 (1991), 323–335

Carlson, T., Modal logics with several operators and provability interpretations, 54

(1986), 14–24

Caro, Y., On a division property of consecutive integers, 33 (1979), 32–36 Caro, Y., Zero-sum subsequences in abelian non-cyclic groups, 92 (1995), 221–233 Carothers, N. L., Rearrangement invariant subspaces of Lorentz function spaces, 40

(1981), 217–228

Carothers, N. L., Haydon, R. and Lin, P.-K., On the isometries of the Lorentz function

spaces, 84 (1993), 265–287

Carothers, N. L. see Alspach, D. E. Carter, J. E. see Vanstone, S. A. Carter, M. W. see Vanstone, S. A. Casadio Tarabusi, E., Cohen, J. M. and Picardello, M. A., The horocyclic Radon

transform on non-homogeneous trees, 78 (1992), 363–380

Casazza, P. G., James’ quasi-reflexive space is primary, 26 (1977), 294–305 Casazza, P. G., Johnson, W. B. and Tzafriri, L., On Tsirelson’s space, 47 (1984),

81–98

Casazza, P. G. and Kalton, N. J., Unconditional bases and unconditional finite-

dimensional decompositions in Banach spaces, 95 (1996), 349–373

Casazza, P. G. and Lin, B. L., On symmetric basic sequences in Lorentz sequence spaces

II, 17 (1974), 191–218

Casazza, P. G., Pengra, R. W. and Sundberg, C., Complemented ideals in the disk

algebra, 37 (1980), 76–83

Casazza, P. G. see Altshuler, Z. Caselles, V., On the peripheral spectrum of positive operators, 58 (1987), 144–160 Castillo, J. M. F. and González, M., The Dunford–Pettis property is not a three-space

property, 81 (1993), 297–299

Cavicchioli, A., Covering numbers of manifolds and critical points of a Morse function,

70 (1990), 279–304

Cawley, E., Marcus, B. and Tuncel, S., Boundary measures of Markov chains, 94

(1996), 111–123

Ceder, J. G., A property of planar convex bodies, 1 (1963), 248–253 Cellina, A. and Marchi, M. V., Non-convex perturbations of maximal monotone

differential inclusions, 46 (1983), 1–11

Censor, Y., On the maximal number of solutions of a problem in linear inequalities, 9

(1971), 27–33

Cernes, A., Ensembles maximaux accrétifs et m-accrétifs, 19 (1974), 335–348 Chabrowski, J., Quasilinear ellipticity and the Dirichlet problem, 63 (1988), 353–379 Chacon, R. V. and Jamison, B., A fundamental property of Markov processes with an

application to equivalence under time changes, 33 (1979), 241–269

Chakerian, G. D., The affine image of a convex body of constant breadth, 3 (1965), 19–22 Chakerian, G. D., The mean volume of boxes and cylinders circumscribed about a convex

body, 12 (1972), 249–256 Chakerian, G. D. and Groemer, H., On classes of convex sets that permit plane

coverings, 19 (1974), 305–312

Chapman, S. T. and Smith, W. W., Factorization in Dedekind domains with finite class

group, 71 (1990), 65–95

Charbonnel, A.-M., Calcul fonctionnel a plusieurs variables pour des opérateurs

pseudodifférentiels dans R^{n}, 45 (1983), 69–89

Chartrand, G., Gould, R. J. and Polimeni, A. D., A note on locally connected and

hamiltonian-connected graphs, 33 (1979), 5–8

Chartrand, G., Kronk, H. V. and Wall, C. E., The point-arboricity of a graph, 6

(1968), 169–175

Chatzidakis, Z. M., A projective profinite group whose smallest embedding cover is not

projective, 85 (1994), 1–9

Chatzidakis, Z. M., Some properties of the absolute Galois group of a Hilbertian field,

55 (1986), 173–183

Chatzidakis, Z. M., Some remarks on profinite HNN extensions, 85 (1994), 11–18 Cheatham, T., The quotient field as a torsion-free covering module, 33 (1979), 172–176 Cheatham, T. J., Enochs, E. E. and Jenda, O. M. G., The structure of injective covers

of special modules, 63 (1988), 237–242

Chen, B.-Y., Interaction of Legendre curves and Lagrangian submanifolds, 99 (1997),

69–108

Chen, B.-Y., Submanifolds in de Sitter space-time satisfying ΔH = λH, 91 (1995),

373–391

Chen, B.-Y. see Blair, D. E. Chen, C. S. see Wang, A. T. Chen, J.-C., Weights and L_{Φ}-boundedness of the Poisson integral operator, 81 (1993),

193–202

Chen, R., A finitely additive version of Kolmogorov’s law of the iterated logarithm, 23

(1976), 209–220

Chen, R., A note on the Dubins-Savage utility of a strategy, 21 (1975), 1–6 Chen, R., Some finitely additive versions of the strong law of large numbers, 24 (1976),

244–259

Chen, Z.-M., A remark on flow invariance for semilinear parabolic equations, 74 (1991),

257–266

Cherlin, G. L., Amalgamation bases for commutative rings without nilpotent elements,

25 (1976), 87–96

Chernousov, V. I. see Rapinchuk, A. S. Chiappinelli, R., Remarks on bifurcation for elliptic operators with odd nonlinearity, 65

(1989), 285–292

Chiappinelli, R. and Edmunds, D. E., Eigenvalue asymptotics and a non-linear

Schrödinger equation, 81 (1993), 179–192

Chicone, C., The topology of stationary curl parallel solutions of Euler’s equations, 39

(1981), 161–166

Chilin, V. I., Dodds, P. G. and Sukochev, F. A., The Kadec–Klee property in

symmetric spaces of measurable operators, 97 (1997), 203–219

Chillag, D., A note on transitive permutation groups of prime degree, 21 (1975), 50–53 Chillag, D., On a class of transitive permutation groups of prime degree p = 4n + 1, 15

(1973), 78–91

Chillag, D. and Macdonald, I. D., Generalized Frobenius groups, 47 (1984), 111–122 Chillag, D., Mann, A. and Manz, O., The co-degrees of irreducible characters, 73

(1991), 207–223

Chillag, D., Mann, A. and Scoppola, C. M., Generalized Frobenius groups. II, 62

(1988), 269–282

Chillag, D. and Sonn, J., Sylow-metacyclic groups and Q-admissibility, 40 (1981), 307–323 Chin, W., Prime ideals in restricted differential operator rings, 60 (1987), 236–256 Chin, W., Spectra of smash products, 72 (1990), 84–98 Chinak, M., The fixed point statements on Bergman type complex manifolds, 97 (1997),

293–304

Chong, C. T., Major subsets of α-recursively enumerable sets, 34 (1979), 106–114 Chothi, V., Everest, G. and Ward, T., Oriented local entropies for expansive actions

by commuting automorphisms, 93 (1996), 281–301

Choucroun, F. M., Sous-groupes discrets des groupes 𝔭-adiques de rang un et arbres de

Bruhat-Tits, 93 (1996), 195–219

Choueka, Y., Bounded sets of nodes in transfinite trees, 11 (1972), 254–257 Chow, Y. S. and Dvoretzky, A., Stopping rules for x_{n}/n and related problems, 7 (1969),

240–248

Chow, Y. S., Moriguti, S., Robbins, H. and Samuels, S. M., Optimal selection based

on relative rank (the “Secretary Problem”), 2 (1964), 81–90

Christensen, J. P. R., Measurability problems in a metrizable convex compact set, 12

(1972), 201–203

Christensen, J. P. R., On sets of Haar measure zero in abelian Polish groups, 13 (1972),

255–260

Christensen, J. P. R. see Andersen, N. J. M. Chuang, C.-L., On invariant additive subgroups, 57 (1987), 116–128 Chuang, C.-L., The additive subgroup generated by a polynomial, 59 (1987), 98–106 Chuaqui, M. and Osgood, B., An extension of a theorem of Gehring and Pommerenke,

91 (1995), 393–407

Cima, J. A. and Rung, D. C., Normal functions and a class of associated boundary

functions, 4 (1966), 119–126

Ćirić, M. and Bogdanović, S., The lattice of positive quasi-orders on a semigroup, 98

(1997), 157–166

Ćirić, M. see Bogdanović, S. Clarke, C. J. S., On rigged immersions, 41 (1982), 347–366 Clarke, C. J. S., The cardinality of manifold atlases, 45 (1983), 9–16 Clément, P., On abstract Volterra equations with kernels having a positive resolvent, 36

(1980), 193–200

Clément, P. and Mitidieri, E., Qualitative properties of solutions of Volterra equations

in Banach spaces, 64 (1988), 1–24

Cobos, F. and Peetre, J., Interpolation of compactness using Aronszajn-Gagliardo

functors, 68 (1989), 220–240

Cochet, Y. and Nivat, M., Une généralisation des ensembles de Dyck, 9 (1971), 389–395 Coelho, Z. and de Faria, E., Limit laws of entrance times for homeomorphisms of the

circle, 93 (1996), 93–112

Coelho, Z. and Parry, W., Central limit asymptotics for shifts of finite type, 69 (1990),

235–249

Cohen, A. M. and Wales, D. B., Embeddings of the group L(2, 13) in groups of Lie type

E_{6}, 82 (1993), 45–86

Cohen, F. R. and Peterson, F. P., On the homology of certain spaces looped beyond

their connectivity, 66 (1989), 105–131

Cohen, J. M. and Colonna, F., Extreme points of the Bloch space of a homogeneous

tree, 94 (1996), 247–271

Cohen, J. M. and Picardello, M. A., The 2-circle and 2-disk problems on trees, 64

(1988), 73–86

Cohen, J. M. see Casadio Tarabusi, E. Cohen, M., Addendum to “Semiprime Goldie centralizers”, 24 (1976), 89–93 Cohen, M., Semiprime Goldie centralizers, 20 (1975), 37–45 Cohen, M., Raianu, S. and Westreich, S., Semiinvariants for Hopf algebra actions, 88

(1994), 279–306 Cohen, M., Westreich, S. and Zhu, S., Determinants, integrality and Noether’s theorem

for quantum commutative algebras, 96 (1996), 185–222

Cohen, M. see Bergen, J. Cohen, S. D., Likeable functions in finite fields, 46 (1983), 123–126 Cohn, H., A ratio limit theorem for the finite nonhomogeneous Markov chains, 19 (1974),

329–334

Cohn, H., On the Borel-Cantelli lemma, 12 (1972), 11–16 Cohn, P. M., On skew field coproducts with a finite extension, 96 (1996), 223–230 Cohn, P. M., Progress in free associative algebras, 19 (1974), 109–151 Colbourn, C. J. see Vanstone, S. A. Collet, P., Martínez, S. and Schmitt, B., The Pianigiani–Yorke measure for topological

Markov chains, 97 (1997), 61–70

Collier, J. B. and Edelstein, M., Common supports for pairs of sets, 19 (1974), 237–240 Collier, J. B. and Edelstein, M., On strongly exposed points and Fréchet differentia-

bility, 17 (1974), 66–68

Colombeau, J. F. and Paques, O. T. W., Structure of spaces of C^{∞}-functions on nuclear

spaces, 45 (1983), 17–32

Colombo, G., Fonda, A. and Ornelas, A., Lower semicontinuous perturbations of

maximal monotone differential inclusions, 61 (1988), 211–218

Colonna, F. see Cohen, J. M. Connes, A. and Weiss, B., Property T and asymptotically invariant sequences, 37 (1980),

209–210

Connes, B., Mesure de Hausdorff des complémentaires des U_{ε} de Zygmund, 23 (1976), 1–7 Converse, G., Sets with a mode, 27 (1977), 124–130 Conze, J. P. and Glasner, S., Almost periodic measures on the torus, 33 (1979), 270–280 Cook, R. J., Compounds of asymmetric convex bodies, 19 (1974), 349–358 Cooper, S. B., Lempp, S. and Watson, P., Weak density and cupping in the d-r.e.

degrees, 67 (1989), 137–152

Cooperstein, B. N., Minimal degree for a permutation representation of a classical group,

30 (1978), 213–235

Cornet, B. and Haddad, G., Théorèmes de viabilité pour les inclusions différentielles du second ordre. [Viability theorems for second-order differential inclusions], 57 (1987), 225–238

Coulhon, T., Noyau de la chaleur et discrétisation d’une variété riemannienne, 80 (1992),

289–300

Coven, E. M. and Paul, M. E., Sofic systems, 20 (1975), 165–177 Covitz, H. and Nadler, Jr., S. B., Multi-valued contraction mappings in generalized

metric spaces, 8 (1970), 5–11

Cowen, R., On the classification of one-sided Markov chains, 67 (1989), 243–250 Craizer, M., Entropy of inner functions, 74 (1991), 129–168 Crandall, M. G., The semigroup approach to first order quasilinear equations in several

space variables, 12 (1972), 108–132

Crandall, M. G. and Evans, L. C., On the relation of the operator ∂/∂s + ∂/∂τ to

evolution governed by accretive operators, 21 (1975), 261–278

Crandall, M. G. and Nohel, J. A., An abstract functional differential equation and a

related nonlinear Volterra equation, 29 (1978), 313–328

Crandall, M. G. and Pazy, A., Nonlinear evolution equations in Banach spaces, 11

(1972), 57–94

Crandall, M. G. and Pazy, A., On the range of accretive operators, 27 (1977), 235–246 Crandall, M. G., Pazy, A. and Tartar, L., Remarks on generators of analytic

semigroups, 32 (1979), 363–374

Crone, L. and Robinson, W., Diagonal maps and diameters in Köthe spaces, 20 (1975), 13–22 Cunningham, F., Jr., Effros, E. G. and Roy, N. M., M-structure in dual Banach

spaces, 14 (1973), 304–308

Currás-Bosch, C., A_{X}-operator on complete Riemannian manifolds, 53 (1986), 315–320 Da Prato, G. and Sinestrari, E., Hölder regularity for non-autonomous abstract

parabolic equations, 42 (1982), 1–19

Dacunha-Castelle, D., Sur un théorème de J. L. Krivine concernant la caracterisation des classes d’espaces isomorphes à des espaces d’Orlicz généralisés et des classes voisines, 13 (1972), 261–276

Dacunha-Castelle, D. and Krivine, J.-L., Sous-espaces de L^{1}, 26 (1977), 320–351 Dalla, L., On a class of some special sets on the k-skeleton of a convex compact set, 68

(1989), 353–364

Damlamian, A. see Attouch, H. Dang-Ngoc, N., Pointwise convergence of martingales in von Neumann algebras, 34

(1979), 273–280

Danhof, K. J. see Buchi, J. R. Dani, S. G., On ergodic quasi-invariant measures of group automorphism, 43 (1982), 62–74 Dani, S. G. and Raghavan, S., Orbits of Euclidean frames under discrete linear groups,

36 (1980), 300–320

Dar, S., On the isotropic constant of non-symmetric convex bodies, 97 (1997), 151–156 Dash, A. T. and Schechter, M., Tensor products and joint spectra, 8 (1970), 191–193 Dashiell, F. K. and Lindenstrauss, J., Some examples concerning strictly convex norms

on C(K) spaces, 16 (1973), 329–342

Datta, S., Secondary flow about a magnetized sphere rotating in viscous conducting fluid,

3 (1965), 89–93

David, M., On a certain type of commutators of operators, 9 (1971), 34–42 David, R., A π_{2}^{1} singleton with no sharp in a generic extension of L^{#}, 31 (1978), 343–352 Davis, B., On the integrability of the martingale square function, 8 (1970), 187–190 Davis, W. J., Embedding spaces with unconditional bases, 20 (1975), 189–191 Davis, W. J., Dean, D. W. and Singer, I., Complemented subspaces and Λ systems in

Banach spaces, 6 (1968), 303–309

Davis, W. J. and Johnson, W. B., Basic sequences and norming subspaces in non-quasi-

reflexive Banach spaces, 14 (1973), 353–367

Davis, W. J., Milman, V. D. and Tomczak-Jaegermann, N., The distance between

certain n-dimensional Banach spaces, 39 (1981), 1–15

Day, M. M., A geometric proof of Asplund’s differentiability theorem, 13 (1972), 277–280 De Concini, C. and Kazhdan, D., Special bases for S_{N} and GL(n), 40 (1981), 275–290 Dean, D. W. see Davis, W. J. Deb, M. see Singal, M. K. Debbi, D. and Gillis, J., An inequality relating to triple product integrals of Laguerre

functions, 18 (1974), 45–52

Defant, A. see Carl, B. Defant, M. see Junge, M. Dekster, B. V., An extension of Jung’s theorem, 50 (1985), 169–180 Dekster, B. V., On reduced convex bodies, 56 (1986), 247–256 deLaubenfels, R., C-semigroups and strongly continuous semigroups, 81 (1993), 227–255 deLaubenfels, R., Huang, Z., Wang, S. and Wang, Y., Laplace transforms of

polynomially bounded vector-valued functions and semigroups of operators, 98 (1997), 189–207

deLeeuw, K. and Katznelson, Y., On certain homomorphisms of quotients of group

algebras, 2 (1964), 120–126

deLeeuw, K. and Katznelson, Y., The two sides of a Fourier–Stieltjes transform and

almost idempotent measures, 8 (1970), 213–229

Deliyanni, I. see Argyros, S. A. Demko, S. see Andrew, A. D.

Denker, M. and Keane, M., Almost topological dynamical systems, 34 (1979), 139–160 Denker, M. and Urbański, M., Hausdorff measures on Julia sets of subexpanding rational

maps, 76 (1991), 193–214

Denker, M. see Burton, R. M. Desch, W., Lasiecka, I. and Schappacher, W., Feedback boundary control problems

for linear semigroups, 51 (1985), 177–207

Deshouillers, J-M. and Freiman, G. A., On an additive problem of Erdős and Straus,

1, 92 (1995), 33–43

Deshpande, M. G. and Feller, E. H., Endomorphism rings of essential extensions of a

Noetherian module, 17 (1974), 46–49

Desmarais, P. C. and Martindale, W. S., 3rd, Generalized rational identities and

rings with involution, 36 (1980), 187–192

Desmarais, P. C. and Martindale, W. S., 3rd, Primitive rings with involution and

pivotal monomials, 22 (1975), 118–126

Deuel, J. and Hess, P., Nonlinear parabolic boundary value problems with upper and

lower solutions, 29 (1978), 92–104

deValcourt, B. A., Measures of axial symmetry for ovals, 4 (1966), 65–82 Deville, R., Geometrical implications of the existence of very smooth bump functions in

Banach spaces, 67 (1989), 1–22

Deville, R., Godefroy, G., Hare, D. E. G. and Zizler, V., Differentiability of convex functions and the convex point of continuity property in Banach spaces, 59 (1987), 245–255

Devlin, K. J., Iterated Souslin forcing, the principles ♢(E) and a generalisation of the

axiom SAD, 31 (1978), 368–382

Devlin, K. J. and Shelah, S., A weak version of ♢ which follows from 2^{ℵ_{0}} < 2^{ℵ_{1}}, 29

(1978), 239–247

Devlin, K. J. see Avraham, U. Dharmadhikari, S. W. and Jogdeo, K., A characterization of convexity and central

symmetry for planar polygonal sets, 15 (1973), 356–366

Di Blasio, G., Kunisch, K. and Sinestrari, E., Stability for abstract linear functional

differential equations, 50 (1985), 231–263

Di Vincenzo, O. M., On the graded identities of M_{1,1}(E), 80 (1992), 323–335 Diaz, J. I. see Bénilan, P. Dicks, W. see Camps, R. Diderrich, G. T., Sums of length t in Abelian groups, 14 (1973), 14–22 Dillen, F., Semi-parallel hypersurfaces of a real space form, 75 (1991), 193–202 Dillen, F., Petrovic, M. and Verstraelen, L., Einstein, conformally flat and semi-

symmetric submanifolds satisfying Chen’s equality, 100 (1997), 163–169

Dilworth, S. J. and Szarek, S. J., The cotype constant and an almost Euclidean

decomposition for finite-dimensional normed spaces, 52 (1985), 82–96

Dilworth, S. J., Universal non-compact operators between super-reflexive Banach spaces

and the existence of a complemented copy of Hilbert space, 52 (1985), 15–27

Dilworth, S. J. and Koldobsky, A. L., The Fourier transform of order statistics with

applications to Lorentz spaces, 92 (1995), 411–425

Dineen, S. and Timoney, R. M., Irreducible domains in Banach spaces, 57 (1987),

327–346

Dineen, S. see Nachbin, L. Dinghas, A., Über das Schwarzsche Lemma und verwandte Sätze, 5 (1967), 157–169 Dinghas, A., Zur Peano-Abbildung einer Strecke, 7 (1969), 211–216 Ditzian, Z., An inversion formula for some convolution transforms, 6 (1968), 217–222 Ditzian, Z., Exponential formulae for semi-group of operators in terms of the resolvent, 9

(1971), 541–553

Ditzian, Z., Polynomials of best approximation in C[−1, 1], 52 (1985), 341–354 Ditzian, Z., Saturation classes for a sequence of convolution operators with limited

oscillation, 17 (1974), 315–324

Ditzian, Z. and Jakimovski, A., Real inversion and jump formula for the Laplace

transform, Part II, 2 (1964), 33–40

Ditzian, Z. and Jakimovski, A., Real inversion and jump formulae for the Laplace

transform. Part I., 1 (1963), 85–104

Djoković, D. Ž., Irreducible connected Lie subgroups of GL_{n}(R) are closed, 28 (1977),

175–176

Dobbs, D. E., Amitsur cohomology of cubic extensions of algebraic integers, 14 (1973),

213–220

Dodds, P. G. and Fremlin, D. H., Compact operators in Banach lattices, 34 (1979),

287–320

Dodds, P. G. see Chilin, V. I. Doerk, K. see Ballester-Bolinches, A. Doi, Y., Hopf extensions of algebras and Maschke type theorems, 72 (1990), 99–108 Domański, P., A note on injective C(T)-spaces and the Amir boundary, 73 (1991), 357–359 Domar, Y., Estimates of ∥e^{itf}∥_{A(Γ)}, when Γ ⊂ ℝ^{n} is a curve and f is a real-valued function,

12 (1972), 184–189

Domshlak, Yu. I. see Bainov, D. Donder, H.-D., Regularity of ultrafilters and the core model, 63 (1988), 289–322 Donder, H.-D. and Levinski, J.-P., On weakly precipitous filters, 67 (1989), 225–242 Donder, H.-D. and Matet, P., Two cardinal versions of diamond, 83 (1993), 1–43 Dong, X. H., On a theorem of Bazilevic for areally mean p-valent functions attaining

maximal growth on k rays, 100 (1997), 327–337

Donin, J. and Gurevich, D., Some Poisson structures associated to Drinfeld–Jimbo

R-matrices and their quantization, 92 (1995), 23–32

Donin, J. and Shnider, S., Quasi-associativity and flatness criteria for quadratic algebra

deformations, 97 (1997), 125–149

Donoghue, W. F., Jr., The theorems of Loewner and Pick, 4 (1966), 153–170 Dor, L. E., A geometric characterization of inner product spaces, 37 (1980), 84–91 Dor, L. E., A note on monotone bases, 14 (1973), 285–286 Dor, L. E., On good L_{p} subspaces of l_{p}, 44 (1983), 303–316 Dor, L. E., Potentials and isometric embeddings in L_{1}, 24 (1976), 260–268 Dor, L. E. see Bennett, G. Dow, A., Some linked subsets of posets, 59 (1987), 353–376 Dow, A., Juhász, I. and Weiss, W., Integer-valued functions and increasing unions of

first countable spaces, 67 (1989), 181–192

Downarowicz, T., The Choquet simplex of invariant measures for minimal flows, 74

(1991), 241–256

Downarowicz, T., The royal couple conceals their mutual relationship: A noncoalescent

Toeplitz flow, 97 (1997), 239–251

Downey, R. and Shore, R. A., Lattice embeddings below a nonlow_{2} recursively enumer-

able degree, 94 (1996), 221–246

Downey, R. and Stob, M., Minimal pairs in initial segments of the recursively enumerable

degrees, 100 (1997), 7–27

Doyle, P. G., Mallows, C., Orlitsky, A. and Shepp, L. A., On the evolution of

islands, 67 (1989), 34–42

Doyle, P. H., On shrinking arcs in metric spaces, 5 (1967), 104–106 Dragomir, S., On submanifolds of Hopf manifolds, 61 (1988), 199–210 Drensky, V., New central polynomials for the matrix algebras, 92 (1995), 235–248 Drensky, V. and Regev, A., Exact asymptotic behaviour of the codimensions of some

P.I. algebras, 96 (1996), 231–242

Dressler, R. E. and Kirk, R. B., Non-measurable sets of reals whose measurable subsets are countable, 11 (1972), 265–270 van den Dries, L. and Miller, C., Correction to “On the real exponential field with

restricted analytic functions” [Israel J. Math. 85 (1994), pp. 19–56], 92 (1995), 427–427

van den Dries, L. and Miller, C., On the real exponential field with restricted analytic

functions, 85 (1994), 19–56

van den Dries, L. see Lubotzky, A. Dror, E., Homology spheres, 15 (1973), 115–129 Dror, E. and Dwyer, W. G., A Whitehead theorem for long towers of spaces, 29 (1978),

141–154

Dror, E. see Brown, K. S. Dror Farjoun, E., A transfer map for singular fibrations, 66 (1989), 132–142 Dror Farjoun, E., Farkas, H. M., Harper, J. R. and Miller, H., Alexander

Zabrodsky 1936–1986, 66 (1989), 7–8

Dror Farjoun, E., Orr, K. and Shelah, S., Bousfield localization as an algebraic

closure of groups, 66 (1989), 143–153

Droste, M., Complete embeddings of linear orderings and embeddings of lattice-ordered

groups, 56 (1986), 315–334

Droste, M. and Shelah, S., A construction of all normal subgroup lattices of 2-transitive

automorphism groups of linearly ordered sets, 51 (1985), 223–261

Dubins, L. E., On roulette when the holes are of various sizes, 11 (1972), 153–158 Dubins, L. E., Subfair casino functions are superadditive, 5 (1967), 182–184 Dubins, L. E., Hirsch, M. W. and Karush, J., Scissor congruence, 1 (1963), 239–247 Dubins, L. E., Maitra, A., Purves, R. A. and Sudderth, W. D., Measurable,

nonleavable gambling problems, 67 (1989), 257–271

Dubinsky, E., Infinite type power series subspaces of finite type power series spaces, 15

(1973), 257–281

Dubinsky, E., Infinite type power series subspaces of infinite type power series spaces, 20

(1975), 359–368

Dugas, M. and Göbel, R., Countable mixed abelian groups with very nice full-rank

subgroups, 51 (1985), 1–12

Dugas, M. and Göbel, R., Separable abelian p-groups having certain prescribed chains,

72 (1990), 289–298

Dugas, M. and Herden, G., Arbitrary torsion classes and almost free abelian groups, 44

(1983), 322–334

Dugas, M. and Irwin, J., On thickness and decomposability of abelian p-groups, 79

(1992), 153–159

Duke, R. A. and Haggard, G., The genus of subgraphs of K_{8}, 11 (1972), 452–455 Duke, R. A. and Winkler, P. M., Degree sets of k-trees: small k, 40 (1981), 296–306 Dumont, J. M., Kamae, T. and Takahashi, S., Minimal cocycles with the scaling

property and substitutions, 95 (1996), 393–410

Dumont, J. M. see Blanchard, F. Durdević, M., Generalized braided quantum groups, 98 (1997), 329–348 Duval, A. M., A combinatorial decomposition of simplicial complexes, 87 (1994), 77–87 Dvoretzky, A., On series in linear topological spaces, 1 (1963), 37–47 Dvoretzky, A., On the oscillation of the Brownian motion process, 1 (1963), 212–214 Dvoretzky, A. see Adelman, O. Dvoretzky, A. see Chow, Y. S. Dwyer, W. G., Miller, H. and Neisendorfer, J. A., Fibrewise completion and

unstable Adams spectral sequences, 66 (1989), 160–178

Dwyer, W. G., R-nilpotency in homotopy equivalences, 66 (1989), 154–159 Dwyer, W. G. see Dror, E. Dyer, J. A., Generalized Markushevich bases, 7 (1969), 51–59 Dyer, J. A., Porcelli, P. and Rosenfeld, M., Spectral characterization of two sided

ideals in B(H), 10 (1971), 26–31 Dym, H., An extremal problem in the theory of Hardy functions, 18 (1974), 391–399 Dym, H., Trace formulas for a class of Toeplitz-like operators, 27 (1977), 21–48 Dyn, N., Jackson, I. R. H., Levin, D. and Ron, A., On multivariate approximation by

integer translates of a basis function, 78 (1992), 95–130

Eckhoff, J., Intersection properties of boxes. Part I: An upper-bound theorem, 62 (1988),

283–301

Eckhoff, J., Intersection properties of boxes. Part II: Extremal families, 73 (1991),

129–149

Eckhoff, J., On a class of convex polytopes, 23 (1976), 332–336 Eda, K., A characterization of ℵ_{1}-free abelian groups and its application to the Chase

radical, 60 (1987), 22–30

Edelson, A. L. see Barnette, D. W. Edelstein, M., Farthest points of sets in uniformly convex Banach spaces, 4 (1966),

171–176

Edelstein, M., Intersections by hyperplanes, 7 (1969), 90–94 Edelstein, M. and Schwarz, B., On the length of linked curves, 23 (1976), 94–95 Edelstein, M. see Collier, J. B. Edmunds, D. E. see Chiappinelli, R. Effros, E. G., On a class of real Banach spaces, 9 (1971), 430–458 Effros, E. G. see Cunningham, F., Jr. Efrat, I., A Galois-theoretic characterization of p-adically closed fields, 91 (1995), 273–284 Efrat, I. and Haran, D., On Galois groups over Pythagorean and semi-real closed fields,

85 (1994), 57–78

Eggleston, H. G., Minimal universal covers in E^{n}, 1 (1963), 149–155 Eggleston, H. G., Sets of constant width in finite dimensional Banach spaces, 3 (1965),

163–172

Ehrenfeucht, A. and Moran, G., Size direction games over the real line. I, 14 (1973),

163–168

Eigen, S. J., On the simplicity of the full group of ergodic transformations, 40 (1981),

345–349

Eigen, S. J., Hajian, A. and Kalikow, S., Ergodic transformations and sequences of

integers, 75 (1991), 119–128

Einstein-Matthews, S. M. and Kasana, H. S., Proximate order and type of entire

functions of several complex variables, 92 (1995), 273–284

Eke, B. G., Typically real mean univalent functions of large growth, 22 (1975), 1–6 Eklof, P. C., Infinitary model theory of abelian groups, 25 (1976), 97–107 Eklof, P. C., Mekler, A. H. and Shelah, S., Almost disjoint abelian groups, 49 (1984),

34–54

Eklof, P. C., Mekler, A. H. and Shelah, S., On strongly-non-reflexive groups, 59

(1987), 283–298

Eklof, P. C., Mekler, A. H. and Shelah, S., Uniformization and the diversity of

Whitehead groups, 80 (1992), 301–321

Eklof, P. C., Mekler, A. H. and Shelah, S., Hereditarily separable groups and mono-

chromatic uniformization, 88 (1994), 213–235

El-Mosalamy, M. S., Free subgroups of small cancellation groups, 56 (1986), 345–348 Ellis, M. H., The d̄-distance between two Markov processes cannot always be attained by

a Markov joining, 24 (1976), 269–273

Ellis, M. H. and Friedman, N. A., Subsequence generators for ergodic group translations,

31 (1978), 115–121

Ellis, R. and Keynes, H. B., Bohr compactifications and a result of Følner, 12 (1972),

314–330

Ellis, R. and Perrizo, W., Unique ergodicity of flows on homogeneous spaces, 29 (1978),

276–284

Ellis, R. see Brezin, J.

Elton, J., Extremely weakly unconditionally convergent series, 40 (1981), 255–258 Enflo, P., Banach spaces which can be given an equivalent uniformly convex norm, 13

(1972), 281–288

Enflo, P., Uniform structures and square roots in topological groups, Part I., 8 (1970),

230–252

Enflo, P., Uniform structures and square roots in topological groups, Part II., 8 (1970),

253–272

Enochs, E. E., Injective and flat covers, envelopes and resolvents, 39 (1981), 189–209 Enochs, E. E. see Cheatham, T. J. Entringer, R. C. and Purdy, G. B., How often is a polygon bounded by three sides?,

43 (1982), 23–27

Epstein, B., On a difference equation arising in a learning-theory model, 4 (1966), 145–152 Epstein, B. see Blum, J. R. Erdélyi, T. see Borwein, P. Erdős, P., On cliques in graphs, 4 (1966), 233–234 Erdős, P., On extremal problems of graphs and generalized graphs, 2 (1964), 183–190 Erdős, P., On some extremal problems in graph theory, 3 (1965), 113–116 Erdős, P., On the multiplicative representation of integers, 2 (1964), 251–261 Erdős, P., On the structure of linear graphs, 1 (1963), 156–160 Erdős, P., On the sum ∑_{d|2^{n}−1} d^{−1}, 9 (1971), 43–48 Erdős, P., Some remarks on number theory, 3 (1965), 6–12 Erdős, P., Some remarks on number theory. II, 5 (1967), 57–64 Erdős, P., Frankl, P. and Füredi, Z., Families of finite sets in which no set is covered

by the union of r others, 51 (1985), 79–89

Erdős, P., Herzog, M. and Schönheim, J., An extremal problem on the set of

noncoprime divisors of a number, 8 (1970), 408–412

Erdős, P. and Netanyahu, E., A remark on polynomials and the transfinite diameter,

14 (1973), 23–25

Erdős, P., Sárközy, A. and Sós, V. T., On sum sets of Sidon sets, II, 90 (1995),

221–233

Erdős, P. and Shelah, S., Separability properties of almost-disjoint families of sets, 12

(1972), 207–214

Erdős, P. and Smith, B., Finite abelian group cohesion, 39 (1981), 177–185 Erdős, P. see Bollobás, B. Escauriaza, L., The L^{p} Dirichlet problem for small perturbations of the Laplacian, 94

(1996), 353–366

Escobedo, M. and Zuazua, E., Self-similar solutions for a convection–diffusion equation

with absorption in R^{N}, 74 (1991), 47–64

Eskin, G., Asymptotics near the boundary of spectral functions of elliptic self-adjoint

boundary problems, 22 (1975), 214–246

Evans, L. C., Nonlinear evolution equations in an arbitrary Banach space, 26 (1977), 1–42 Evans, L. C., On solving certain nonlinear partial differential equations by accretive

operator methods, 36 (1980), 225–247

Evans, L. C. see Crandall, M. G. Evans, R. J., Character sum analogues of constant term identities for root systems, 46

(1983), 189–196

Evans, S. N. and Perkins, E., Measure-valued Markov branching processes conditioned

on non-extinction, 71 (1990), 329–337

Everest, G. see Chothi, V. Evyatar, A. A., On polynomial equations, 10 (1971), 321–326 Evyatar, A. A. and Villamayor, O. E., Taylor series and divisibility, 18 (1974), 309–319 Ewald, G., On shortness exponents of families of graphs, 16 (1973), 53–61 Ewald, G. see Bokowski, J.

Ezin, J.-P., Rigoli, M. and Salavessa, I. M. C., Spherical-type hypersurfaces in a

Riemannian manifold, 71 (1990), 193–209

Ezquerro, L. M. see Ballester-Bolinches, A. Fabec, R. C., Normal ergodic actions and extensions, 40 (1981), 175–186 Fabian, M. and Troyanski, S. L., A Banach space admits a locally uniformly rotund

norm if its dual is a Vašák space, 69 (1990), 214–224

Fabian, M., Whitfield, J. H. M. and Zizler, V., Norms with locally Lipschitzian

derivatives, 44 (1983), 262–276

Facchini, A., Torsion-free covers and pure-injective envelopes over valuation domains, 52

(1985), 129–139

Faith, C., Semiperfect Prüfer rings and FPF rings, 26 (1977), 166–177 Fajuyigbe, O., An extension theorem for periodic transformations of surfaces. I, 34 (1979),

12–20

Falkowitz, M., On weakly mixing Markov processes, 14 (1973), 221–227 Faltings, G. and Jordan, B. W., Crystalline cohomology and GL(2,ℚ), 90 (1995), 1–66 Fan, K., Invariant cross-sections and invariant linear subspaces, 2 (1964), 19–26 Fan, P. and Stampfli, J. G., On the density of hyponormal operators, 45 (1983), 255–256 Farahat, J. E., On the problem of k structure, 28 (1977), 141–150 Farber, M., Minimal Seifert manifolds and the knot finiteness theorem, 66 (1989), 179–215 de Faria, E. see Coelho, Z. Farkas, H. M., Period relations for hyperelliptic Riemann surfaces, 10 (1971), 289–301 Farkas, H. M. and Kopeliovich, Y., New theta constant identities, 82 (1993), 133–140 Farkas, H. M. and Kra, I., Automorphic forms for subgroups of the modular group, 82

(1993), 87–131

Farkas, H. M. and Kra, I., Branched two-sheeted covers, 74 (1991), 169–197 Farkas, H. M. see Dror Farjoun, E. Farmer, J. D., Polynomial reflexivity in Banach spaces, 87 (1994), 257–273 Fathi, A., Le groupe des transformations de [0, 1] qui préservent la mesure de Lebesgue est

un groupe simple, 29 (1978), 302–308

Faudree, R. J., Rousseau, C. C., Schelp, R. H. and Schuster, S., Panarboreal

graphs, 35 (1980), 177–185

Feder, M. and Saphar, P., Spaces of compact operators and their dual spaces, 21 (1975),

38–49

Fefferman, C., A note on spherical summation multipliers, 15 (1973), 44–52 Fefferman, C., L^{p} bounds for pseudo-differential operators, 14 (1973), 413–417 Fein, B., Hales, A. and Schacher, M., Brauer groups are not characterized by Ulm

invariants, 55 (1986), 363–379

Fein, B. and Schacher, M., Sums of corestrictions of cyclic algebras, 96 (1996), 243–258 Feintuch, A. and Rosenthal, P., Remarks on reductive operator algebras, 15 (1973),

130–136

Feit, P., Følner sequences and Hilbert’s irreducibility theorem over Q, 69 (1990), 289–320 Feit, W., Schur indices of characters of groups related to finite sporadic simple groups, 93

(1996), 229–251

Feit, W., The computations of some Schur indices, 46 (1983), 274–300 Feit, W., The K-admissibility of 2A_{6} and 2A_{7}, 82 (1993), 141–156 Feldman, J., A generalization of a result of R. Lyons about measures on [0, 1), 81 (1993),

281–287

Feldman, J., A short proof of the Levy continuity theorem in Hilbert space, 3 (1965),

99–103

Feldman, J., New K-automorphisms and a problem of Kakutani, 24 (1976), 16–38 Feldman, J., r-entropy, equipartition, and Ornstein’s isomorphism theorem in R^{n}, 36

(1980), 321–345

Felix, Y., Halperin, S. and Thomas, J.-C., The category of a map and the grade of a module, 78 (1992), 177–196 Feller, E. H. see Boyle, A. K. Feller, E. H. see Deshpande, M. G. Felzenszwalb, B., Rings radical over subrings, 23 (1976), 156–164 Ferenczi, S., Existence d’un générateur suivant une sous-suite de densité supérieure nulle,

42 (1982), 334–342

Ferenczi, S., Measure-theoretic complexity of ergodic systems, 100 (1997), 189–207 Feres, R., Connection preserving actions of lattices in SL_{n}ℝ, 79 (1992), 1–21 Fernández, J. L., A note on entropy and inner functions, 53 (1986), 158–162 Fernández-Alcober, G. A. see Vera López, A. Ferrero, M. and Puczyłowski, E. R., Prime ideals and radicals of centred extensions

and tensor products, 94 (1996), 381–401

Fialkow, L. A., A note on norm ideals and the operator X ⟶ AX − XB, 32 (1979),

331–348

Fieldsteel, A., Factor orbit equivalence of compact group extensions, 38 (1981), 289–303 Fieldsteel, A., The relative isomorphism theorem for Bernoulli flows, 40 (1981), 197–216 Figiel, T. and Johnson, W. B., Large subspaces of l_{∞}^{n} and estimates of the Gordon–Lewis

constant, 37 (1980), 92–112

Figiel, T. and Tomczak-Jaegermann, N., Projections onto Hilbertian subspaces of

Banach spaces, 33 (1979), 155–171

Figiel, T. see Bourgain, J. Filippi, M., Variété de non-synthèse spectrale sur un groupe abelien localement compact,

3 (1965), 43–60

Filliman, P., The largest projections of regular polytopes, 64 (1988), 207–228 Finet, C., Subspaces of Asplund Banach spaces with the point continuity property, 60

(1987), 191–198

Finet, C., Uniform convexity properties of norms on a super-reflexive Banach space, 53

(1986), 81–92

Firey, W. J., A functional characterization of certain mixed volumes, 24 (1976), 274–281 Firey, W. J., Intermediate Christoffel-Minkowski problems for figures of revolution, 8

(1970), 384–390

Firey, W. J., The mixed area of a convex body and its polar reciprocal, 1 (1963), 201–202 Firey, W. J. and Grünbaum, B., Addition and decomposition of convex polytopes, 2

(1964), 91–100

Fischman, D. see Bahturin, Y. Fischman, D. see Bergen, J. Fisher, E. R. and Robinson, A., Inductive theories and their forcing companions, 12

(1972), 95–107

Fisher, E. R. see Barwise, J. Fisher, S. D., Singular semi-linear equations in L^{1}(R), 28 (1977), 129–140 Fishman, K. and Moskovitz, H., Conditions for extension of some classes of square-

summable analytic functions, 37 (1980), 273–280

Fisman, E. and Schaps, M., The combinatorics of alternative factorizations, 48 (1984),

79–95

Fitzsimmons, P. J., Construction and regularity of measure-valued Markov branching

processes, 64 (1988), 337–361

Fitzsimmons, P. J., Correction to “Construction and regularity of measure-valued Markov branching processes” [Israel J. Math, 64 (1988), no. 3, 337–361], 73 (1991), 127–127

Flatto, L., A limit theorem for random coverings of a circle, 15 (1973), 167–184 Flytzanis, E., Rigid factors of ergodic transformations, 35 (1980), 323–326 Foguel, S. R., A generalized 0-2 law, 45 (1983), 219–224 Foguel, S. R., Algebraic functions of normal operators, 6 (1968), 199–201 Foguel, S. R., Ergodic decomposition of a topological space, 7 (1969), 164–167 Foguel, S. R., Existence of a σ finite invariant measure for a Markov process on a locally

compact space, 6 (1968), 1–4

Foguel, S. R., Harris operators, 33 (1979), 281–309 Foguel, S. R., Invariant subspaces of a measure preserving transformation, 2 (1964),

198–200

Foguel, S. R., On a paper of A. Feldzamen, 1 (1963), 133–138 Foguel, S. R., On order preserving contractions, 1 (1963), 54–59 Foguel, S. R., On spectrality criterion for operators on a direct sum of Hilbert spaces, 3

(1965), 248–250

Foguel, S. R., On the “zero-two” law, 10 (1971), 275–280 Foguel, S. R., Ratio limit theorems for Markov processes, 7 (1969), 384–392 Foguel, S. R., Remarks on conservative Markov processes, 6 (1968), 381–383 Foguel, S. R., Strong convergence of the iterates of an operator, 16 (1973), 159–161 Foguel, S. R., The ergodic theorem for Markov processes, 4 (1966), 11–22 Foias, C. see Bercovici, H. Fonda, A. see Colombo, G. Fonf, V. P., Boundedly complete basic sequences, c_{0}-subspaces, and injections of Banach

spaces, 89 (1995), 173–188

Ford, T. J., The Brauer group of a curve over a strictly local discrete valuation ring, 96

(1996), 259–266

Formanek, E. and Lawrence, J., The group algebra of the infinite symmetric group, 23

(1976), 325–331

Formanek, E. and Lichtman, A. I., Ideals in group rings of free products, 31 (1978),

101–104

Forrest, A. H., K-groups associated with substitution minimal systems, 98 (1997),

101–139

Forrest, A. H., The construction of a set of recurrence which is not a set of strong

recurrence, 76 (1991), 215–228

Förster, P., Pronormal subgroups and homomorphs in finite groups, 55 (1986), 94–108 Fournier, J. J. F., Majorants and L^{p} norms, 18 (1974), 157–166 Frank, L. S., Algèbre des opérateurs aux différences finies, 13 (1972), 24–55 Frankl, P. see Erdős, P. Franklin, S. P., Compactness and semi-continuity, 3 (1965), 13–14 Fransman, A. see Amberg, B. Freedman, M. A., Product integrals of continuous resolvents: existence and nonexistence,

46 (1983), 145–160

Frehse, J., On systems of second-order variational inequalities, 15 (1973), 421–429 Freiman, G. A. and Schein, B. M., Structure of R(3, 3)-groups, 77 (1992), 17–31 Freiman, G. A. see Deshouillers, J-M. Fremlin, D. H. and Shelah, S., On partitions of the real line, 32 (1979), 299–304 Fremlin, D. H. see Dodds, P. G. Freud, R. and Pálfy, P. P., On the possible number of elements of given order in a finite

group, 93 (1996), 345–358

Frey, G., Curves with infinitely many points of fixed degree, 85 (1994), 79–83 Friberg, B., Canonical isotopies in Euclidean space, 16 (1973), 398–403 Fried, M. D., On the Sprindžuk–Weissauer approach to universal Hilbert subsets, 51

(1985), 347–363

Fried, M. D. and Friedland, S., A discriminant criteria for reducibility of a polynomial,

54 (1986), 25–32

Fried, M. D., Guralnick, R. M. and Saxl, J., Schur covers and Carlitz’s conjecture,

82 (1993), 157–225

Fried, M. D., Haran, D. and Jarden, M., Effective counting of the points of definable sets over finite fields, 85 (1994), 103–133 Fried, M. D. and Völklein, H., The absolute Galois group of a Hilbertian PRC field,

85 (1994), 85–101

Friedland, D., Irreflexive Banach spaces are imperfect, 8 (1970), 30–33 Friedland, S., A characterization of normal operators, 42 (1982), 235–240 Friedland, S., Extremal eigenvalue problems for convex sets of symmetric matrices and

operators, 15 (1973), 311–331

Friedland, S., Matrices with prescribed off-diagonal elements, 11 (1972), 184–189 Friedland, S., On an inverse problem for nonnegative and eventually nonnegative matrices,

29 (1978), 43–60

Friedland, S., On pointwise and analytic similarity of matrices, 35 (1980), 89–108 Friedland, S. see Fried, M. D. Friedlander, J., On exact coverings of the integers, 12 (1972), 299–305 Friedlander, L., On the second eigenvalue of the Dirichlet Laplacian, 79 (1992), 23–32 Friedman, A., Probabilistic methods in partial differential equations, 13 (1972), 56–64 Friedman, H., Beth’s theorem in cardinality logics, 14 (1973), 205–212 Friedman, N. A., Eventually independent sequences, 33 (1979), 310–316 Friedman, N. A., Mixing and sweeping out, 68 (1989), 365–375 Friedman, N. A. and King, J. L., Rank one lightly mixing, 73 (1991), 281–288 Friedman, N. A. see Ellis, M. H. Friedman, S. D., Uncountable admissibles II: Compactness, 40 (1981), 129–149 Friedman, Y. see Arazy, J. Frumkin, A., The irreducible characters of the Lie superalgebras of type A(n, m) and

filtrations of their Kac modules, 96 (1996), 267–279

Frumkin, A., Theorem about the conjugacy representation of S_{n}, 55 (1986), 121–128 Frumkin, A. and Yakir, A., Rank of inclusion matrices and modular representation

theory, 71 (1990), 309–320

Frumkin, A. see Adin, R. M. Fuchs, L. and Magidor, M., Butler groups of arbitrary cardinality, 84 (1993), 239–263 Fuhrmann, P. A., A functional calculus in Hilbert space based on operator valued analytic

functions, 6 (1968), 267–278

Fuhrmann, P. A., On a class of finite dimensional contractive perturbations of restricted

shifts of finite multiplicity, 16 (1973), 162–175

Fuhrmann, P. A., On generalized Hankel operators induced by sums and products, 21

(1975), 279–295

Fuhrmann, P. A., On maximal ideals in two algebras of operator valued analytic functions,

8 (1970), 23–27

Füredi, Z. see Erdős, P. Furstenberg, H. and Katznelson, Y., Idempotents in compact semigroups and Ramsey

theory, 68 (1989), 257–270

Furstenberg, H., Keynes, H. B. and Shapiro, L., Prime flows in topological dynamics,

14 (1973), 26–38

Furstenberg, H. and Kifer, Y., Random matrix products and measures on projective

spaces, 46 (1983), 12–32

Furstenberg, H. and Tzkoni, I., Spherical functions and integral geometry, 10 (1971),

327–338

Furstenberg, H. and Weiss, B., On almost 1–1 extensions, 65 (1989), 311–322 Furstenberg, H. see Bellow, A. Furuya, K., Analyticity of nonlinear semigroups, 68 (1989), 89–101 Gabbay, D. M., Decidability results in non-classical logic. III. (Systems with statability

operators), 10 (1971), 135–146

Gabbay, D. M., On decidable, finitely axiomatizable, modal and tense logics without the

finite model property. Part I, 10 (1971), 478–495

Gabbay, D. M., On decidable, finitely axiomatizable, modal and tense logics without the finite model property. Part II, 10 (1971), 496–503 Gabber, O., Affine analog of the proper base change theorem, 87 (1994), 325–335 Gábor, L. see Lükő, G. Gaboriau, D., Levitt, G. and Paulin, F., Pseudogroups of isometries of ℝ and Rips’

theorem on free actions on ℝ-trees, 87 (1994), 403–428

Gaifman, H., A generalization of Malho’s method for obtaining large cardinal numbers, 5

(1967), 188–200

Gaifman, H., Concerning measures in first order calculi, 2 (1964), 1–18 Gaifman, H., Finiteness is not a Σ_{0}-property, 19 (1974), 359–368 Gaifman, H., Global and local choice functions, 22 (1975), 257–265 Gal, S., A general search game, 12 (1972), 32–45 Galaktionov, V. A. and Levine, H. A., On critical Fujita exponents for heat equations

with nonlinear flux conditions on the boundary, 94 (1996), 125–146

Galeeva, R., Martens, M. and Tresser, C., Inducing, slopes, and conjugacy classes,

99 (1997), 123–147

Gallardo, D., Weighted weak type integral inequalities for the Hardy–Littlewood maximal

operator, 67 (1989), 95–108

Gallivan, S., On the number of disjoint increasing paths in the one-skeleton of a convex

body leading to a given exposed face, 32 (1979), 282–288

Gamelin, T. W., Wolff’s proof of the Corona theorem, 37 (1980), 113–119 Gamlen, J. L. B. and Gaudet, R. J., On subsequences of the Haar system in L_{p}

[0, 1], (1 < p < ∞), 15 (1973), 404–413

García Hernández, J. L. and Gómez Pardo, J. L., Self-injective and PF endomorphism

rings, 58 (1987), 324–350

Garcia-Cuerva, J., Macias, R. and Torrea, J. L., The Hardy–Littlewood property of

Banach lattices, 83 (1993), 177–201

Gårding, L., Local hyperbolicity, 13 (1972), 65–81 Garling, D. J. H., Convexity, smoothness and martingale inequalities, 29 (1978), 189–198 Garling, D. J. H. and Gordon, Y., Relations between some constants associated with

finite dimensional Banach spaces, 9 (1971), 346–361

Garling, D. J. H. and Tomczak-Jaegermann, N., The cotype and uniform convexity

of unitary ideals, 45 (1983), 175–197

Gauchman, H., On locally symmetric vector fields on Riemannian manifolds, 33 (1979),

37–51

Gaudet, R. J. see Gamlen, J. L. B. Gay, R. see Berenstein, C. A. Gelbart, S., Howe, R. E. and Piatetski-Shapiro, I. I., Uniqueness and existence of

Whittaker models for the metaplectic group, 34 (1979), 21–37

Gelbart, S. and Piatetski-Shapiro, I. I., Some remarks on metaplectic cusp forms and

the correspondences of Shimura and Waldspurger, 44 (1983), 97–126

Gelbart, S., Rogawski, J. and Soudry, D., On periods of cusp forms and algebraic

cycles for U(3), 83 (1993), 213–252

Genel, A. and Lindenstrauss, J., An example concerning fixed points, 22 (1975), 81–86 Gerber, M., A zero-entropy mixing transformation whose product with itself is loosely

Bernoulli, 38 (1981), 1–22

Gerber, M., Factor orbit equivalence of compact group extensions and classification of

finite extensions of ergodic automorphisms, 57 (1987), 28–48

Gerritzen, L., Equations defining the periods of totally degenerate curves, 77 (1992),

187–210

Geyer, W.-D., Algebraic function fields of one variable over finite fields are stable, 68

(1989), 102–108

Geyer, W.-D., Galois groups of intersections of local fields, 30 (1978), 382–396 Geyer, W.-D., On the irreducibility of sums of rational functions with separated variables,

85 (1994), 135–168 Geyer, W.-D. and Jarden, M., On the normalizer of finitely generated subgroups of

absolute Galois groups of uncountable Hilbertian fields of characteristic 0, 63 (1988), 323–334

Geyer, W.-D. and Jarden, M., Torsion points of elliptic curves over large algebraic

extensions of finitely generated fields, 31 (1978), 257–297

Giambruno, A. and Valenti, A., Central polynomials and matrix invariants, 96 (1996),

281–297

Giambruno, A. see Berele, A. Giesy, D. P., B-convexity and reflexivity, 15 (1973), 430–436 Gil-Medrano, O. and Michor, P. W., Geodesics on spaces of almost hermitian structures,

88 (1994), 319–332

Gilat, D., On the ratio of the expected maximum of a martingale and the L_{p}-norm of its

last term, 63 (1988), 270–280

Gilat, D. and Weiss, E.-A., Jr., On roulette which allows stakes on infinitely many

holes, 24 (1976), 282–285

Gillet, H. and Soulé, C., On the number of lattice points in convex symmetric bodies

and their duals, 74 (1991), 347–357

Gillis, J. see Debbi, D. Giménez, F., Comparison theorems for the volume of a complex submanifold of a Kaehler

manifold, 71 (1990), 239–255

Giménez, F. and Miquel, V., On nonimmersibility of compact hypersurfaces into a ball

of a simply connected space form, 94 (1996), 21–27

Ginzburg, A. and Tamari, D., Representation of generalized groups by families of binary

relations, 7 (1969), 33–45

Ginzburg, A. see Tamari, D. Ginzburg, D., A Rankin–Selberg integral for the adjoint L-function of Sp_{4}, 95 (1996),

301–339

Ginzburg, D., On the symmetric fourth power L-function of GL_{2}, 92 (1995), 157–184 Ginzburg, D., Rallis, S. and Soudry, D., On the automorphic theta representation for

simply laced groups, 100 (1997), 61–116

Girardi, M. and Johnson, W. B., Universal non-completely-continuous operators, 99

(1997), 207–219

Girbau, J., Vanishing cohomology theorems and stability of complex analytic foliations,

40 (1981), 235–254

Girela, D., Integral means of univalent functions with restricted Hayman index, 65 (1989),

44–58

Gitik, M., All uncountable cardinals can be singular, 35 (1980), 61–88 Gitik, M., Changing cofinalities and the nonstationary ideal, 56 (1986), 280–314 Gitik, M., Some results on the nonstationary ideal, 92 (1995), 61–112 Gitik, M., Some results on the nonstationary ideal II, 99 (1997), 175–188 Gitik, M., The nonstationary ideal on ℵ_{2}, 48 (1984), 257–288 Gitik, M. and Shelah, S., Forcings with ideals and simple forcing notions, 68 (1989),

129–160

Gitik, M. and Shelah, S., On the 𝕀-condition, 48 (1984), 148–158 Gitik, M. and Witzany, J., Consistency strength of the axiom of full reflection at large

cardinals, 93 (1996), 113–124

Gitik, R. and Rips, E., A necessary condition for A ∗_{a=b} B to be LERF, 73 (1991), 123–125

Glasner, E., Minimal nil-transformations of class two, 81 (1993), 31–51 Glasner, E., Host, B. and Rudolph, D. J., Simple systems and their higher order

self-joinings, 78 (1992), 131–142

Glasner, E. see Aaronson, J. Glasner, M. and Nakai, M., Singular points on compactifications of a Riemann surface,

55 (1986), 15–32 Glasner, M. and Nakai, M., The roles of sets of nondensity points, 36 (1980), 1–12 Glasner, S., A metric minimal flow whose enveloping semigroup contains finitely many

minimal ideals is PI, 22 (1975), 87–92

Glasner, S., Almost periodic sets and measures on the torus, 32 (1979), 161–172 Glasner, S., Proximal flows of Lie groups, 45 (1983), 97–99 Glasner, S., Quasi-factors in ergodic theory, 45 (1983), 198–208 Glasner, S. and Weiss, B., Interpolation sets for subalgebras of l^{∞}(Z), 44 (1983),

345–360

Glasner, S. and Weiss, B., Minimal transformations with no common factor need not be

disjoint, 45 (1983), 1–8

Glasner, S. and Weiss, B., On the construction of minimal skew products, 34 (1979),

321–336

Glasner, S. see Conze, J. P. Glass, A. M. W. and Pierce, K. R., Existentially complete lattice-ordered groups, 36

(1980), 257–272

Gleit, A., On the existence of simplex spaces, 9 (1971), 199–209 Glicksberg, I., A remark on almost periodic transition operators, 3 (1965), 71–74 Glicksberg, I., A remark on eigenvalues of certain positive operators, 7 (1969), 147–150 Glicksberg, I., Some remarks on ideals in function algebras, 8 (1970), 413–418 Glicksberg, I., Two remarks on Fourier–Stieltjes transforms, 32 (1979), 1–13 Gluskin, E. D., Meyer, M. and Pajor, A., Zeros of analytic functions and norms of

inverse matrices, 87 (1994), 225–242

Gluskin, E. D., Tomczak-Jaegermann, N. and Tzafriri, L., Subspaces of l_{p}^{N} of small

codimension, 79 (1992), 173–192

Göbel, R. see Dugas, M. Godefroy, G., On Riesz subsets of abelian discrete groups, 61 (1988), 301–331 Godefroy, G., Points de Namioka. Espaces normants. Applications à la théorie isométrique

de la dualité, 38 (1981), 209–220

Godefroy, G., Symétries isométriques. Applications à la dualité des espaces réticulés, 44

(1983), 61–74

Godefroy, G. and Louveau, A., Axioms of determinacy and biorthogonal systems, 67

(1989), 109–116

Godefroy, G. and Talagrand, M., Espaces de Banach représentables, 41 (1982),

321–330

Godefroy, G. see Deville, R. Goel, V. K., Jain, S. K. and Singh, S., Semiprime rings with finite length w.r.t. an

idempotent kernel functor, 28 (1977), 110–112

Gohberg, I. and Rodman, L., On spectral analysis of non-monic matrix and operator

polynomials, I. Reduction to monic polynomials, 30 (1978), 133–151

Gohberg, I. and Rodman, L., On spectral analysis of non-monic matrix and operator

polynomials, II. Dependence on the finite spectral data, 30 (1978), 321–334

Golan, J. S., A Krull-like dimension for noncommutative rings, 19 (1974), 297–304 Golan, J. S., Characterization of rings using quasiprojective modules, 8 (1970), 34–38 Golan, J. S., On the endomorphism ring of a module Noetherian with respect to a torsion

theory, 45 (1983), 257–264

Golan, J. S. and Teply, M. L., Torsion-free covers, 15 (1973), 237–256 Goldberger, J., Lind, D. and Smorodinsky, M., The entropies of renewal systems, 75

(1991), 49–64

Goldfeld, M., A large sieve for a class of non-abelian L-functions, 14 (1973), 39–49 Goldie, A., Some recent developments in ring theory, 19 (1974), 153–168 Goldrei, D. C., Macintyre, A. and Simmons, H., The forcing companions of number

theories, 14 (1973), 317–337

Goldring, N., Projecting precipitousness, 74 (1991), 13–31 Gol’dshtein, V., Gurov, L. and Romanov, A., Homeomorphisms that induce monomor-

phisms of Sobolev spaces, 91 (1995), 31–60

Goldstein, L. J., On a formula of Hecke, 17 (1974), 283–301 Goldstein, M. S. and Grabarnik, G. Ya., Almost sure convergence theorems in von

Neumann algebras, 76 (1991), 161–182

Goldstein, S., Entropy increase in dynamical systems, 38 (1981), 241–256 Goldstern, M. and Judah, H., Iteration of Souslin forcing, projective measurability and

the Borel conjecture, 78 (1992), 335–362

v. Golitschek, M. see Bak, J. Gómez, F., De Rham homology for networks of manifolds, 54 (1986), 110–128 Gómez Pardo, J. L., The rational Loewy series and nilpotent ideals of endomorphism

rings, 60 (1987), 315–332

Gómez Pardo, J. L. see García Hernández, J. L. Gómez-Mont, C., On local Torelli for a class of extremal varieties, 71 (1990), 353–377 Gonçalves, J. Z. and Mandel, A., Are there free groups in division rings?, 53 (1986),

69–80

Gonshor, H., Projective covers as subquotients of enlargements, 14 (1973), 257–261 González, M. see Castillo, J. M. F. Gonzalo, J. and Varela, F., Characterization of the projective space Pr_{3} by a contact

form, 53 (1986), 346–354

Gonzalo, R. and Jaramillo, J. A., Smoothness and estimates of sequences in Banach

spaces, 89 (1995), 321–341

Goodaire, E. G. and Parmenter, M. M., Units in alternative loop rings, 53 (1986),

209–216

Goodey, P. R., Hamiltonian circuits in polytopes with even sided faces, 22 (1975), 52–56 Goodey, P. R., Mean square inequalities for chords of convex sets, 42 (1982), 132–150 Goodey, P. R., Some upper bounds for the diameters of convex polytopes, 11 (1972),

380–385

Goodman, J. E. and Pach, J., Cell decomposition of polytopes by bending, 64 (1988),

129–138

Goodman, V. see Bennett, G. Goodwyn, L. W. see Adler, R. L. Gootman, E. C. and Lazar, A. J., Crossed products of type I AF algebras by abelian

groups, 56 (1986), 267–279

Góra, P. and Boyarsky, A., Absolutely continuous invariant measures for piecewise

expanding C^{2} transformation in R^{N}, 67 (1989), 272–286

Gordon, Y., A note on the gl constant of E ˇ⊗F, 39 (1981), 141–144 Gordon, Y., Asymmetry and projection constants of Banach spaces, 14 (1973), 50–62 Gordon, Y., On p-absolutely summing constants of Banach spaces, 7 (1969), 151–163 Gordon, Y., On the distance coefficient between isomorphic function spaces, 8 (1970),

391–397

Gordon, Y., On the projection and Macphail constants of l_{n}^{p} spaces, 6 (1968), 295–302 Gordon, Y., Properties of generalised juxtapolynomials, 4 (1966), 177–188 Gordon, Y., Some inequalities for Gaussian processes and applications, 50 (1985), 265–289 Gordon, Y. and Lewis, D. R., Isometries of diagonally symmetric Banach spaces, 28

(1977), 45–67

Gordon, Y., Lewis, D. R. and Retherford, J. R., Banach ideals of operators with applications to the finite dimensional structure of Banach spaces, 13 (1972), 348–360

Gordon, Y. and Reisner, S., Some geometrical properties of Banach spaces of polynomials,

42 (1982), 99–116

Gordon, Y. and Witsenhausen, H. S., On extensions of the Gale-Berlekamp switching

problem and constants of l_{p}-spaces, 11 (1972), 216–229

Gordon, Y. see Garling, D. J. H.

Gorelik, E., The uniform nonequivalence of L_{p} and l_{p}, 87 (1994), 1–8 Gorenstein, D., Finite simple groups and their classification, 19 (1974), 5–66 Gorenstein, D. and Lyons, R., On Aschbacher’s local C(G; T) theorem, 82 (1993),

227–279

Gotsman, C. see Bollobás, B. Gottlieb, D. H., Splitting off tori and the evaluation subgroup of the fundamental group,

66 (1989), 216–222

Gould, R. J. see Chartrand, G. Goullet de Rugy, A., Espaces de fonctions pondérables, 12 (1972), 147–160 Goupil, A. see Barcelo, H. Gowers, W. T., A finite-dimensional normed space with two non-equivalent symmetric

bases, 87 (1994), 143–151

Gowers, W. T., Symmetric block bases in finite-dimensional normed spaces, 68 (1989),

193–219

Gowers, W. T., Symmetric block bases of sequences with large average growth, 69 (1990),

129–151

Grabarnik, G. Ya. and Hrushovski, E., Singular compactness and the Neveu decom-

position, 89 (1995), 135–139

Grabarnik, G. Ya. and Katz, A., Ergodic type theorems for finite von Neumann algebras,

90 (1995), 403–422

Grabarnik, G. Ya. see Goldstein, M. S. Granirer, E. E., On amenable groups of automorphisms on Von Neumann algebras, 69

(1990), 321–336

Grant, D. D., On linear forms whose coefficients are in arithmetic progression, 15 (1973),

204–209

Gregory, M. B., p-Helson sets, 1 < p < 2, 12 (1972), 356–368 Greiner, G. and van Neerven, J. M. A. M., Adjoints of semigroups acting on vector-

valued function spaces, 77 (1992), 305–333

Grekas, S., Remarks on the strong lifting property for products, 58 (1987), 198–204 Grekas, S., Structural properties of compact groups with measure-theoretic applications,

87 (1994), 89–95

Grimmer, R. and Zeman, M., Nonlinear Volterra integrodifferential equations in a Banach

space, 42 (1982), 162–176

Gripenberg, G., An abstract nonlinear Volterra equation, 34 (1979), 198–212 Gritzmann, P., Upper and lower bounds of the valence-functional, 43 (1982), 237–243 Gritzmann, P. see Betke, U. Grobler, J. J., The zero–two law in Banach lattice algebras, 64 (1988), 32–38 Groemer, H., Stability theorems for projections of convex sets, 60 (1987), 177–190 Groemer, H. see Chakerian, G. D. Groh, U., Uniformly ergodic maps on C^{∗}-algebras, 47 (1984), 227–235 Grønbæk, N., Johnson, B. E. and Willis, G. A., Amenability of Banach algebras of

compact operators, 87 (1994), 289–324

Grossberg, R. and Shelah, S., On universal locally finite groups, 44 (1983), 289–302 Grossman, R. and Larson, R. G., Hopf-algebraic structure of combinatorial objects and

differential operators, 72 (1990), 109–117

Grosswald, E., On some generalizations of theorems by Landau and Pólya, 3 (1965),

211–220

Groszek, M. and Shore, R. A., Initial segments of the degrees of constructibility, 63

(1988), 149–177

Groves, J. R. J., Regular groups in varieties of p-groups, 19 (1974), 382–388 Groves, J. R. J., Regular p-groups and words giving rise to commutative group operations,

24 (1976), 73–77

Grubb, G., On coerciveness and semiboundedness of general boundary problems, 10 (1971), 32–95 Gruber, P. M., Isometries of the space of convex bodies contained in a Euclidean ball, 42

(1982), 277–283

Grünbaum, B., Acyclic colorings of planar graphs, 14 (1973), 390–408 Grünbaum, B., Some analogues of Eberhard’s theorem on convex polytopes, 6 (1968),

398–411

Grünbaum, B., Strictly antipodal sets, 1 (1963), 5–10 Grünbaum, B., Unambiguous polyhedral graphs, 1 (1963), 235–238 Grünbaum, B. and Kelly, L. M., Metrically homogeneous sets, 6 (1968), 183–197 Grünbaum, B. and Kelly, L. M., “Metrically homogeneous sets”: Corrigendum, 8 (1970),

93–95

Grünbaum, B. see Firey, W. J. Grunewald, F., Mennicke, J. and Vaserstein, L., On the groups SL_{2}(ℤ[x]) and

SL_{2}(k[x, y]), 86 (1994), 157–193

Grzeszczuk, P., On derivations of finite index chain conditions, dimensions and radicals,

58 (1987), 54–66

Grzeszczuk, P. and Matczuk, J., Goldie conditions for constants of algebraic derivations

of semiprime algebras, 83 (1993), 329–342

Grzeszczuk, P. see Bergen, J. Guay, M. D. and Kay, D. C., On sets having finitely many points of local nonconvexity

and property P_{m}, 10 (1971), 196–209

Guay, M. D. see Kay, D. C. Guerre, S., Sur les suites presque echangeables dans L^{q}, 1 ≤ q < 2, 56 (1986), 361–380 Guerre, S., Types et suites symétriques dans L^{p}, 1 ≤ p < +∞, p ≠ 2, 53 (1986), 191–208 Guerre, S. and Lapresté, J. T., Quelques propriétés des espaces de Banach stables, 39

(1981), 247–254

Guggenheimer, H., A formula of Furstenberg-Tzkoni type, 14 (1973), 281–282 Guggenheimer, H., Correction to: “Polar reciprocal convex bodies” (Israel J. Math. 14

(1973), no. 3, 309–316), 29 (1978), 312–312

Guggenheimer, H., Finite sets on curves and surfaces, 3 (1965), 104–112 Guggenheimer, H., Homogeneous linear differential equations with only periodic solutions,

9 (1971), 49–52

Guggenheimer, H., Polar reciprocal convex bodies, 14 (1973), 309–316 Guivarc’h, Y. and Raugi, A., Propriétés de contraction d’un semi-groupe de matrices

inversibles. Coefficients de Liapunoff d’un produit de matrices aléatoires indépendantes, 65 (1989), 165–196

Guralnick, R. M. and Robinson, G. R., On extensions of the Baer–Suzuki theorem, 82

(1993), 281–297

Guralnick, R. M. see Fried, M. D. Gurevič, R., Löwenheim–Skolem problem for functors, 42 (1982), 273–276 Gurevich, D. see Donin, J. Gurevich, Y., Monadic theory of order and topology, 1, 27 (1977), 299–319 Gurevich, Y., Monadic theory of order and topology II, 34 (1979), 45–71 Gurevich, Y. and Shelah, S., The monadic theory and the “next world”, 49 (1984),

55–68

Gurov, L. see Gol’dshtein, V. Gustafson, K. E. and Rejtö, P. A., Some essentially self-adjoint Dirac operators with

spherically symmetric potentials, 14 (1973), 63–75

Gutfreund, H. see Blatt, J. M. Gutman, S. and Martin, R. H., Jr., The porous medium equation with nonlinear

absorption and moving boundaries, 54 (1986), 81–109

Haagerup, U., A new upper bound for the complex Grothendieck constant, 60 (1987),

199–224

Haddad, G., Monotone trajectories of differential inclusions and functional differential inclusions with memory, 39 (1981), 83–100 Haddad, G. see Baillon, J.-B. Haddad, G. see Cornet, B. Hadwiger, H., Zur axiomatischen Charakterisierung des Steinerpunktes konvexer Körper,

7 (1969), 168–176

Hadwiger, H., Zur axiomatischen Charakterisierung des Steinerpunktes konvexer Körper;

Berichtigung und Nachtrag, 9 (1971), 466–472

Haggard, G. see Duke, R. A. Häggström, O., On phase transitions for subshifts of finite type, 94 (1996), 319–352 Hagler, J. and Johnson, W. B., On Banach spaces whose dual balls are not weak*

sequentially compact, 28 (1977), 325–330

Hahn, F., Discrete real time flows with quasi-discrete spectra and algebras generated by

exp q(t), 16 (1973), 20–37

Haile, D. E., Knus, M-A., Rost, M. and Tignol, J-P., Algebras of odd degree with

involution, trace forms and dihedral extensions, 96 (1996), 299–340

Hajian, A., Ito, Y. and Kakutani, S., Orbits, sections, and induced transformations,

18 (1974), 97–115

Hajian, A. see Eigen, S. J. Hajnal, A., Embedding finite graphs into graphs colored with infinitely many colors, 73

(1991), 309–319

Hakimi, S. L. and Schmeichel, E. F., Bounds on the number of cycles of length three in

a planar graph, 41 (1982), 161–180

Hales, A. see Fein, B. Halfant, M., Analytic properties of Rényi’s invariant density, 27 (1977), 1–20 Halperin, S. see Felix, Y. Hamachi, T., Suborbits and group extensions of flows, 100 (1997), 249–283 Hanani, H., Netanyahu, E. and Reichaw-Reichbach, M., The sphere in the image, 1

(1963), 188–195

Hanani, H. and Schönheim, J., Addendum: “On Steiner systems”, Israel Journal of

Mathematics 2 (2) (1964), 4 (1966), 144–144

Hanani, H. and Schönheim, J., On Steiner systems, 2 (1964), 139–142 Handel, M., Kitchens, B. and Rudolph, D. J., Metrics and entropy for non-compact

spaces, 91 (1995), 253–271

Handelman, D., Space-time boundaries for random walks obtained from diffuse measures,

86 (1994), 107–156

Handelman, D. see Boyle, M. Hansen, A. B. and Sternfeld, Y., On the characterization of the dimension of a compact

metric space K by the representing matrices of C(K), 22 (1975), 148–167

Har-El, J. and Kaniel, S., Linear programming method for rational approximation, 16

(1973), 343–349

Haran, D., Non-free torsion-free profinite groups with open free subgroups, 50 (1985),

350–352

Haran, D. and Lauwers, L., Galois stratification over e-fold ordered Frobenius fields, 85

(1994), 169–197

Haran, D. and Lubotzky, A., Embedding covers and the theory of Frobenius fields, 41

(1982), 181–202

Haran, D. and Völklein, H., Galois groups over complete valued fields, 93 (1996), 9–27 Haran, D. see Efrat, I. Haran, D. see Fried, M. D. Harase, T., Algebraic elements in formal power series rings, 63 (1988), 281–288 Harase, T., Algebraic elements in formal power series rings II, 67 (1989), 62–66 Haraux, A. see Brezis, H. Harboure, E., Macías, R. A., Segovia, C. and Torrea, J. L., Some estimates for

maximal functions on Köthe function spaces, 90 (1995), 349–371

Hare, D. E. G. see Deville, R.

Harel, D., Hamiltonian paths in infinite graphs, 76 (1991), 317–336 Harmelin, R., Aharonov invariants and univalent functions, 43 (1982), 244–254 Harmelin, R., Generalized Grunsky coefficients and inequalities, 57 (1987), 347–364 Harmelin, R., Hyperbolic metric, curvature of geodesics and hyperbolic discs in hyperbolic

plane domains, 70 (1990), 111–128

Harmelin, R., Locally convex functions and the Schwarzian derivative, 67 (1989), 367–379 Harmelin, R. and Minda, D., Quasi-invariant domain constants, 77 (1992), 115–127 Harnik, V., A two cardinal theorem for sets of formulas in a stable theory, 21 (1975), 7–23 Harnik, V. and Makkai, M., New axiomatizations for logics with generalized quantifiers,

32 (1979), 257–281

Harnik, V. and Ressayre, J.-P., Prime extensions and categoricity in power, 10 (1971),

172–185

de la Harpe, P., Robertson, A. G. and Valette, A., On the spectrum of the sum of

generators for a finitely generated group, 81 (1993), 65–96

Harper, J. R., Co-H-maps to spheres, 66 (1989), 223–237 Harper, J. R., The work of Alexander Zabrodsky 1936–1986, 66 (1989), 9–22 Harper, J. R. see Dror Farjoun, E. Harrington, L. see Shelah, S. Harris, L. A., Analytic invariants and the Schwarz–Pick inequality, 34 (1979), 177–197 Hart, B. see Shelah, S. Hart, S. see Aumann, R. J. Hartley, B. and Zalesskiĭ, A. E., On simple periodic linear groups—dense subgroups,

permutation representations, and induced modules, 82 (1993), 299–327

Harvey, C., On the Cauchy problem for hyperbolic differential equations with multiple

characteristics, 7 (1969), 217–226

Hass, J. and Scott, P., Intersections of curves on surfaces, 51 (1985), 90–120 Hawkins, J., Non-ITPFI diffeomorphisms, 42 (1982), 117–131 Haxell, P. E. and Kohayakawa, Y., The size-Ramsey number of trees, 89 (1995),

261–274

Haydn, N. T. A., The product structure of finitely presented dynamical systems, 73

(1991), 343–356

Haydon, R., A non-reflexive Grothendieck space that does not contain l_{∞}, 40 (1981),

65–73

Haydon, R., On Banach spaces which contain l^{1}(τ) and types of measures on compact

spaces, 28 (1977), 313–324

Haydon, R., On dual L^{1}-spaces and injective bidual Banach spaces, 31 (1978), 142–152 Haydon, R., Odell, E. and Sternfeld, Y., A fixed point theorem for a class of

star-shaped sets in c_{0}, 38 (1981), 75–81

Haydon, R. see Carothers, N. L. Hechler, S. H., Classifying almost-disjoint families with applications to βN − N, 10

(1971), 413–432

Hechler, S. H., Directed graphs over topological spaces: Some set theoretical aspects, 11

(1972), 231–248

Hechler, S. H., Exponents of some N-compact spaces, 15 (1973), 384–395 Hechler, S. H., Large superuniversal metric spaces, 14 (1973), 115–148 Hechler, S. H., On monads in saturated enlargements, 12 (1972), 49–50 Hechler, S. H. and Kainen, P. C., On the immersion of digraphs in cubes, 18 (1974),

221–233

Hector, G. see Alcalde Cuesta, F. Heilper, A., The zeros of functions in Nevanlinna’s area class, 34 (1979), 1–11 Heinrich, K. see Alspach, B. Hellekalek, P. and Larcher, G., On the ergodicity of a class of skew products, 54

(1986), 301–306

Henle, J. M., γ-Ramsey and γ-ineffable cardinals, 30 (1978), 85–98 Henson, C. W., Nonstandard hulls of Banach spaces, 25 (1976), 108–144 Henson, C. W., When do two Banach spaces have isometrically isomorphic nonstandard

hulls?, 22 (1975), 57–67

Herden, G. see Dugas, M. Hering, F., A generalization of the arithmetic-geometric mean inequality and an application

to finite sequences of zeros and ones, 11 (1972), 14–30

Hering, F., Nested bipartite graphs, 9 (1971), 403–417 Hernández, F. L. and Peirats, V., Orlicz function spaces without complemented copies

of l^{p}, 56 (1986), 355–360

Hernández, F. L. and Rodríguez-Salinas, B., Lattice-embedding L^{p} into Orlicz spaces,

90 (1995), 167–188

Hernández, F. L. and Rodríguez-Salinas, B., On l^{p}-complemented copies in Orlicz

spaces, 62 (1988), 37–55

Hernández, F. L. and Rodríguez-Salinas, B., On l^{p}-complemented copies in Orlicz

spaces II, 68 (1989), 27–55

Herod, J. V. and McKelvey, R. W., A Hille–Yosida theory for evolutions, 36 (1980),

13–40

Herrero, M. A. and Velázquez, J. J. L., Plane structures in thermal runaway, 81

(1993), 321–341

Hershkowitz, D. and Schneider, H., Matrices with a sequence of accretive powers, 55

(1986), 327–344

Hershkowitz, D. and Schneider, H., On the existence of matrices with prescribed height

and level characteristics, 75 (1991), 105–117

Herstein, I. N., Invariant subrings of a certain kind, 26 (1977), 205–208 Herstein, I. N., Multiplicative commutators in division rings, 31 (1978), 180–188 Herstein, I. N. and Small, L. W., The intersection of the powers of the radical in

Noetherian P.I. rings, 16 (1973), 176–180

Herzog, F. see Stewart, B. M. Herzog, M., A characterization of the simple groups PSL(2, p), p > 3, 5 (1967), 79–85 Herzog, M., Groups with strongly self-centralizing 3-centralizers, 9 (1971), 507–510 Herzog, M., On 2-Sylow intersections, 11 (1972), 326–327 Herzog, M., On finite groups with cyclic Sylow subgroups for all odd primes, 6 (1968),

206–216

Herzog, M., On simple groups with a cyclic maximal 2-Sylow intersection, 15 (1973),

350–355

Herzog, M., On simple groups with a quaternion maximal 2-Sylow intersection, 19 (1974),

225–227

Herzog, M., Longobardi, P. and Maj, M., On a combinatorial problem in group theory,

82 (1993), 329–340

Herzog, M. and Reid, K. B., Permutation groups generated by cycles of fixed length, 26

(1977), 221–231

Herzog, M. see Erdős, P. Hess, P. see Alikakos, N. D. Hess, P. see Deuel, J. Hill, P., Pride, S. J. and Vella, A. D., On the T (q)-conditions of small cancellation

theory, 52 (1985), 293–304

Hill, W. M., Frattini subgroups and supernilpotent groups, 26 (1977), 68–74 Hill, W. M., Frattini subgroups of p-closed and p-supersolvable groups, 19 (1974), 208–211 Hill, W. M. and Parker, D. B., The nilpotence class of the Frattini subgroup, 15 (1973),

211–215

Hiller, H. L. and Shelah, S., Singular cohomology in L, 26 (1977), 313–319 Hillman, J. A., On 4-manifolds homotopy equivalent to circle bundles over 3-manifolds,

75 (1991), 277–287

Hillman, J. A., On L^{2}-homology and asphericity, 99 (1997), 271–283 Hilton, P., On the genus of nilpotent groups and spaces, 54 (1986), 1–13 Hinich, V., On Brieskorn’s theorem, 76 (1991), 153–160 Hinich, V., On the singularities of nilpotent orbits, 73 (1991), 297–308 Hinman, P. G., Finite termination games with tie, 12 (1972), 17–22 Hirsberg, B., M-ideals in complex function spaces and algebras, 12 (1972), 133–146 Hirsch, M. W. see Dubins, L. E. Hirschfeld, J., Models of arithmetic and recursive functions, 20 (1975), 111–126 Hirschfeld, Y., Non standard analysis and the compactification of groups, 25 (1976),

145–153

Hirschman, Jr., I. I. and Katznelson, Y., Sets of uniqueness and multiplicity for l^{p,α},

3 (1965), 221–231

Hitczenko, P., Domination inequality for martingale transforms of a Rademacher sequence,

84 (1993), 161–178

Ho, A., The Radon–Nikodým property and weighted trees in Banach spaces, 32 (1979),

59–66

Hofbauer, F., Generic properties of invariant measures for simple piecewise monotonic

transformations, 59 (1987), 64–80

Hofbauer, F., On intrinsic ergodicity of piecewise monotonic transformations with positive

entropy, 34 (1979), 213–237

Hofbauer, F., On intrinsic ergodicity of piecewise monotonic transformations with positive

entropy II, 38 (1981), 107–115

Hoffman, A. J., On eigenvalues of symmetric (+1, −1) matrices, 17 (1974), 69–75 Holgate, P., Recurrence of sums of multiple Markov sequences, 4 (1966), 208–212 Holland, S. S., Jr., Baer ordered ∗-fields of the first kind, 57 (1987), 365–374 Holub, J. R., On generalized shift bases for the Wiener disc algebra, 52 (1985), 231–238 Hopkins, M. J., Nilpotence and finite H-spaces, 66 (1989), 238–246 Horiuchi, R., A description of the Weierstrass gap sequence by means of the Riemann

theta function, 53 (1986), 129–136

Hörmander, L., Lower bounds at infinity for solutions of differential equations with

constant coefficients, 16 (1973), 103–116

Hörmander, L., On the singularities of solutions of partial differential equations with

constant coefficients, 13 (1972), 82–105

Horowitz, C., Coefficients of nonvanishing functions in H^{∞}, 30 (1978), 285–291 Horowitz, C., Iterated logarithms of entire functions, 29 (1978), 31–42 Horowitz, J., The Hausdorff dimension of the sample path of a subordinator, 6 (1968),

176–182

Horowitz, S., L_{∞}-limit theorems for Markov processes, 7 (1969), 60–62 Horowitz, S., Markov processes on a locally compact space, 7 (1969), 311–324 Horowitz, S., On σ-finite invariant measures for Markov processes, 6 (1968), 338–345 Horowitz, S., Pointwise convergence of the iterates of a Harris-recurrent Markov operator,

33 (1979), 177–180

Horowitz, S., Some limit theorems for Markov processes, 6 (1968), 107–118 Host, B., Mixing of all orders and pairwise independent joinings of systems with singular

spectrum, 76 (1991), 289–298

Host, B., Nombres normaux, entropie, translations, 91 (1995), 419–428 Host, B. see Glasner, E. Howe, R. E. and Rosenberg, J., The unitary representation theory of GL(n) of an

infinite discrete field, 67 (1989), 67–81

Howe, R. E. see Gelbart, S. Howie, J. see Brodsky, S. D. Hrushovski, E., A non-PAC field whose maximal purely inseparable extension is PAC, 85

(1994), 199–202

Hrushovski, E., Strongly minimal expansions of algebraically closed fields, 79 (1992), 129–151 Hrushovski, E. and Pillay, A., Groups definable in local fields and pseudo-finite fields,

85 (1994), 203–262

Hrushovski, E. and Shelah, S., Stability and omitting types, 74 (1991), 289–321 Hrushovski, E. see Grabarnik, G. Ya. Huang, Z. see deLaubenfels, R. Hubbuck, J. R. and Kane, R. M., The homotopy types of compact Lie groups, 51 (1985),

20–26

Huber Dyson, V., On the decision problem for theories of finite models, 2 (1964), 55–70 Huberich, M., Non-regular ultrafilters, 87 (1994), 275–288 Hudson, R. H., On the first occurrence of certain patterns of quadratic residues and

non-residues, 44 (1983), 23–32

Hudson, R. H. and Williams, K. S., A new criterion for 7 to be a fourth power (mod p),

38 (1981), 221–230

Hustad, O., A note on complex P_{1} spaces, 16 (1973), 117–119 Hutchinson, J. E., Model theory via set theory, 24 (1976), 286–304 Hutchinson, J. J. and Teply, M. L., A module as a torsion-free cover, 46 (1983),

305–312

Iacob, F. see Pavel, N. H. Idowu, E. A., p-saturated formations, 30 (1978), 307–312 Ihara, Y., On the stable derivation algebra associated with some braid groups, 80 (1992),

135–153

Ilani, I., Counting finite index subgroups and the P. Hall enumeration principle, 68 (1989),

18–26

Impens, C., Local microcontinuity of nonstandard polynomials, 59 (1987), 81–97 Imrich, W., On products of graphs and regular groups, 11 (1972), 258–264 Iranzo, M. J., Navarro, G. and Pérez Monasor, F., A conjecture on the number of

conjugacy classes in a p-solvable group, 93 (1996), 185–188

Iranzo, M. J. and Pérez Monasor, F., Existence of N -injectors in a not central normal

Fitting class, 48 (1984), 123–128

Iranzo, M. J. and Pérez Monasor, F., Fitting classes F such that all finite groups have

F-injectors, 56 (1986), 97–101

Irving, R. S., Some more primitive group rings, 37 (1980), 331–350 Irwin, J. see Dugas, M. Isaac, R., Theorems for conditional expectations, with applications to Markov processes,

16 (1973), 362–374

Isaacs, I. M., Solvable groups contain large centralizers, 55 (1986), 58–64 Isaacson, D. L. and Madsen, R. W., L_{1} ergodic behavior of non-negative kernels, 17

(1974), 350–363

Isbell, J. R., Contract law and the value of a game, 5 (1967), 135–144 Isbell, J. R., Functorial implicit operations, 15 (1973), 185–188 Ish-Shalom, A., On 2-Sylow intersections, 18 (1974), 235–242 Ish-Shalom, A., On central 2-Sylow intersections, 27 (1977), 339–347 Ito, T. see Brown, L. Ito, Y. see Hajian, A. Ivanov, A. A. and Shpectorov, S. V., The last flag-transitive P -geometry, 82 (1993),

341–362

Ivorra, C. see Blasco, J. L. Iwanik, A., Lemańczyk, M. and Rudolph, D. J., Absolutely continuous cocycles over

irrational rotations, 83 (1993), 73–95

Iyahen, S. O., On the closed graph theorem, 10 (1971), 96–105 Jablow, E. R., An analogue of the Rauch variational formula for Prym differentials, 65

(1989), 323–355

Jabotinsky, E., On the use of spherical means in hydrodynamics. Application to the study of the inception of turbulence, 9 (1971), 3–19 Jackson, I. R. H. see Dyn, N. Jacob, B. and Wadsworth, A. R., Division algebras with no common subfields, 83

(1993), 353–360

Jacquet, H., The continuous spectrum of the relative trace formula for GL(3) over a

quadratic extension, 89 (1995), 1–59

Jain, N. C. and Orey, S., On the range of random walk, 6 (1968), 373–380 Jain, N. C. and Orey, S., Vague convergence of sums of independent random variables,

33 (1979), 317–348

Jain, S. K. see Goel, V. K. Jakimovski, A. and Levikson, B., A new approximation operator of the Arato-Renyi

type, 23 (1976), 39–52

Jakimovski, A. and Livne, A., Approximation operators and Tauberian constants, 7

(1969), 263–292

Jakimovski, A. see Ditzian, Z. Jakóbczak, P., Division by a holomorphic matrix in strictly pseudoconvex domains, 71

(1990), 145–159

Jakóbczak, P., Highly nonintegrable functions in the unit ball, 97 (1997), 175–181 James, G. D., A note on the T -ideal generated by s_{3}[x_{1}, x_{2}, x_{3}], 29 (1978), 105–112 James, R. C., A counterexample for a sup theorem in normed spaces, 9 (1971), 511–512 James, R. C., A nonreflexive Banach space that is uniformly nonoctahedral, 18 (1974),

145–155

James, R. C., Nonreflexive spaces of type 2, 30 (1978), 1–13 James, R. C., Reflexivity and the sup of linear functionals, 13 (1972), 289–300 James, R. C., Weak compactness and reflexivity, 2 (1964), 101–119 James, R. C. and Schäffer, J. J., Super-reflexivity and the girth of spheres, 11 (1972),

398–404

Jamison, B. see Chacon, R. V. Jara, P. and Santos, E., Completions of commutative rings and modules, 89 (1995),

141–155

Jara, P. see Bueso, J. L. Jaramillo, J. A. see Gonzalo, R. Jarchow, H. and Swart, J., On Mackey convergence in locally convex spaces, 16 (1973),

150–158

Jarden, M., Algebraic extensions of finite corank of Hilbertian fields, 18 (1974), 279–307 Jarden, M., The inverse Galois problem over formal power series fields, 85 (1994), 263–275 Jarden, M. and Razon, A., Pseudo algebraically closed fields over rings, 86 (1994), 25–59 Jarden, M. and Ritter, J., The Frattini subgroup of the absolute Galois group of a local

field, 74 (1991), 81–90

Jarden, M. see Fried, M. D. Jarden, M. see Geyer, W.-D. Jarosz, K., Any Banach space has an equivalent norm with trivial isometries, 64 (1988),

49–56

Jaworski, W., The asymptotic σ-algebra of a recurrent random walk on a locally compact

group, 94 (1996), 201–219

Jech, T. J., ω_{1} can be measurable, 6 (1968), 363–367 Jech, T. J., A maximal definable σ-ideal over ω_{1}, 39 (1981), 155–160 Jech, T. J., Simple complete Boolean algebras, 18 (1974), 1–10 Jech, T. J. and Shelah, S., A note on canonical functions, 68 (1989), 376–380 Jekel, S. M., A simplicial formula and bound for the Euler class, 66 (1989), 247–259 Jenda, O. M. G. see Cheatham, T. J. Jensen, A. see Beaulieu, R. A. Jeroslow, R. G. and Kortanek, K. O., On semi-infinite systems of linear inequalities,

10 (1971), 252–258 Jespers, E., Jacobson rings and rings strongly graded by an abelian group, 63 (1988),

67–78

Jiang, Q. and Peng, L., Phase space, wavelet transform and Toeplitz–Hankel type

operators, 89 (1995), 157–171

Jin, R. see Shelah, S. Jockusch, C. G., Jr., Upward closure and cohesive degrees, 15 (1973), 332–335 Jockusch, C. G., Jr. and Posner, D. B., Automorphism bases for degrees of unsolv-

ability, 40 (1981), 150–164

Jockusch, C. G., Jr. and Solovay, R. M., Fixed points of jump preserving automor-

phisms of degrees, 26 (1977), 91–94

Jogdeo, K. see Dharmadhikari, S. W. John, K. and Zizler, V., A renorming of dual spaces, 12 (1972), 331–336 John, K. and Zizler, V., On the heredity of weak compact generating, 20 (1975), 228–236 Johnson, A. S. A., Measures on the circle invariant under multiplication by a nonlacunary

subsemigroup of the integers, 77 (1992), 211–240

Johnson, B. E. see Grønbæk, N. Johnson, D. L., A note on the Fibonacci groups, 17 (1974), 277–282 Johnson, D. L. and Wamsley, J. W., Minimal relations for certain finite p-groups, 8

(1970), 349–356

Johnson, L. see Brown, E. Johnson, R. A., Minimal functions with unbounded integral, 31 (1978), 133–141 Johnson, W. B., A complementary universal conjugate Banach space and its relation to

the approximation problem, 13 (1972), 301–310

Johnson, W. B., Factoring compact operators, 9 (1971), 337–345 Johnson, W. B. and Lindenstrauss, J., Correction to: “Some remarks on weakly

compactly generated Banach spaces” [Israel J. Math. 17 (1974), 219–230], 32 (1979), 382–383

Johnson, W. B. and Lindenstrauss, J., Some remarks on weakly compactly generated

Banach spaces, 17 (1974), 219–230

Johnson, W. B., Lindenstrauss, J. and Schechtman, G., Extensions of Lipschitz

maps into Banach spaces, 54 (1986), 129–138

Johnson, W. B., Lindenstrauss, J. and Schechtman, G., On the relation between

several notions of unconditional structure, 37 (1980), 120–129

Johnson, W. B., Rosenthal, H. P. and Zippin, M., On bases, finite dimensional

decompositions and weaker structures in Banach spaces, 9 (1971), 488–506

Johnson, W. B. and Schechtman, G., Computing p-summing norms with few vectors,

87 (1994), 19–31

Johnson, W. B. and Schechtman, G., Martingale inequalities in rearrangement invariant

function spaces, 64 (1988), 267–275

Johnson, W. B. and Tzafriri, L., On the local structure of subspaces of Banach lattices,

20 (1975), 292–299

Johnson, W. B. and Zippin, M., On subspaces of quotients of (∑ G_{n})_{l_{p}} and (∑ G_{n})_{c_{0}},

13 (1972), 311–316

Johnson, W. B. and Zippin, M., Separable L_{1} preduals are quotients of C(Δ), 16 (1973),

198–202

Johnson, W. B. and Zippin, M., Subspaces and quotient spaces of (∑ G_{n})_{l_{p}} and (∑ G_{n})_{c_{0}},

17 (1974), 50–55

Johnson, W. B. see Bennett, G. Johnson, W. B. see Casazza, P. G. Johnson, W. B. see Davis, W. J. Johnson, W. B. see Figiel, T. Johnson, W. B. see Girardi, M. Johnson, W. B. see Hagler, J. Jøndrup, S. see Beaulieu, R. A.

Jones, R. L. and Olsen, J., Subsequence pointwise ergodic theorems for operators in L^{p},

77 (1992), 33–54

Jordan, B. W. see Faltings, G. Jordan, D. A., Zero-divisors in completions of non-commutative rings, 80 (1992), 351–358 Jorgensen, P. E. T., Unitary dilations and the C^{∗}-algebra O_{2}, 56 (1986), 129–142 Jorgenson, J., On directional derivatives of the theta function along its divisor, 77 (1992),

273–284

Josefson, B., A Banach space containing non-trivial limited sets but no non-trivial

bounding sets, 71 (1990), 321–327

Joseph, A., A generalization of Quillen’s lemma and its application to the Weyl algebras,

28 (1977), 177–192

Joseph, A. and Small, L. W., An additivity principle for Goldie rank, 31 (1978), 105–114 Joyner, D., On the metaplectic analog of Kazhdan’s “endoscopic” lifting, 61 (1988),

113–154

Judah, H. and Repický, M., No random reals in countable support iterations, 92 (1995),

349–359

Judah, H. and Shelah, S., Baire property and axiom of choice, 84 (1993), 435–450 Judah, H. and Shelah, S., MA(σ-centered): Cohen reals, strong measure zero sets and

strongly meager sets, 68 (1989), 1–17

Judah, H. see Brendle, J. Judah, H. see Goldstern, M. Juhász, A., Engel groups I, 69 (1990), 1–24 Juhász, A., On finite groups with a Sylow p-subgroup of type (m, n), 36 (1980), 133–168 Juhász, A., Small cancellation theory with a weakened small cancellation hypothesis. I.

The basic theory, 55 (1986), 65–93

Juhász, A., Small cancellation theory with a weakened small cancellation hypothesis. II.

The word problem, 58 (1987), 19–36

Juhász, A., Small cancellation theory with a weakened small cancellation hypothesis. III.

The conjugacy problem, 58 (1987), 37–53

Juhász, I., On the weight-spectrum of a compact space, 81 (1993), 369–379 Juhász, I. and Shelah, S., How large can a hereditarily separable or hereditarily Lindelöf

space be?, 53 (1986), 355–364

Juhász, I., Shelah, S. and Soukup, L., More on countably compact, locally countable

spaces, 62 (1988), 302–310

Juhász, I. see Dow, A. del Junco, A., A family of counterexamples in ergodic theory, 44 (1983), 160–188 del Junco, A. and Rudolph, D. J., Residual behavior of induced maps, 93 (1996),

387–398

Junge, M. and Defant, M., Some estimates on entropy numbers, 84 (1993), 417–433 Just, W. see Burke, M. R. Kahane, J.-P., Une inégalité du type de Slepian et Gordon sur les processus gaussiens, 55

(1986), 109–110

Kahane, J.-P. and Katznelson, Y., Contribution à deux problèmes, concernant les

fonctions de la classe A, 1 (1963), 110–131

Kahane, J.-P. and Katznelson, Y., Lignes de niveau et series de Fourier absolument

convergentes, 6 (1968), 346–353

Kahane, J.-P. and Katznelson, Y., Mauvaise approximation par polynomes trigono-

métriques dans certaines algèbres de convolution, 31 (1978), 217–223

Kahn, J. see Bourgain, J. Kaimanovich, V., The Poisson boundary of covering Markov operators, 89 (1995), 77–134 Kainen, P. C. see Hechler, S. H. Kakutani, S. see Hajian, A. Kalai, G., Characterization of f-vectors of families of convex sets in R^{d}. Part I: Necessity

of Eckhoff’s conditions, 48 (1984), 175–195 Kalai, G., Enumeration of Q-acyclic simplicial complexes, 45 (1983), 337–351 Kalai, G., Intersection patterns of convex sets, 48 (1984), 161–174 Kalai, G. see Bourgain, J. Kalf, H., The quantum mechanical virial theorem and the absence of positive energy

bound states of Schrödinger operators, 20 (1975), 57–69

Kalikow, S., Random Markov processes and uniform martingales, 71 (1990), 33–54 Kalikow, S., Katznelson, Y. and Weiss, B., Finitarily deterministic generators for

zero entropy systems, 79 (1992), 33–45

Kalikow, S. and Weiss, B., When are random graphs connected, 62 (1988), 257–268 Kalikow, S. see Bramson, M. Kalikow, S. see Eigen, S. J. Kaliman, S., Exotic analytic structures and Eisenman intrinsic measures, 88 (1994),

411–423

Kallay, M., Indecomposable polytopes, 41 (1982), 235–243 Kallay, M., Reconstruction of a plane convex body from the curvature of its boundary,

17 (1974), 149–161

Kallay, M., The extreme bodies in the set of plane convex bodies with a given width

function, 22 (1975), 203–207

Kallay, M., The minimal number of circuits in a finite set in R^{2k−1}, 49 (1984), 360–364 Kalton, N. J., An F -space with trivial dual where the Krein–Milman theorem holds, 36

(1980), 41–49

Kalton, N. J., Banach spaces embedding into L_{0}, 52 (1985), 305–319 Kalton, N. J., Compact and strictly singular operators on Orlicz spaces, 26 (1977),

126–136

Kalton, N. J., Compact convex sets and complex convexity, 59 (1987), 29–40 Kalton, N. J., Embedding L_{1} in a Banach lattice, 32 (1979), 209–220 Kalton, N. J., Subseries convergence in topological groups and vector spaces, 10 (1971),

402–412

Kalton, N. J., Leranoz, C. and Wojtaszczyk, P., Uniqueness of unconditional bases

in quasi-Banach spaces with applications to Hardy spaces, 72 (1990), 299–311

Kalton, N. J. and Shapiro, J. H., An F -space with trivial dual and non-trivial compact

endomorphisms, 20 (1975), 282–291

Kalton, N. J. see Casazza, P. G. Kamae, T., A simple proof of the ergodic theorem using nonstandard analysis, 42 (1982),

284–290

Kamae, T., Cyclic extensions of odometer transformations and spectral disjointness, 59

(1987), 41–63

Kamae, T., Subsequences of normal sequences, 16 (1973), 121–149 Kamae, T. and Mendes France, M., Van der Corput’s difference theorem, 31 (1978),

335–342

Kamae, T. and Weiss, B., Normal numbers and selection rules, 21 (1975), 101–110 Kamae, T. see Dumont, J. M. Kamburelis, A., A new proof of the Gitik–Shelah theorem, 72 (1990), 373–380 Kamenomostskaya, S., The asymptotic behaviour of the solution of the filtration equation,

14 (1973), 76–87

Kamin, S. and Peletier, L. A., Large time behaviour of solutions of the porous media

equation with absorption, 55 (1986), 129–146

Kamin, S. and Peletier, L. A., Source-type solutions of degenerate diffusion equations

with absorption, 50 (1985), 219–230

Kamiński, B. see Binkowska, M. Kammeyer, J. W., The isomorphism theorem for relatively finitely determined Z^{n}-actions,

69 (1990), 117–127

Kanamori, A. and Taylor, A. D., Separating ultrafilters on uncountable cardinals, 47 (1984), 131–138 Kane, R. M. see Hubbuck, J. R. Kaniel, S., A sufficient condition for smoothness of solutions of Navier-Stokes equations,

6 (1968), 354–358

Kaniel, S., Construction of a fixed point for contractions in Banach space, 9 (1971),

535–540

Kaniel, S., On the approximation of symmetric operators by operators of finite rank, 3

(1965), 1–5

Kaniel, S. see Har-El, J. Kannai, Y., An unsolvable hypoelliptic differential operator, 9 (1971), 306–315 Kannai, Y., Existence of a utility in infinite dimensional partially ordered spaces, 1 (1963),

229–234

Kannai, Y., Hypoelliptic ordinary differential operators, 13 (1972), 106–134 Kannai, Y., Nonexistence for a boundary value problem arising in parabolic theory, 71

(1990), 349–351

Kannai, Y., Values of games with a continuum of players, 4 (1966), 54–58 Kannai, Y. see Agmon, S. Kano, H. and Shiokawa, I., Rings of normal and nonnormal numbers, 84 (1993), 403–416 Kantor, W. M., Generalized quadrangles having a prime parameter, 23 (1976), 8–18 Kantor, W. M., Line-transitive collineation groups of finite projective spaces, 14 (1973),

229–235

Kantor, W. M., On homologies of finite projective planes, 16 (1973), 351–361 Kantor, W. M., On point-transitive affine planes, 42 (1982), 227–234 Kantor, W. M., Primitive groups having transitive subgroups of smaller, prime power

degree, 18 (1974), 141–143

Kantor, W. M., Strongly regular graphs defined by spreads, 41 (1982), 298–312 Karlin, S. and Micchelli, C. A., The fundamental theorem of algebra for monosplines

satisfying boundary conditions, 11 (1972), 405–451

Karlin, S. and Ziegler, Z., Generalized absolutely monotone functions, 3 (1965), 173–180 Karush, J. see Dubins, L. E. Kasana, H. S. see Einstein-Matthews, S. M. Katchalski, M., The dimension of intersections of convex sets, 10 (1971), 465–470 Katchalski, M., The dimension of intersections of convex sets II, 26 (1977), 209–213 Kato, T., Schrödinger operators with singular potentials, 13 (1972), 135–148 Katok, A., Interval exchange transformations and some special flows are not mixing, 35

(1980), 301–310

Katok, A. and Lewis, J., Global rigidity results for lattice actions on tori and new

examples of volume-preserving actions, 93 (1996), 253–280

Katok, A. and Lewis, J., Local rigidity for certain groups of toral automorphisms, 75

(1991), 203–241

Katok, S. and Sarnak, P., Heegner points, cycles and Maass forms, 84 (1993), 193–227 Katz, A. see Grabarnik, G. Ya. Katz, M., Rearrangements of (0 − 1) matrices, 9 (1971), 53–72 Katznelson, Y., Ergodic automorphisms of T^{n} are Bernoulli shifts, 10 (1971), 186–195 Katznelson, Y., On αβ-sets, 33 (1979), 1–4 Katznelson, Y. and Körner, T. W., An algebra dense in its tilde algebra, 17 (1974),

248–260

Katznelson, Y. and Malliavin, P., Analyse harmonique dans quelques algèbres ho-

mogènes, 5 (1967), 107–117

Katznelson, Y. and McGehee, O. C., Some sets obeying harmonic synthesis, 23 (1976),

88–93

Katznelson, Y. and Weiss, B., A simple proof of some ergodic theorems, 42 (1982),

291–296

Katznelson, Y. and Weiss, B., Commuting measure-preserving transformations, 12 (1972), 161–173 Katznelson, Y. and Weiss, B., The construction of quasi-invariant measures, 12 (1972),

1–4

Katznelson, Y. see Auslander, J. Katznelson, Y. see Bourgain, J. Katznelson, Y. see deLeeuw, K. Katznelson, Y. see Furstenberg, H. Katznelson, Y. see Hirschman, Jr., I. I. Katznelson, Y. see Kahane, J.-P. Katznelson, Y. see Kalikow, S. Kaufman, R. P., A functional method for linear sets, 5 (1967), 185–187 Kaufman, R. P., A functional method for linear sets, II, 7 (1969), 293–298 Kaufman, R. P., A note on tilde algebras, 67 (1989), 117–122 Kaufman, R. P., On some operators in c_{0}, 50 (1985), 353–356 Kaufman, R. P., On the symbolic calculus of Bernoulli convolutions, 6 (1968), 30–35 Kaufman, R. P., Sobolev spaces and symbolic calculus, 35 (1980), 222–224 Kaufman, R. P. see Ash, J. M. Kaul, R. N., Union curvature of a vector field in a subspace, 3 (1965), 15–18 Kay, D. C. and Guay, M. D., Convexity and a certain property P_{m}, 8 (1970), 39–52 Kay, D. C. see Breen, M. Kay, D. C. see Guay, M. D. Kazhdan, D., On ε-representations, 43 (1982), 315–323 Kazhdan, D., On arithmetic varieties II, 44 (1983), 139–159 Kazhdan, D., Proof of Springer’s hypothesis, 28 (1977), 272–286 Kazhdan, D. and Lusztig, G., Fixed point varieties on affine flag manifolds, 62 (1988),

129–168

Kazhdan, D. see De Concini, C. Keane, M., Non-ergodic interval exchange transformations, 26 (1977), 188–196 Keane, M. and Smorodinsky, M., A class of finitary codes, 26 (1977), 352–371 Keane, M. and Smorodinsky, M., Finitary isomorphisms of irreducible Markov shifts,

34 (1979), 281–286

Keane, M. see Aaronson, J. Keane, M. see Burton, R. M. Keane, M. see Denker, M. Kechris, A. S., Hereditary properties of the class of closed sets of uniqueness for

trigonometric series, 73 (1991), 189–198

Kechris, A. S. and Moschovakis, Y. N., Two theorems about projective sets, 12 (1972),

391–399

Kechris, A. S. and Solecki, S., Approximation of analytic by Borel sets and definable

countable chain conditions, 89 (1995), 343–356

Keene, F. W., Lipsman, R. L. and Wolf, J. A., The Plancherel formula for parabolic

subgroups, 28 (1977), 68–90

Keisler, H. J., Some model theoretic results for ω-logic, 4 (1966), 249–261 Keisler, H. J. and Morley, M. D., Elementary extensions of models of set theory, 6

(1968), 49–65

Keisler, H. J. and Morley, M. D., On the number of homogeneous models of a given

power, 5 (1967), 73–78

Kella, O., Star search—a different show, 81 (1993), 145–159 Keller, G., Exponential weak Bernoulli mixing for Collet–Eckmann maps, 86 (1994),

301–310

Keller, J. see Kharchenko, V. K. Kelly, L. M. see Grünbaum, B. Kemer, A. R., Remarks on the prime varieties, 96 (1996), 341–356 Kemer, A. R., The standard identity in characteristic p: A conjecture of I. B. Volichenko,

81 (1993), 343–355 Kenmochi, N., Some nonlinear parabolic variational inequalities, 22 (1975), 304–331 Kenyon, R., Projecting the one-dimensional Sierpinski gasket, 97 (1997), 221–238 Kenyon, R. and Peres, Y., Correction to “Hausdorff dimensions of sofic affine-invariant

sets” [Israel J. Math. 94 (1996) 157–178], 97 (1997), 347–347

Kenyon, R. and Peres, Y., Hausdorff dimensions of sofic affine-invariant sets, 94 (1996),

157–178

Kerr, J. W., Very long chains of annihilator ideals, 46 (1983), 197–204 Kertz, R. P. and Rösler, U., Martingales with given maxima and terminal distributions,

69 (1990), 173–192

Kertz, R. P. and Rösler, U., Stochastic and convex orders and lattices of probability

measures, with a martingale interpretation, 77 (1992), 129–164

Kesten, H., A Tauberian theorem for random walk, 6 (1968), 279–294 Kesten, H., The speed of convergence of a martingale, 32 (1979), 83–96 Ketonen, J. see Benda, M. Keynes, H. B. and Newton, D., The structure of ergodic measures for compact group

extensions, 18 (1974), 363–389

Keynes, H. B. see Ellis, R. Keynes, H. B. see Furstenberg, H. Kharchenko, V. K., Keller, J. and Rodrígues-Romo, S., Prime rings with PI rings

of constants, 96 (1996), 357–377

Kiefer, J. E. and Weiss, G. H., Some asymptotic Bessel function ratios, 12 (1972), 46–48 Kieffer, J. C., A direct proof that VWB processes are closed in the d̄-metric, 41 (1982),

154–160

Kieffer, J. C. and Rahe, M., Finitely fixed implies loosely Bernoulli, a direct proof, 43

(1982), 97–104

Kifer, Y., A lower bound for Hausdorff dimensions of harmonic measures on negatively

curved manifolds, 71 (1990), 339–348

Kifer, Y., Bounded harmonic functions on nonamenable covers of compact manifolds, 61

(1988), 170–178

Kifer, Y., Principal eigenvalues, topological pressure, and stochastic stability of equilibrium

states, 70 (1990), 1–47

Kifer, Y., Spectrum, harmonic functions, and hyperbolic metric spaces, 89 (1995), 377–428 Kifer, Y., The exit problem for small random perturbations of dynamical systems with a

hyperbolic fixed point, 40 (1981), 74–96

Kifer, Y., The inverse problem for small random perturbations of dynamical systems, 40

(1981), 165–174

Kifer, Y. and Newhouse, S. E., A global volume lemma and applications, 74 (1991),

209–223

Kifer, Y. see Furstenberg, H. Kim, J-H. see Alon, N. Kim, J. M., Bae, S. H. and Lee, I.-S., Cyclotomic units in Z_{p}-extensions, 75 (1991),

161–165

Kinderlehrer, D., Some open questions about variational inequalities, 13 (1972), 149–154 Kinderlehrer, D., Variational inequalities with lower dimensional obstacles, 10 (1971),

339–348

King, J. H. see Steinhorn, C. I. King, J. L., A lower bound on the rank of mixing extensions, 59 (1987), 377–380 King, J. L., Ergodic properties where order 4 implies infinite order, 80 (1992), 65–86 King, J. L., For mixing transformations rank(T^{k}) = k · rank(T), 56 (1986), 102–122 King, J. L., Lightly mixing is closed under countable products, 62 (1988), 341–346 King, J. L. see Akcoglu, M. A. King, J. L. see Friedman, N. A. Kinney, J. R., A thin set of lines, 8 (1970), 97–102 Kinney, J. R., Convex hull of Brownian motion in d-dimensions, 4 (1966), 139–143 Kinnunen, J., The Hardy–Littlewood maximal function of a Sobolev function, 100 (1997),

117–124

Kirk, R. B. see Dressler, R. E. Kirk, W. A., Fixed point theorems for non-Lipschitzian mappings of asymptotically

nonexpansive type, 17 (1974), 339–346

Kirk, W. A. and Schöneberg, R., Zeros of m-accretive operators in Banach spaces, 35

(1980), 1–8

Kirkman, E. E. and Small, L. W., q-analogs of harmonic oscillators and related rings,

81 (1993), 111–127

Kiss, G. see Bezdek, K. Kitchens, B. and Tuncel, S., Semi-groups and graphs for sofic systems, 53 (1986),

231–255

Kitchens, B. see Handel, M. Klee, V., Connectedness in topological linear spaces, 2 (1964), 127–131 Klee, V., On a conjecture of Lindenstrauss, 1 (1963), 1–4 Klee, V., Unions of increasing and intersections of decreasing sequences of convex sets, 12

(1972), 70–78

Klee, V., Utility functions and the ‘lin’ operation for convex sets, 2 (1964), 191–197 Klee, V., Maluta, E. and Zanco, C., Inspheres and inner products, 55 (1986), 1–14 Klee, V. see Bessaga, C. Klein, A. A., A note about two properties of matrix rings, 8 (1970), 90–92 Klein, A. A., The sum of nil one-sided ideals of bounded index of a ring, 88 (1994), 25–30 Klein, A. A. and Makar-Limanov, L., Skew fields of differential operators, 72 (1990),

281–287

Klein, A. A. and Regev, A., The codimensions of a PI-algebra, 12 (1972), 421–426 Klein, T., Groups whose proper factors are all abelian, 9 (1971), 362–366 Kleinschmidt, P., Pachner, U. and Wills, J. M., On L. Fejes Tóth’s “sausage-

conjecture”, 47 (1984), 216–226

Kleinschmidt, P. see Bokowski, J. Klemes, I. and Reinhold, K., Rank one transformations with singular spectral type, 98

(1997), 1–14

Knaust, H. and Odell, E., On c_{0} sequences in Banach spaces, 67 (1989), 153–169 Knight, J. F., Theories whose resplendent models are homogeneous, 42 (1982), 151–161 Knörr, R., Lempken, W. and Thielcke, B., The S_{3}-conjecture for solvable groups, 91

(1995), 61–76

Knus, M-A. see Haile, D. E. Kohayakawa, Y. see Haxell, P. E. Kohlberg, E. and Neyman, A., Asymptotic behavior of nonexpansive mappings in

normed linear spaces, 38 (1981), 269–275

Kojman, M., Perles, M. A. and Shelah, S., Sets in a Euclidean space which are not a

countable union of convex subsets, 70 (1990), 313–342

Kojman, M. and Shelah, S., Universal abelian groups, 92 (1995), 113–124 Koldobsky, A. L. see Dilworth, S. J. Kolesnik, G. see Berend, D. Komjáth, P., Connectivity and chromatic number of infinite graphs, 56 (1986), 257–266 Komjáth, P., Consistency results on infinite graphs, 61 (1988), 285–294 Komjáth, P., Mekler, A. H. and Pach, J., Some universal graphs, 64 (1988), 158–168 Komorowski, R. A. and Tomczak-Jaegermann, N., Banach spaces without local

unconditional structure, 89 (1995), 205–226

König, H., Spaces with large projection constants, 50 (1985), 181–188 König, H., Type constants and (q, 2)-summing norms defined by n vectors, 37 (1980),

130–138

Kononenko, A. and Yue, C. B., Cohomology and rigidity of Fuchsian groups, 97 (1997), 51–59 Kopeliovich, Y. see Farkas, H. M. Koppel, M., Towards a new theory of confirmation, 50 (1985), 207–218 Koren, M., Realization of a sum of sequences by a sum graph, 15 (1973), 396–403 Körner, T. W. see Blei, R. C. Körner, T. W. see Katznelson, Y. Kornfeld, I. and Lin, M., Coboundaries of irreducible Markov operators on C(K), 97

(1997), 189–202

Koronel, J. D., Continuity and kth order differentiability in Orlicz–Sobolev spaces:

W^{k}L_{A}, 24 (1976), 119–138

Kortanek, K. O. see Jeroslow, R. G. Kottman, C. A., A characterization of the Euclidean topology among the affine topologies,

10 (1971), 212–217

Koufogiorgos, T. see Blair, D. E. Kovács, L. G., Primitive permutation groups of simple diagonal type, 63 (1988), 119–127 Kra, B., The conjugating map for commutative groups of circle diffeomorphisms, 93 (1996),

303–316

Kra, I., Families of univalent functions and Kleinian groups, 60 (1987), 89–127 Kra, I., On new kinds of Teichmüller spaces, 16 (1973), 237–257 Kra, I., On Teichmüller’s theorem on the quasi-invariance of cross ratios, 30 (1978),

152–158

Kra, I. see Farkas, H. M. Kraft, R., Riemann functions for a linear system of hyperbolic form, 5 (1967), 53–56 Kraft, R., Riemann functions for a system of hyperbolic form in three independent

variables, 5 (1967), 44–52

Krattenthaler, C., On bideterminantal formulas for characters of classical groups, 90

(1995), 67–79

Krause, G., Flat embeddings of Noetherian algebras in Artinian rings, 77 (1992), 97–114 Kräuter, A. R. and Seifter, N., On some questions concerning permanents of (1, −1)-

matrices, 45 (1983), 53–62

Krengel, U., An example concerning the nonlinear pointwise ergodic theorem, 58 (1987),

193–197

Krengel, U. and Lin, M., Order preserving nonexpansive operators in L_{1}, 58 (1987),

170–192

Krengel, U., Lin, M. and Wittmann, R., A limit theorem for order preserving

nonexpansive operators in L_{1}, 71 (1990), 181–191

Krengel, U. see Akcoglu, M. A. Krieger, H. A., A new look at Bergström’s theorem on convergence in distribution for

sums of dependent random variables, 47 (1984), 32–64

Krieger, W., On sofic systems I, 48 (1984), 305–330 Krieger, W., On sofic systems II, 60 (1987), 167–176 Krivine, J.-L. see Dacunha-Castelle, D. Krivine, J.-L. and Maurey, B., Espaces de Banach stables, 39 (1981), 273–295 Kronk, H. V. see Chartrand, G. Kudla, S. S., Splitting metaplectic covers of dual reductive pairs, 87 (1994), 361–401 Kudla, S. S. and Rallis, S., Degenerate principal series and invariant distributions, 69

(1990), 25–45

Kudla, S. S. and Rallis, S., Ramified degenerate principal series representations for

Sp(n), 78 (1992), 209–256

Kudla, S. S. and Sweet, W. J., Jr., Degenerate principal series representations for

U(n, n), 98 (1997), 253–306

Kuhlmann, F.-V., Quantifier elimination for Henselian fields relative to additive and

multiplicative congruences, 85 (1994), 277–306

Kühnel, W. and Lassmann, G., Neighborly combinatorial 3-manifolds with dihedral automorphism group, 52 (1985), 147–166 Kuiper, N. H., Double normals of convex bodies, 2 (1964), 71–80 Kulkarni, R. S., Infinite families of surface symmetries, 76 (1991), 337–343 Külshammer, B., Donovan’s conjecture, crossed products and algebraic group actions, 92

(1995), 295–306

Kunen, K. see Barwise, J. Kunisch, K. see Di Blasio, G. Kupka, J., Strong liftings with application to measurable cross sections in locally compact

groups, 44 (1983), 243–261

Kurz, M. see Aumann, R. J. Kwiatkowski, J., Lemańczyk, M. and Rudolph, D. J., A class of real cocycles having

an analytic coboundary modification, 87 (1994), 337–360

Kwiatkowski, J., Lemańczyk, M. and Rudolph, D. J., Weak isomorphism of measure-

preserving diffeomorphisms, 80 (1992), 33–64

Kwon, K.-H., On the growth of entire functions, 86 (1994), 409–427 Lachlan, A. H., On countable stable structures which are homogeneous for a finite

relational language, 49 (1984), 69–153

Lachlan, A. H. and Shelah, S., Stable structures homogeneous for a finite binary

language, 49 (1984), 155–180

Lacroix, M.-T., Sur les normes normales au sens de Benilan–Crandall et Pazy, 59 (1987),

213–224

Lafuente, J., On the second Loewy term of projectives of a group algebra, 67 (1989),

170–180

Lai, P. T., Classe de compacité d’opérateur intervenant dans une classe de problèmes

elliptiques dégénérés, 17 (1974), 364–379

Lai, P. T., Comportement asymptotique du noyau de la résolvante et des valeurs propres

d’un opérateur elliptique non nécessairement auto-adjoint, 23 (1976), 221–250

Lai, P. T. and Robert, D., Sur un problème aux valeurs propres non linéaire. [On a

nonlinear eigenvalue problem], 36 (1980), 169–186

Lam, T. Y. and Leroy, A., Recognition and computations of matrix rings, 96 (1996),

379–397

Landver, A., Finite combinations of Baire numbers, 81 (1993), 289–296 Lange, G. L., Two-generator Frattini subgroups of finite p-groups, 29 (1978), 357–360 Lannes, J. and Schwartz, L., Sur les groupes d’homotopie des espaces dont la cohomologie

modulo 2 est nilpotente, 66 (1989), 260–273

Lanski, C., Invariant additive subgroups of a certain kind, 61 (1988), 192–198 Lanski, C., Minimal differential identities in prime rings, 56 (1986), 231–246 Lapidot, E., A note on generalized absolutely monotone functions and a Besicovitch-type

problem, 17 (1974), 271–276

Lapresté, J. T. see Guerre, S. Larcher, G. see Hellekalek, P. Larman, D. G., The existence of d-polytopes whose facets have prescribed areas, 30

(1978), 51–53

Larrea, M^{a} C. see Vera López, A. Larsen, M., On the conjugacy of element-conjugate homomorphisms, 88 (1994), 253–277 Larsen, M. and Pink, R., A connectedness criterion for ℓ-adic Galois representations, 97

(1997), 1–10

Larson, J. A., A counter-example in the partition calculus for an uncountable ordinal, 36

(1980), 287–299

Larson, R. G. and Taft, E. J., The algebraic structure of linearly recursive sequences

under Hadamard product, 72 (1990), 118–132

Larson, R. G. see Grossman, R. Larsson, J.-O., Embeddings of l_{p}^{m} into l_{∞}^{n}, 1 ≤ p ≤ 2, 55 (1986), 153–161 Lascar, D., Ranks and definability in superstable theories, 23 (1976), 53–87 Lasiecka, I. see Desch, W.

Lasry, J. M. and Lions, P. L., A remark on regularization in Hilbert spaces, 55 (1986),

257–266

Lassak, M., Reduced convex bodies in the plane, 70 (1990), 365–379 Lassmann, G. see Kühnel, W. Lau, K.-S., Farthest points in weakly compact sets, 22 (1975), 168–174 Läuchli, H., Coloring infinite graphs and the Boolean prime ideal theorem, 9 (1971),

422–429

Laue, H., A cocycle construction for groups, 47 (1984), 314–322 Laue, R., On normal p-subgroups with large centers which cannot be contained in the

Frattini subgroup, 29 (1978), 155–166

Laufer, H. B., On normal two-dimensional double point singularities, 31 (1978), 315–334 Lausch, H., Conjugacy classes of maximal nilpotent subgroups, 47 (1984), 29–31 Lauwers, L. see Haran, D. Laver, R., Making the supercompactness of κ indestructible under κ-directed closed

forcing, 29 (1978), 385–388

Laver, R., Precipitousness in forcing extensions, 48 (1984), 97–108 Lavie, M., Analytic iteration and differential equations, 5 (1967), 86–92 Lawler, G. F., Intersections of random walks with random sets, 65 (1989), 113–132 Lawler, G. F. and Puckette, E. E., The disconnection exponent for simple random

walk, 99 (1997), 109–121

Lawrence, J. see Formanek, E. Lax, R. F., On the distribution of Weierstrass points on singular curves, 57 (1987), 107–115 Lazar, A. J., Affine functions on simplexes and extreme operators, 5 (1967), 31–43 Lazar, A. J., Extreme boundaries of convex bodies in l_{2}, 20 (1975), 369–374 Lazar, A. J., Polyhedral Banach spaces and extensions of compact operators, 7 (1969),

357–364

Lazar, A. J. and Lindenstrauss, J., On Banach spaces whose duals are L_{1} spaces, 4

(1966), 205–207

Lazar, A. J. and Zippin, M., On finite dimensional subspaces of Banach spaces, 3 (1965),

147–156

Lazar, A. J. see Gootman, E. C. Le Bruyn, L., Centers of generic division algebras, the rationality problem 1965–1990, 76

(1991), 97–111

Le Bruyn, L., The functional equation for Poincaré series of trace rings of generic 2 × 2

matrices, 52 (1985), 355–360

Le Merdy, C., Analytic factorizations and completely bounded maps, 88 (1994), 381–409 Lebowitz, A., A remark on degeneration of a compact Riemann surface of genus 2, 18

(1974), 349–351

Lebowitz, A., Degeneration of a compact Riemann surface of genus 2, 12 (1972), 223–236 Lebowitz, A., Handle removal on a compact Riemann surface of genus 2, 15 (1973),

189–192

Ledger, A. J., Congruence theorems for Riemannian submanifolds, 83 (1993), 295–304 Ledger, A. J., Naturally reductive locally 4-symmetric spaces, 63 (1988), 223–236 Ledoux, M., A remark on hypercontractivity and the concentration of measure phenomenon

in a compact Riemannian manifold, 69 (1990), 361–370

Ledoux, M., Sur une inégalité de H. P. Rosenthal et le théorème limite central dans les

espaces de Banach, 50 (1985), 290–318

Ledrappier, F., Harmonic measures and Bowen–Margulis measures, 71 (1990), 275–287 Ledrappier, F., Poisson boundaries of discrete groups of matrices, 50 (1985), 319–336 Lee, C. W., Two combinatorial properties of a class of simplicial polytopes, 47 (1984),

261–269

Lee, I.-S. see Kim, J. M. Lee, S. T. and Zhu, C.-B., Degenerate principal series and local theta correspondence II,

100 (1997), 29–59 Lehman, A., A result on rearrangements, 1 (1963), 22–28 Lehmann, D., LR(k) grammars and deterministic languages, 10 (1971), 526–530 Lehrer, E., Topological mixing and uniquely ergodic systems, 57 (1987), 239–255 Lehrer, E. and Pant, H., On strictly ergodic models which are not almost topologically

conjugate, 73 (1991), 1–15

Lemańczyk, M. and de Sam Lazaro, J., Spectral analysis of certain compact factors for

Gaussian dynamical systems, 98 (1997), 307–328

Lemańczyk, M. see Iwanik, A. Lemańczyk, M. see Kwiatkowski, J. Lemberg, H., Nouvelle démonstration d’un théorème de J. L. Krivine sur la finie

représentation de l_{p} dans un espace de Banach, 39 (1981), 341–348

Lempken, W. see Knörr, R. Lempp, S. see Cooper, S. B. Lenagan, T. H., Localization at height one prime ideals in affine prime PI rings, 67 (1989),

345–352

Leranoz, C. see Kalton, N. J. Lerer, L. E., On approximating the spectrum of convolution-type operators, 1. Wiener–

Hopf matricial integral operators, 30 (1978), 339–362

Lerman, M., Shore, R. A. and Soare, R. I., r-Maximal major subsets, 31 (1978), 1–18 Lerman, M. and Simpson, S. G., Maximal sets in α-recursion theory, 14 (1973), 236–247 Leron, U. and Vapne, A., Polynomial identities of related rings, 8 (1970), 127–137 Leroy, A. see Lam, T. Y. Lesigne, E. see Assani, I. Leung, D. H., Factoring operators through Hilbert space, 71 (1990), 225–227 Leung, D. H., Some isomorphically polyhedral Orlicz sequence spaces, 87 (1994), 117–128 Leung, K. H., On a generalization of Albert’s theorem, 69 (1990), 337–350 Leviatan, D., A generalized moment problem, 5 (1967), 97–103 Leviatan, D., A generalized moment problem for self-adjoint operators, 4 (1966), 113–118 Leviatan, D., Absolute Tauberian conditions for absolute Hausdorff and quasi-Hausdorff

methods, 8 (1970), 138–146

Leviatan, D., Some vector-valued Laplace transforms, 16 (1973), 73–86 Leviatan, D. see Bak, J. Levikson, B. see Jakimovski, A. Levin, A., The distribution of the poles of the best approximating rational functions and

the analytical properties of the approximated function, 24 (1976), 139–144

Levin, D. see Dyn, N. Levin, G., On the complement of the Mandelbrot set, 88 (1994), 189–212 Levin, G. and Przytycki, F., External rays to periodic points, 94 (1996), 29–57 Levin, M., Dimension and superposition of continuous functions, 70 (1990), 205–218 Levin, M., Olsen, G. H. and Sternfeld, Y., Finite dimensional Polish spaces are

extreme boundaries of convex bodies in Euclidean space, 71 (1990), 135–143

Levine, H. A. and Meier, P., A blowup result for the critical exponent in cones, 67

(1989), 129–136

Levine, H. A. see Galaktionov, V. A. Levine, J., Finitely-presented groups with long lower central series, 73 (1991), 57–64 Levinski, J.-P., Instances of the conjecture of Chang, 48 (1984), 225–243 Levinski, J.-P., Magidor, M. and Shelah, S., Chang’s conjecture for ℵ_{ω}, 69 (1990),

161–172

Levinski, J.-P. see Donder, H.-D. Levitt, G. see Gaboriau, D. Levitzki, J., On nil subrings, 1 (1963), 215–216 Lévy, A. and Solovay, R. M., Measurable cardinals and the continuum hypothesis, 5

(1967), 234–248

Lewin, M., Matching-perfect and cover-perfect graphs, 18 (1974), 345–347 Lewis, D. R. and Smith, R. R., Finite rank operators with large trace, 51 (1985), 262–272 Lewis, D. R. see Gordon, Y. Lewis, J. see Katok, A. Lewittes, J., Construction of modular forms, 24 (1976), 145–183 Liao, G.-F. and Wang, L.-Y., Almost periodicity, chain recurrence and chaos, 93 (1996),

145–156

Liardet, P., Propriétés génériques de processus croisés, 39 (1981), 303–325 Liardet, P. and Volný, D., Sums of continuous and differentiable functions in dynamical

systems, 98 (1997), 29–60

Lichtman, A. I., Growth in enveloping algebras, 47 (1984), 296–304 Lichtman, A. I., On embedding of group rings in division rings, 23 (1976), 288–297 Lichtman, A. I., On linear groups over a field of fractions of a polycyclic group ring, 42

(1982), 318–326

Lichtman, A. I., The residual nilpotency of the augmentation ideal and the residual

nilpotency of some classes of groups, 26 (1977), 276–293

Lichtman, A. I. see Formanek, E. Liggett, T. see Aaronson, J. Lima, Å., Complex Banach spaces whose duals are L_{1}-spaces, 24 (1976), 59–72 Lima, Å., The metric approximation property, norm-one projections and intersection

properties of balls, 84 (1993), 451–475

Lima, Å. and Uttersrud, U., Centers of symmetry in finite intersections of balls in

Banach spaces, 44 (1983), 189–200

Lin, B. L. see Altshuler, Z. Lin, B. L. see Casazza, P. G. Lin, J. P. and Williams, F., On 14-connected finite H-spaces, 66 (1989), 274–288 Lin, M., Conservative Markov processes on a topological space, 8 (1970), 165–186 Lin, M., Mixed ratio limit theorems for Markov processes, 8 (1970), 357–366 Lin, M., Mixing of Cartesian squares of positive operators, 11 (1972), 349–354 Lin, M., On local ergodic convergence of semi-groups and additive processes, 42 (1982),

300–308

Lin, M., Quasi-compactness and uniform ergodicity of positive operators, 29 (1978),

309–311

Lin, M. and Wittmann, R., Convergence of representation averages and of convolution

powers, 88 (1994), 125–157

Lin, M. see Aaronson, J. Lin, M. see Kornfeld, I. Lin, M. see Krengel, U. Lin, P. K. see Benyamini, Y. Lin, P.-K., B-convexity of the space of 2-summing operators, 37 (1980), 139–150 Lin, P.-K. see Carothers, N. L. Lind, D. A., Ergodic affine transformations are loosely Bernoulli, 30 (1978), 335–338 Lind, D. A., Ergodic automorphisms of the infinite torus are Bernoulli, 17 (1974), 162–168 Lind, D. A., Ergodic group automorphisms are exponentially recurrent, 41 (1982), 313–320 Lind, D. A., Split skew products, a related functional equation, and specification, 30

(1978), 236–254

Lind, D. A., The structure of skew products with ergodic group automorphisms, 28 (1977),

205–248

Lind, D. A. and Schmidt, K., Bernoullicity of solenoidal automorphisms and global fields,

87 (1994), 33–35

Lind, D. A. see Goldberger, J. Lindenstrauss, A., Deformation retracts and the Hochschild homology of polynomial

rings, 93 (1996), 317–332

Lindenstrauss, J., A remark on L_{1} spaces, 8 (1970), 80–82 Lindenstrauss, J., A remark on symmetric bases, 13 (1972), 317–320 Lindenstrauss, J., A selection theorem, 2 (1964), 201–204 Lindenstrauss, J., Notes on Klee’s paper “Polyhedral sections of convex bodies”, 4 (1966),

235–242

Lindenstrauss, J., On a problem of Nachbin concerning extension of operators, 1 (1963),

75–84

Lindenstrauss, J., On complemented subspaces of m, 5 (1967), 153–156 Lindenstrauss, J., On extreme points in l_{1}, 4 (1966), 59–61 Lindenstrauss, J., On James’s paper “Separable conjugate spaces”, 9 (1971), 279–284 Lindenstrauss, J., On operators which attain their norm, 1 (1963), 139–148 Lindenstrauss, J., On reflexive spaces having the metric approximation property, 3

(1965), 199–204

Lindenstrauss, J., On subspaces of Banach spaces without quasicomplements, 6 (1968),

36–38

Lindenstrauss, J. and Phelps, R. R., Extreme point properties of convex bodies in

reflexive Banach spaces, 6 (1968), 39–48

Lindenstrauss, J. and Rosenthal, H. P., Automorphisms in c_{0}, l_{1} and m, 7 (1969),

227–239

Lindenstrauss, J. and Rosenthal, H. P., The L_{p} spaces, 7 (1969), 325–349 Lindenstrauss, J. and Tzafriri, L., On Orlicz sequence spaces, 10 (1971), 379–390 Lindenstrauss, J. and Tzafriri, L., On Orlicz sequence spaces III, 14 (1973), 368–389 Lindenstrauss, J. and Tzafriri, L., On Orlicz sequence spaces. II, 11 (1972), 355–379 Lindenstrauss, J. and Tzafriri, L., On the complemented subspaces problem, 9 (1971),

263–269

Lindenstrauss, J. and Tzafriri, L., The uniform approximation property in Orlicz

spaces, 23 (1976), 142–155

Lindenstrauss, J. see Aharoni, I. Lindenstrauss, J. see Benyamini, Y. Lindenstrauss, J. see Bourgain, J. Lindenstrauss, J. see Dashiell, F. K. Lindenstrauss, J. see Genel, A. Lindenstrauss, J. see Johnson, W. B. Lindenstrauss, J. see Lazar, A. J. Lindqvist, P., On Liouville’s theorem for locally quasiregular mappings in R^{n}, 47 (1984),

291–295

Lindqvist, P., On p-harmonic functions in the complex plane and curvature, 63 (1988),

257–269

Ling, S., On the Q-rational cuspidal subgroup and the component group of J_{0}(p^{r}) , 99

(1997), 29–54

Ling, S., The old subvariety of J_{0}(pq) and the Eisenstein kernel in Jacobians, 84 (1993),

365–384

Linial, N. see Bourgain, J. Lions, J. L., Remarks on some nonlinear evolution problems arising in Bingham flows, 13

(1972), 155–172

Lions, P. L., Une méthode itérative de résolution d’une inéquation variationnelle, 31

(1978), 204–208

Lions, P. L. see Brezis, H. Lions, P. L. see Lasry, J. M. Lipsman, R. L., Characters of Lie groups: traceability and certain semidirect products, 24

(1976), 45–58

Lipsman, R. L., Harmonic analysis on non-semisimple symmetric spaces, 54 (1986),

335–350

Lipsman, R. L. see Keene, F. W. Liskevich, V. A. and Tuv, E. W., A priori estimates for second order elliptic equations

and their applications, 81 (1993), 257–263 Litman, A., On the monadic theory of ω_{1} without A.C., 23 (1976), 251–266 Litman, A. and Shelah, S., Models with few isomorphic expansions, 28 (1977), 331–338 Littman, W., Decay at infinity of solutions to higher order partial differential equations:

Removal of the curvature assumption, 8 (1970), 403–407

Liu, K. and Shelah, S., Cofinalities of elementary substructures of structures on ℵ_{ω}, 99

(1997), 189–205

Livne, A. see Jakimovski, A. Livné, R., Motivic orthogonal two-dimensional representations of Gal(ℚ̄/ℚ), 92 (1995),

149–156

Lluch, A. and Miquel, V., Comparing the relative volume with a revolution manifold as

a model, 84 (1993), 385–401

Lochter, M., On equivalent number fields with special Galois groups, 84 (1993), 89–96 Loeb, P. A., Applications of nonstandard analysis to ideal boundaries in potential theory,

25 (1976), 154–187

Loeb, P. A. see Brown, D. J. Loebl, M. see Aharoni, R. Lomonosov, V., An extension of Burnside’s theorem to infinite-dimensional spaces, 75

(1991), 329–339

London, D., Nonnegative matrices with stochastic powers, 2 (1964), 237–244 Longobardi, P. see Herzog, M. Lorentz, G. G., Incomplete polynomials of best approximation, 29 (1978), 132–140 Lorenz, M. and Passman, D. S., Addendum – Prime ideals in crossed products of finite

groups, 35 (1980), 311–322

Lorenz, M. and Passman, D. S., Prime ideals in crossed products of finite groups, 33

(1979), 89–132

Lotz, H. P. and Rosenthal, H. P., Embeddings of C(Δ) and L^{1}[0, 1] in Banach lattices,

31 (1978), 169–179

Louveau, A., Une méthode topologique pour l’etude de la propriété de Ramsey, 23 (1976),

97–116

Louveau, A. see Godefroy, G. Lövblom, G.-M., Almost isometries on the unit ball of l_{1}, 63 (1988), 129–138 Lövblom, G.-M., Isometries and almost isometries between spaces of continuous functions,

56 (1986), 143–159

Lovelady, D. L., Addition in a class of nonlinear Stieltjes integrators, 10 (1971), 391–396 Loveys, J. and Peterzil, Y., Linear o-minimal structures, 81 (1993), 1–30 Lowig, H. F. J., Note on the self-duality of the unrestricted distributive law in complete

lattices, 2 (1964), 170–172

Lubarsky, R. S., Admissibility spectra through ω_{1}, 60 (1987), 225–235 Lubarsky, R. S., Playing with admissibility spectra, 63 (1988), 243–256 Lubotzky, A., Mann, A. and Segal, D., Finitely generated groups of polynomial

subgroup growth, 82 (1993), 363–371

Lubotzky, A., Pyber, L. and Shalev, A., Discrete groups of slow subgroup growth, 96

(1996), 399–418

Lubotzky, A. and Shalev, A., On some Λ-analytic pro-p groups, 85 (1994), 307–337 Lubotzky, A. and van den Dries, L., Subgroups of free profinite groups and large

subfields of Q̃, 39 (1981), 25–45

Lubotzky, A. and Zimmer, R. J., Variants of Kazhdan’s property for subgroups of

semisimple groups, 66 (1989), 289–299

Lubotzky, A. see Bass, H. Lubotzky, A. see Haran, D. Lükő, G., On the mean length of the chords of a closed curve, 4 (1966), 23–32 Lunardi, A., Differentiability with respect to (t, s) of the parabolic evolution operator, 68

(1989), 161–184 Lunardi, A., On the evolution operator for abstract parabolic equations, 60 (1987),

281–314

Lusky, W., Every L_{1}-predual is complemented in a simplex space, 64 (1988), 169–178 Lusky, W., On the primariness of the Poulsen simplex space, 37 (1980), 151–163 Lust-Piquard, F., A Grothendieck factorization theorem on 2-convex Schatten spaces, 79

(1992), 331–365

Lusztig, G. see Kazhdan, D. Lutwak, E., Mixed width-integrals of convex bodies, 28 (1977), 249–253 Lutwak, E., Rotation means of projections, 58 (1987), 161–169 Luxemburg, W. A. J., On a class of valuation fields introduced by A. Robinson, 25

(1976), 189–201

Lyons, R., On measures simultaneously 2- and 3-invariant, 61 (1988), 219–224 Lyons, R. see Adams, S. Lyons, R. see Gorenstein, D. Ma, W. and Minda, D., Behavior of domain constants under conformal mappings, 91

(1995), 157–171

Maass, A. see Blanchard, F. Maass, W., Shore, R. A. and Stob, M., Splitting properties and jump classes, 39

(1981), 210–224

Macías, R. A. see Harboure, E. Macdonald, I. D., More on p-groups of Frobenius type, 56 (1986), 335–344 Macdonald, I. D., Some p-groups of Frobenius and extra-special type, 40 (1981), 350–364 Macdonald, I. D. see Chillag, D. Macias, R. see Garcia-Cuerva, J. Macintyre, A., Existentially closed structures and Jensen’s principle ♢, 25 (1976),

202–210

Macintyre, A. and Simmons, H., Algebraic properties of number theories, 22 (1975),

7–27

Macintyre, A. see Goldrei, D. C. Madras, N. and Tanny, D., Oscillating random walk with a moving boundary, 88 (1994),

333–365

Madsen, R. W. see Isaacson, D. L. Magid, A. R., Invariant subrings of separable algebras, 16 (1973), 306–313 Magid, A. R., Local structure of representation varieties, 62 (1988), 326–340 Magidor, M., Decomposition theorems for finite sequential machines, 6 (1968), 246–260 Magidor, M., On the role of supercompact and extendible cardinals in logic, 10 (1971),

147–157

Magidor, M., On the singular cardinals problem I, 28 (1977), 1–31 Magidor, M., Precipitous ideals and Σ_{4}^{1} sets, 35 (1980), 109–134 Magidor, M., There are many normal ultrafilters corresponding to a supercompact

cardinal, 9 (1971), 186–192

Magidor, M. and Moran, G., Probabilistic tree automata and context free languages, 8

(1970), 340–348

Magidor, M. see Fuchs, L. Magidor, M. see Levinski, J.-P. Maharam, D., A “planar” representation for generalized transition kernels, 78 (1992),

281–307

Maharam, D., Invertible liftings of measurable transformations, 73 (1991), 225–245 Maharam, D., Spectral representation of generalized transition kernels, 98 (1997), 15–28 Maier, B. J., On existentially closed and generic nilpotent groups, 46 (1983), 170–188 Maitra, A. and Sudderth, W. D., An operator solution of stochastic games, 78 (1992),

33–49

Maitra, A. see Dubins, L. E. Maj, M. see Herzog, M.

Majid, S., More examples of bicrossproduct and double cross product Hopf algebras, 72

(1990), 133–148

Makan, A. R., On an embedding of certain p-groups, 21 (1975), 31–37 Makar-Limanov, L., On group rings of nilpotent groups, 48 (1984), 244–248 Makar-Limanov, L., On groups of automorphisms of a class of surfaces, 69 (1990),

250–256

Makar-Limanov, L., On the hypersurface x+x^{2}y +z^{2} +t^{3} = 0 in C^{4} or a C^{3}-like threefold

which is not C^{3}, 96 (1996), 419–429

Makar-Limanov, L. see Klein, A. A. Makkai, M., A survey of basic stability theory, with particular emphasis on orthogonality

and regular types, 49 (1984), 181–238

Makkai, M. see Harnik, V. Makkai, M. see Shelah, S. Malcolmson, P. see Bokut’, L. Malle, G. and Sonn, J., Covering groups of almost simple groups as Galois groups over

ℚ^{ab}(t), 96 (1996), 431–444

Malliavin, P. see Katznelson, Y. Mallows, C. see Doyle, P. G. Maluta, E. see Klee, V. Mammone, P. and Merkurjev, A., On the corestriction of p^{n}-symbol, 76 (1991), 73–79 Mandel, A. see Gonçalves, J. Z. Mandelbrojt, S., Considérations arithmétiques dans la théorie des fonctions analytiques,

6 (1968), 5–12

Mandelker, M., Round z-filters and round subsets of βX, 7 (1969), 1–8 Manevitz, L. M., Applied model theory and metamathematics: An Abraham Robinson

memorial problem list, 49 (1984), 3–14

Manevitz, L. M., Robinson forcing is not absolute, 25 (1976), 211–232 Manevitz, L. M. and Miller, A. W., Lindelöf models of the reals: Solution to a problem

of Sikorski, 45 (1983), 209–218

Manevitz, L. M. and Weinberger, S., Discrete circle actions: A note using non-standard

analysis, 94 (1996), 147–155

Mangheni, P. J., The classification of injective Banach lattices, 48 (1984), 341–347 Mani, P., On angle sums and Steiner points of polyhedra, 9 (1971), 380–388 Mani-Levitska, P., Convex surfaces which intersect each congruent copy of themselves in

a connected set, 54 (1986), 71–80

Mankiewicz, P., A superreflexive Banach space X with L(X) admitting a homomorphism

onto the Banach algebra C(βN), 65 (1989), 1–16

Mankiewicz, P. and Nielsen, N. J., A superreflexive Banach space with a finite dimensional decomposition so that no large subspace has a basis, 70 (1990), 188–204

Mankiewicz, P. and Tomczak-Jaegermann, N., A solution of the finite-dimensional

homogeneous Banach space problem, 75 (1991), 129–159

Mann, A., Conjugacy classes in finite groups, 31 (1978), 78–84 Mann, A., Generators of 2-groups, 10 (1971), 158–159 Mann, A., Groups with dense normal subgroups, 6 (1968), 13–25 Mann, A., Injectors and normal subgroups of finite groups, 9 (1971), 554–558 Mann, A., On subgroups of finite soluble groups III, 16 (1973), 446–451 Mann, A., Philip Hall’s ‘rather curious’ formula for abelian p-groups, 96 (1996), 445–448 Mann, A., Regular p-groups, 10 (1971), 471–477 Mann, A., Regular p-groups. II, 14 (1973), 294–303 Mann, A., Shimshon Avraham Amitsur (1921–1994), 96 (1996), ix–xxvii Mann, A., Simple groups having p-nilpotent 2nd-maximal subgroups, 6 (1968), 233–245 Mann, A., Soluble subgroups of symmetric and linear groups, 55 (1986), 162–172 Mann, A., Some finite groups with large conjugacy classes, 71 (1990), 55–63 Mann, A., Some properties of polynomial subgroup growth groups, 82 (1993), 373–380 Mann, A. and Shalev, A., Simple groups, maximal subgroups, and probabilistic aspects

of profinite groups, 96 (1996), 449–468

Mann, A. and Shalev, A., The nilpotency class of the unit group of a modular group

algebra II, 70 (1990), 267–277

Mann, A. see Chillag, D. Mann, A. see Lubotzky, A. Mansfield, R., A Souslin operation for Π_{2}^{1}, 9 (1971), 367–379 Mansfield, R., How many slopes in a polygon?, 39 (1981), 265–272 Manz, O. see Chillag, D. Marchi, M. V. see Cellina, A. Marcus, B., Unique ergodicity of some flows related to Axiom A diffeomorphisms, 21

(1975), 111–132

Marcus, B., Unique ergodicity of the horocycle flow: variable negative curvature case, 21

(1975), 133–144

Marcus, B. see Bowen, R. Marcus, B. see Cawley, E. Marcus, F., On a class of Ribaucour’s normal congruences, 12 (1972), 190–200 Marcus, F., Sur le théorème de Fubini concernant la réciproque du théorème de per-

mutabilité de Bianchi, 22 (1975), 185–202

Marcus, L., A minimal prime model with an infinite set of indiscernibles, 11 (1972),

180–183

Marcus, L., Minimal models of theories of one function symbol, 18 (1974), 117–131 Margolis, S. W. and Sapir, M. V., Quasi-identities of finite semigroups and symbolic

dynamics, 92 (1995), 317–331

Margulis, G. A. see Abels, H. Markley, N. G., Substitution-like minimal sets, 22 (1975), 332–353 Markley, N. G. and Paul, M. E., Almost automorphic symbolic minimal sets without

unique ergodicity, 34 (1979), 259–272

Marrero, O., A property of (v, k, λ)-designs, 12 (1972), 277–278 Martens, M. see Galeeva, R. Martin, R. H., Jr. see Gutman, S. Martindale, W. S., 3rd see Desmarais, P. C. Martinez, C. and Zelmanov, E. I., Products of powers in finite simple groups, 96 (1996),

469–479

Martinez Hernandez, J. see Asensio Mayor, J. Martio, O. and Srebro, U., Universal radius of injectivity for locally quasiconformal

mappings, 29 (1978), 17–23

Martínez, S. see Collet, P. Marton, K., On weak asymptotic isomorphy of memoryless correlated sources, 58 (1987),

288–304

Marton, K. and Shields, P. C., The positive-divergence and blowing-up properties, 86

(1994), 331–348

Maruyama, K. and Mimura, M., Nilpotent subgroups of the group of self-homotopy

equivalences, 72 (1990), 313–319

Maschler, M., The inequalities that determine the bargaining set M_{1}^{(i)}, 4 (1966), 127–134 Mascioni, V., s-numbers of projections in Banach spaces, 67 (1989), 82–94 Maskit, B. and Swarup, G., Two parabolic generator Kleinian groups, 64 (1988), 257–266 Mason, A. W., Quotients of the congruence kernels of SL_{2} over arithmetic Dedekind

domains, 91 (1995), 77–91

Mastrangelo, L. B., On radicals of triangular operator algebras, 84 (1993), 33–51 Masuoka, A., Existence of a unique maximal subcoalgebra whose action is inner, 72

(1990), 149–157 Masuoka, A., Semisimple Hopf algebras of dimension 6, 8, 92 (1995), 361–373 Matczuk, J. see Grzeszczuk, P. Matelski, J. P., The classification of discrete 2-generator subgroups of PSL(2, R), 42

(1982), 309–317

Matet, P. see Donder, H.-D. Mathews, J. H., A bounded normal light interior function that possesses no point

asymptotic values, 7 (1969), 381–383

Mathur, A. see Singal, M. K. Matlis, E., The ring as a torsion-free cover, 37 (1980), 211–230 Matous̆ek, J., On the distortion required for embedding finite metric spaces into normed

spaces, 93 (1996), 333–344

Matsuoka, C., A certain class of Jensen measures for uniform algebras, 60 (1987), 129–145 Matsuoka, C., Factoring functions of Maurey’s factorization theorem, 57 (1987), 318–326 Matsuoka, C., Jensen measures of uniform algebras, 61 (1988), 155–169 Mattila, K., On normal and compact operators on Banach spaces, 37 (1980), 164–170 Mauldin, R. D. and Monticino, M. G., Randomly generated distributions, 91 (1995),

215–237

Maurey, B. see Aharoni, I. Maurey, B. see Krivine, J.-L. Mayer-Wolf, E., Isometries between Banach spaces of Lipschitz functions, 38 (1981),

58–74

McCarthy, C. A., c_{p}, 5 (1967), 249–271 McConnell, J. C., I-adic completions of noncommutative rings, 32 (1979), 305–310 McCullough, D. and Rubin, L. R., Sequentially equidistant points in metric spaces, 69

(1990), 75–93

McGehee, O. C., Sets of uniqueness and sets of multiplicity, 4 (1966), 83–96 McGehee, O. C. see Katznelson, Y. McKean, H. P., The central limit theorem for Carleman’s equation, 21 (1975), 54–92 McKelvey, R. W. see Herod, J. V. McLaughlin, T. G., Embeddings of and into Nerode semirings, 60 (1987), 65–88 McMullen, P., Indecomposable convex polytopes, 58 (1987), 321–323 McMullen, P., On a problem of Klee concerning convex polytopes, 8 (1970), 1–4 McMullen, P., On the inner parallel body of a convex body, 19 (1974), 217–219 McMullen, P., Polytopes with centrally symmetric faces, 8 (1970), 194–196 McMullen, P., Polytopes with centrally symmetric facets, 23 (1976), 337–338 McMullen, P., The numbers of faces of simplicial polytopes, 9 (1971), 559–570 McMullen, P., Schulz, C. and Wills, J. M., Equivelar polyhedral manifolds in E^{3}, 41

(1982), 331–346

McMullen, P., Schulz, C. and Wills, J. M., Polyhedral 2-manifolds in E^{3} with

unusually large genus, 46 (1983), 127–144

McMullen, P., Schulz, C. and Wills, J. M., Two remarks on equivelar manifolds, 52

(1985), 28–32

Meester, R. W. J. and Nowicki, T., Infinite clusters and critical values in two-

dimensional circle percolation, 68 (1989), 63–81

Megiddo, N., The kernel and the nucleolus of a product of simple games, 9 (1971), 210–221 Meier, P. see Levine, H. A. Meilijson, I., Mixing properties of a class of skew-products, 19 (1974), 266–270 Meilijson, I., The average of the values of a function at random points, 15 (1973), 193–203 Meilijson, I., The expected value of some functions of the convex hull of a random set of

points sampled in R^{d}, 72 (1990), 341–352

Meir, A., Analytic continuation by summation-methods, 1 (1963), 224–228 Meir, A., Tauberian theorems, 1 (1963), 29–36 Meixner, T., Power automorphisms of finite p-groups, 38 (1981), 345–360 Meixner, T., The Fitting length of solvable H_{p^{n}}-groups, 51 (1985), 68–78 Mejzler, D., A note on Erlang’s formulas, 3 (1965), 157–162 Mejzler, D., Extreme value limit laws in the nonidentically distributed case, 57 (1987),

1–27

Mejzler, D., On a certain class of infinitely divisible distributions, 16 (1973), 1–19 Mekler, A. H. and Shelah, S., Almost free algebras, 89 (1995), 237–259 Mekler, A. H. and Shelah, S., Every coseparable group may be free, 81 (1993), 161–178 Mekler, A. H. and Shelah, S., The consistency strength of “every stationary set reflects”,

67 (1989), 353–366

Mekler, A. H., Shelah, S. and Spinas, O., The essentially free spectrum of a variety,

93 (1996), 1–8

Mekler, A. H. see Eklof, P. C. Mekler, A. H. see Komjáth, P. Melkman, A. A., Interpolation by splines satisfying mixed boundary conditions, 19 (1974),

369–381

Melkman, A. A., The Budan-Fourier theorem for splines, 19 (1974), 256–263 Mena-Jurado, J. F., Paya-Albert, R. and Rodriguez-Palacios, A., Semisummands

and semiideals in Banach spaces, 51 (1985), 33–67

Mendes France, M. see Kamae, T. Menegazzo, F. and Tomkinson, M. J., Groups with trivial virtual automorphism group,

71 (1990), 297–308

Menke, K., Zur Approximation des transfiniten Durchmessers bei bis auf Ecken analytischen

geschlossenen Jordankurven, 17 (1974), 136–141

Mennicke, J. see Grunewald, F. Merkurjev, A. see Mammone, P. Merrill, K. D., Cohomology of step functions under irrational rotations, 52 (1985),

320–340

Merrill, K. see Baggett, L. Merris, R., Extensions of the Hadamard determinant theorem, 46 (1983), 301–304 Mertens, J.-F. and Zamir, S., The maximal variation of a bounded martingale, 27

(1977), 252–276

Merzbach, E. and Nualart, D., Different kinds of two-parameter martingales, 52 (1985),

193–208

Meyer, J.-P., Mappings of bar constructions, 48 (1984), 331–339 Meyer, M., Une caracterisation volumique de certains espaces normes de dimension finie,

55 (1986), 317–326

Meyer, M. see Gluskin, E. D. Meyer, W., Maximum face size in an arrangement of curves, 14 (1973), 88–91 Meyer, W., On ordinary points in arrangements, 17 (1974), 124–135 Meyers, N. G., Continuity of Bessel potentials, 11 (1972), 271–283 Micchelli, C. A. and Rivlin, T. J., Numerical integration rules near Gaussian quadrature,

16 (1973), 287–299

Micchelli, C. A. see Karlin, S. Michael, E., A note on closed maps and compact sets, 2 (1964), 173–176 Michael, E. and Pełczyński, A., Separable Banach spaces which admit l_{n}^{∞} approxima-

tions, 4 (1966), 189–198

Michael, E. see Burke, D. Michor, P. W. see Gil-Medrano, O. Miles, R. E., A simple derivation of a formula of Furstenberg and Tzkoni, 14 (1973),

278–280

Milgram, M., The nonadditivity of the genus, 19 (1974), 201–207 Militaru, G. see Caenepeel, S. Miller, A. W. see Manevitz, L. M. Miller, C., An extension of Hölder’s theorem on the Gamma function, 97 (1997), 183–187 Miller, C. see van den Dries, L.

Miller, H. see Dror Farjoun, E. Miller, H. see Dwyer, W. G. Miller, R. W. and Turnidge, D. R., Co-Artinian rings and Morita duality, 15 (1973),

12–26

Millet, A. and Sucheston, L., On the existence of σ-finite invariant measures for

operators, 33 (1979), 349–367

Mills, T. M. and Varma, A. K., A new proof of A. F. Timan’s approximation theorem,

18 (1974), 39–44

Milman, P. D. see Bierstone, E. Milman, V. D., Some remarks about embeddings of l_{1}^{k} in finite-dimensional spaces, 43

(1982), 129–138

Milman, V. D. and Pisier, G., Banach spaces with a weak cotype 2 property, 54 (1986),

139–158

Milman, V. D. and Sharir, M., A new proof of the Maurey–Pisier theorem, 33 (1979),

73–87

Milman, V. D. and Wolfson, H., Minkowski spaces with extremal distance from the

Euclidean space, 29 (1978), 113–131

Milman, V. D. see Alon, N. Milman, V. D. see Amir, D. Milman, V. D. see Bourgain, J. Milman, V. D. see Davis, W. J. Mimura, M. see Maruyama, K. Minc, H., (0, 1)-matrices with minimal permanents, 15 (1973), 27–30 Minchev, E. see D. Bainov Minda, D. see Harmelin, R. Minda, D. see Ma, W. Miniowitz, R., Spherically mean p-valent quasiregular mappings, 38 (1981), 199–208 Miquel, V. see Giménez, F. Miquel, V. see Lluch, A. Misiurewicz, M. see Aaronson, J. Mitchem, J., Uniquely k-arborable graphs, 10 (1971), 17–25 Mitidieri, E. see Clément, P. Mityagin, B. S. see Aharoni, I. Miyazaki, Y., The eigenvalue distribution for a degenerate elliptic operator

−div{(1 − |x|^{2})grad·}, 71 (1990), 17–27

Mizohata, S., On evolution equations with finite propagation speed, 13 (1972), 173–187 Mizuki, M., Analyses of subensembles of M/G/1 queueing sequences, 3 (1965), 236–247 Mochizuki, K. and Suzuki, R., Critical exponent and critical blow-up for quasilinear

parabolic equations, 98 (1997), 141–156

Mohar, B. see Alon, N. Mollard, M. see Bezdek, K. Monk, J. D., A very rigid Boolean algebra, 35 (1980), 135–150 Montague, J. S. and Plemmons, R. J., Doubly stochastic matrix equations, 15 (1973),

216–229

Monterde, J. and Sánchez-Valenzuela, O. A., Graded metrics adapted to splittings,

99 (1997), 231–270

Monterde, J. and Sánchez-Valenzuela, O. A., The exterior derivative as a Killing

vector field, 93 (1996), 157–170

Montgomery, S., Biinvertible actions of Hopf algebras, 83 (1993), 45–71 Montgomery, S., Centralizers satisfying polynomial identities, 18 (1974), 207–219 Montgomery, S. and Passman, D. S., Crossed products over prime rings, 31 (1978),

224–256

Montgomery, S. and Passman, D. S., X-inner automorphisms of crossed products and semi-invariants of Hopf algebras, 55 (1986), 33–57 Montgomery, S. and Smith, S. P., Skew derivations and U_{q}(sl(2)), 72 (1990), 158–166 Montgomery, S. see Bahturin, Y. Montgomery, S. see Bergen, J. Montgomery-Smith, S. J., The Gaussian cotype of operators from C(K), 68 (1989),

123–128

Monticino, M. G. see Mauldin, R. D. Moody, R. V. see Berman, S. Moon, J. W. and Moser, L., On cliques in graphs, 3 (1965), 23–28 Moon, J. W. and Moser, L., On Hamiltonian bipartite graphs, 1 (1963), 163–165 Moore, C. C., Amenable subgroups of semi-simple groups and proximal flows, 34 (1979),

121–138

Moore, J. C. and Neisendorfer, J. A., Equivalence of Toda–Hopf invariants, 66 (1989),

300–318

Moran, G., Existence of nondetermined sets for some two person games over reals, 9

(1971), 316–329

Moran, G., Order-two continuous Hausdorff images of compact ordinals, 42 (1982), 75–86 Moran, G., Size direction games over the real line. II, 14 (1973), 418–441 Moran, G., The products of conjugacy classes in some infinite simple groups, 50 (1985),

54–74

Moran, G., Trees and the bireflection property, 41 (1982), 244–260 Moran, G. and Shelah, S., Size direction games over the real line. III, 14 (1973), 442–449 Moran, G. see Ehrenfeucht, A. Moran, G. see Magidor, M. Moran, L. E. and Tench, R. N., Normal complements in mod p-envelopes, 27 (1977),

331–338

Moran, W. and Pollington, A. D., The discrimination theorem for normality to

non-integer bases, 100 (1997), 339–347

Morandi, P. J. and Wadsworth, A. R., Baer orderings with noninvariant valuation

ring, 68 (1989), 241–255

Moriah, Y. and Shpilrain, V., Non-tame automorphisms of extensions of periodic groups,

84 (1993), 17–31

Moriguti, S. see Chow, Y. S. Morley, M. D., Countable models of ℵ_{1}-categorical theories, 5 (1967), 65–72 Morley, M. D., Decidable models, 25 (1976), 233–240 Morley, M. D. see Keisler, H. J. Mormul, P. and Zhitomirskii, M., Modules of vector fields, differential forms and

degenerations of differential systems, 95 (1996), 411–428

Moroz, B. Z., Estimates for character sums in number fields, 60 (1987), 1–21 Morrisey, T. J., A degenerate Hartman theorem, 95 (1996), 157–167 Morrow, J. R., On perfectly homogeneous bases in F -spaces, 29 (1978), 199–204 Mosak, R. D. and Moskowitz, M., Analytic density in Lie groups, 58 (1987), 1–9 Mosak, R. D. and Moskowitz, M., Zariski density in Lie groups, 52 (1985), 1–14 Moschovakis, Y. N. see Kechris, A. S. Moscovici, H., A reciprocity theorem for unitary representations of Lie groups, 15 (1973),

230–236

Moser, L. see Moon, J. W. Moskovitz, H. see Fishman, K. Moskowitz, M. see Mosak, R. D. Mossino, J., Application des inéquations quasi-variationnelles à quelques problèmes non

linéaires de la physique des plasmas, 30 (1978), 14–50

Mostow, G. D., Abraham Robinson, 1918–1974, 25 (1976), 4–14 Mountford, T. S., A solution to a problem of Dubins and Savage, 83 (1993), 253–256 Mozes, S., Actions of Cartan subgroups, 90 (1995), 253–294 Mozes, S., On closures of orbits and arithmetic of quaternions, 86 (1994), 195–209 Muhly, P. S., Isometries of ergodic Hardy spaces, 36 (1980), 50–74 Muhly, P. S. and Solel, B., On triangular subalgebras of groupoid C^{∗}-algebras, 71

(1990), 257–273

Müller, P., Reducibility behavior of polynomials with varying coefficients, 94 (1996),

59–91

Müller, P. F. X., A local version of a result of Gamlen and Gaudet, 63 (1988), 212–222 Müller, P. F. X., Classification of the isomorphic types of martingale-H^{1} spaces, 59

(1987), 195–212

Müller, P. F. X., On the span of some three valued martingale difference sequences in

L^{p} (1 < p < ∞) and H^{1}, 60 (1987), 39–53

Murata, M. and Shibata, Y., Lower bounds at infinity of solutions of partial differential

equations in the exterior of a proper cone, 31 (1978), 193–203

Murthy, M. K. V. and Stampacchia, G., A variational inequality with mixed boundary

conditions, 13 (1972), 188–224

Musson, I. M., Links between cofinite prime ideals in quantum function algebras, 100

(1997), 285–308

Myers, D., The Boolean algebra of the theory of linear orders, 35 (1980), 234–256 Myshkis, A. see D. Bainov Näätänen, M., Dilatation estimates for quasiconformal extensions, 29 (1978), 346–356 Nachbin, L. and Dineen, S., Entire functions of exponential type bounded on the real

axis and Fourier transforms of distributions with bounded supports, 13 (1972), 321–326

Nadel, M. E., An application of set theory to model theory, 11 (1972), 386–393 Nadel, M. E., More Lowenheim-Skolem results for admissible sets, 18 (1974), 53–64 Nadel, M. E., Some Lowenheim-Skolem results for admissible sets, 12 (1972), 427–432 Nadel, M. E., The completeness of Peano multiplication, 39 (1981), 225–233 Nadler, Jr., S. B. see Covitz, H. Nagel, R. and Räbiger, F., Superstable operators on Banach spaces, 81 (1993), 213–226 Nair, R., On LeVeque’s theorem about the uniform distribution (mod 1) of (a_{j} cos a_{j}x)_{j=1}^{∞},

65 (1989), 96–112

Nakai, M. see Glasner, M. Nakamura, H. see Asada, M. Nakao, M., Decay of solutions of the wave equation with a local degenerate dissipation,

95 (1996), 25–42

Namioka, I., Neighborhoods of extreme points, 5 (1967), 145–152 Namioka, I. and Pol, R., Mappings of Baire spaces into function spaces and Kadeč

renorming, 78 (1992), 1–20

Namioka, I. see Preiss, D. Nathanson, M. B., Multiplicative representations of integers, 57 (1987), 129–136 Nauwelaerts, E. and Van Oystaeyen, F., Zariski extensions and biregular rings, 37

(1980), 315–326

Navarro, G. see Iranzo, M. J. Navarro-Pascual, J. C., Extreme points and retractions in Banach spaces, 99 (1997),

335–342

Nebbia, C., Classification of all irreducible unitary representations of the stabilizer of the

horicycles of a tree, 70 (1990), 343–351

Ne’eman, Y. see Brustein, R. van Neerven, J. M. A. M. see Greiner, G. Negrepontis, S. and Tsarpalias, A., A non-linear version of the Amir–Lindenstrauss

method, 38 (1981), 82–94

Negrepontis, S. see Argyros, S. Neisendorfer, J. A. see Dwyer, W. G. Neisendorfer, J. A. see Moore, J. C.

Nerurkar, M. G. and Sussmann, H. J., Construction of minimal cocycles arising from

specific differential equations, 100 (1997), 309–326

Netanyahu, E. and Schiffer, M. M., On the monotonicity of some functionals in the

family of univalent functions, 32 (1979), 14–26

Netanyahu, E. see Erdős, P. Netanyahu, E. see Hanani, H. Neuberger, J. W., Tensor products and successive approximations for partial differential

equations, 6 (1968), 121–132

Neuwirth, J. H., Ergodicity of some mappings of the circle and the line, 31 (1978),

359–367

Nevanlinna, O. and Reich, S., Strong convergence of contraction semigroups and of

iterative methods for accretive operators in Banach spaces, 32 (1979), 44–58

Neveu, J., The filling scheme and the Chacon–Ornstein theorem, 33 (1979), 368–377 Nevo, A., A structure theorem for boundary-transitive graphs with infinitely many ends,

75 (1991), 1–19

Newelski, L., Free sets for some special set mappings, 58 (1987), 213–224 Newelski, L., Independence results for uncountable superstable theories, 65 (1989), 59–78 Newhouse, S. E. see Kifer, Y. Newman, C. M. see Bennett, G. Newman, D. J. see Bak, J. Newman, D. J. see Beck, A. Newman, M., Invariant factors of combinatorial matrices, 10 (1971), 126–130 Newton, D. see Keynes, H. B. Neyman, A., Decomposition of ranges of vector measures, 40 (1981), 54–64 Neyman, A., Representation of L_{p}-norms and isometric embedding in L_{p}-spaces, 48 (1984),

129–138

Neyman, A. see Kohlberg, E. Nicolau, M. and Reventós, A., On some geometrical properties of Seifert bundles, 47

(1984), 323–334

Nielsen, M. J., On two questions concerning tilings, 81 (1993), 129–143 Nielsen, N. J., On the Orlicz function spaces L_{M}(0, ∞), 20 (1975), 237–259 Nielsen, N. J., The positive approximation property of Banach lattices, 62 (1988), 99–112 Nielsen, N. J. see Mankiewicz, P. Niimura, M., Cluster sets in the large of meromorphic functions on Riemann surfaces, 37

(1980), 327–330

Niimura, M., The boundary behavior of functions meromorphic in bounded plane regions,

43 (1982), 161–165

Nishiwada, K. and Tintarev, K., Lacunas and saturation properties for hyperbolic

differential operators with analytic coefficients, 49 (1984), 281–306

Nissenzweig, A., w^{∗} sequential convergence, 22 (1975), 266–272 Nivat, M. see Cochet, Y. Nogueira, A., The three-dimensional Poincaré continued fraction algorithm, 90 (1995),

373–401

Nohel, J. A. see Crandall, M. G. Norberg, T., On the existence of ordered couplings of random sets—with applications,

77 (1992), 241–264

Notbohm, D., Kernels of maps between classifying spaces, 87 (1994), 243–256 Nowicki, T. see Meester, R. W. J. Nualart, D. see Merzbach, E. Nussbaum, R. D., Estimates of the periods of periodic points for nonexpansive operators,

76 (1991), 345–380

Obaya, R. see Alonso, A. I. Oberlin, D. M., M_{p}(G) ≠ M_{q}(G) (p^{−1}+ q^{−1} = 1), 22 (1975), 175–179 Oberlin, D. M., Singular measures and l^{p} Fourier transforms, 34 (1979), 115–120 Oberlin, D. M., The size of sums of sets, II, 55 (1986), 305–316 O’Brien, G. L., The road-colouring problem, 39 (1981), 145–154 Odell, E., A nonseparable Banach space not containing a subsymmetric basic sequence,

52 (1985), 97–109

Odell, E., On complemented subspaces of (∑ l_{2})_{l_{p}}, 23 (1976), 353–367 Odell, E. and Rosenthal, H. P., A double-dual characterization of separable Banach

spaces containing l^{1}, 20 (1975), 375–384

Odell, E., Rosenthal, H. P. and Schlumprecht, Th., On weakly null FDD’s in

Banach spaces, 84 (1993), 333–351

Odell, E. see Haydon, R. Odell, E. see Knaust, H. Okazawa, N. see Takenaka, T. Olsen, G. H. see Levin, M. Olsen, J., Dominated estimates of convex combinations of commuting isometries, 11

(1972), 1–13

Olsen, J. see Jones, R. L. Olsson, J. B. and Regev, A., On the T -ideal generated by a standard identity, 26 (1977),

97–104

Olszak, Z., On the existence of generalized complex space forms, 65 (1989), 214–218 Opolka, H., Cyclotomic splitting fields and strict cohomological dimension, 52 (1985),

225–230

Oren, I., Admissible functions with multiple discontinuities, 42 (1982), 353–360 Oren, I., Ergodicity of cylinder flows arising from irregularities of distribution, 44 (1983),

127–138

Orey, S. see Jain, N. C. Orlitsky, A. see Doyle, P. G. Orlov, V. P., Local solvability of one-dimensional problem of thermoviscoelasticity, 78

(1992), 51–64

Ornelas, A. see Colombo, G. Ørno, P., On Banach lattices of operators, 19 (1974), 264–265 Ornstein, D. S., Factors of Bernoulli shifts, 21 (1975), 145–153 Ornstein, D. S., Guessing the next output of a stationary process, 30 (1978), 292–296 Ornstein, D. S. and Smorodinsky, M., Continuous speed changes for flows, 31 (1978),

161–168

Ornstein, D. S. and Smorodinsky, M., Ergodic flows of positive entropy can be time

changed to become K-flows, 26 (1977), 75–83

Ornstein, D. S. and Weiss, B., Every transformation is bilaterally deterministic, 21

(1975), 154–158

Ornstein, D. S. and Weiss, B., Finitely determined implies very weak Bernoulli, 17

(1974), 94–104

Ornstein, D. S. and Weiss, B., Geodesic flows are Bernoullian, 14 (1973), 184–198 Ornstein, D. S. and Weiss, B., Subsequence ergodic theorems for amenable groups, 79

(1992), 113–127

Ornstein, D. S. and Weiss, B., The Shannon–McMillan–Breiman theorem for a class of

amenable groups, 44 (1983), 53–60

Ornstein, D. S. and Weiss, B., Unilateral codings of Bernoulli systems, 21 (1975),

159–166

Ornstein, D. S. see Arnoux, P. Orr, K. see Dror Farjoun, E. Osgood, B. see Chuaqui, M. Osterburg, J., A semiprime Morita context related to finite automorphism groups of

rings, 44 (1983), 88–96

Osterburg, J. and Passman, D. S., Computing the Connes spectrum of a Hopf algebra, 80 (1992), 225–253 Pach, J. see Goodman, J. E. Pach, J. see Komjáth, P. Pachner, U. see Kleinschmidt, P. Pajor, A., Quotient volumique et espaces de Banach de type 2 faible, 57 (1987), 101–106 Pajor, A. see Gluskin, E. D. Pálfy, P. P. see Freud, R. Palmon, O., The only convex body with extremal distance from the ball is the simplex,

80 (1992), 337–349

Pandey, B. S., On the harmonic summability of Laguerre series, 7 (1969), 299–302 Pansu, P. and Zimmer, R. J., Rigidity of locally homogeneous metrics of negative

curvature on the leaves of a foliation, 68 (1989), 56–62

Pant, H. see Lehrer, E. Papantoniou, B. J. see Blair, D. E. Paques, O. T. W. see Colombeau, J. F. Park, J. K. see Armendariz, E. P. Park, K. K., A special family of ergodic flows and their d̄-limits, 42 (1982), 343–352 Parker, D. B. see Hill, W. M. Parmenter, M. M. see Goodaire, E. G. Parry, W., A sharper Tits alternative for 3-manifold groups, 77 (1992), 265–271 Parry, W., An analogue of the prime number theorem for closed orbits of shifts of finite

type and their suspensions, 45 (1983), 41–52

Parry, W., Cohomology of permutative cellular automata, 99 (1997), 315–333 Parry, W., Endomorphisms of a Lebesgue space III, 21 (1975), 167–172 Parry, W., Notes on a posthumous paper by F. Hahn, 16 (1973), 38–45 Parry, W., The classification of topological Markov chains: Adapted shift equivalence, 38

(1981), 335–344

Parry, W., The information cocycle and ε-bounded codes, 29 (1978), 205–220 Parry, W. see Coelho, Z. Partington, J. R., Equivalent norms on spaces of bounded functions, 35 (1980), 205–209 Partington, J. R., Norm attaining operators, 43 (1982), 273–276 Pascuas, D. see Berenstein, C. A. Pask, D. A., Ergodicity of certain cylinder flows, 76 (1991), 129–152 Pask, D. A., Skew products over the irrational rotation, 69 (1990), 65–74 Passman, D. S., Advances in group rings, 19 (1974), 67–107 Passman, D. S., The semiprimitivity problem for group algebras of locally finite groups,

96 (1996), 481–509

Passman, D. S. see Bergen, J. Passman, D. S. see Lorenz, M. Passman, D. S. see Montgomery, S. Passman, D. S. see Osterburg, J. Paul, M. E. see Coven, E. M. Paul, M. E. see Markley, N. G. Paulin, F. see Gaboriau, D. Pavel, N. H., Semilinear equations with dissipative time-dependent domain perturbations,

46 (1983), 103–122

Pavel, N. H. and Iacob, F., Invariant sets for a class of perturbed differential equations

of retarded type, 28 (1977), 254–264

Pawlikowski, J., A characterization of strong measure zero sets, 93 (1996), 171–183 Payá, R., A counterexample on numerical radius attaining operators, 79 (1992), 83–101 Payá, R. see Acosta, M. D. Paya-Albert, R. see Mena-Jurado, J. F. Payne, L. E. and Sattinger, D. H., Saddle points and instability of nonlinear hyperbolic

equations, 22 (1975), 273–303 Paz, A., Graph-theoretic and algebraic characterizations of some Markov processes, 1

(1963), 169–180

Pazy, A., A class of semi-linear equations of evolution, 20 (1975), 23–36 Pazy, A., Asymptotic behavior of contractions in Hilbert space, 9 (1971), 235–240 Pazy, A., On the asymptotic behavior of iterates of nonexpansive mappings in Hilbert

space, 26 (1977), 197–204

Pazy, A. see Brezis, H. Pazy, A. see Crandall, M. G. Peck, N. T., Nearly exotic topologies on normed spaces, 10 (1971), 397–401 Peck, N. T., Twisted sums and a problem of Klee, 81 (1993), 357–368 Peck, N. T. and Talagrand, M., Banach sublattices of weak L_{1}, 59 (1987), 257–271 Peck, N. T. see Berg, I. D. Peetre, J. see Cobos, F. Peirats, V. see Hernández, F. L. Pełczyński, A., Some problems on bases in Banach and Fréchet spaces, 2 (1964), 132–138 Pełczyński, A. see Bessaga, C. Pełczyński, A. see Michael, E. Peleg, B., Bargaining sets of cooperative games without side payments, 1 (1963), 197–200 Peleg, B., Efficiency prices for optimal consumption plans. II, 9 (1971), 222–234 Peleg, B., Quota games with a continuum of players, 1 (1963), 48–53 Peleg, B., The kernel of the composition of characteristic function games, 3 (1965),

127–138

Peleg, R., Some extensions of weakly mixing flows, 9 (1971), 330–336 Peletier, L. A. see Kamin, S. Peng, L. see Jiang, Q. Pengra, R. W. see Casazza, P. G. Peres, Y., Application of Banach limits to the study of sets of integers, 62 (1988), 17–31 Peres, Y., Points of increase for random walks, 95 (1996), 341–347 Peres, Y. see Benjamini, I. Peres, Y. see Kenyon, R. Perets, G. S., d-critical modules of length 2 over Weyl algebras, 83 (1993), 361–368 Pérez Monasor, F. see Iranzo, M. J. Perez-Ramos, M. D., A characteristic subgroup of N -stable groups, 54 (1986), 51–59 Perkel, M., Groups of type F^{a,b,−c}, 52 (1985), 167–176 Perkins, E. see Evans, S. N. Perles, M. A., A proof of Dilworth’s decomposition theorem for partially ordered sets, 1

(1963), 105–107

Perles, M. A., On Dilworth’s theorem in the infinite case, 1 (1963), 108–109 Perles, M. A. and Shelah, S., A closed (n + 1)-convex set in R^{2} is a union of n^{6} convex

sets, 70 (1990), 305–312

Perles, M. A. see Altshuler, A. Perles, M. A. see Kojman, M. Perrizo, W. see Ellis, R. Pestov, V., On a valuation field invented by A. Robinson and certain structures connected

with it, 74 (1991), 65–79

Peters, J., A discrete analogue of a theorem of Katznelson, 37 (1980), 251–255 Peterson, F. P. see Cohen, F. R. Peterzil, Y. see Loveys, J. Petrovic, M. see Dillen, F. Pfeffer, W. F. and Prikry, K., A characterization of cardinals κ such that 2^{λ} = 2^{κ}

whenever κ ≤ λ < 2^{κ}, 65 (1989), 293–301

Pfluger, A. see Boutellier, R. Phelps, R. R., Some topological properties of support points of convex sets, 13 (1972),

327–336 Phelps, R. R., Weak^{∗} support points of convex sets in E^{∗}, 2 (1964), 177–182 Phelps, R. R. see Lindenstrauss, J. Phelps, R. R. see Preiss, D. Pianigiani, G., First return map and invariant measures, 35 (1980), 32–48 Piatetski-Shapiro, I. I. and Soudry, D., Special representations of rank one orthogonal

groups, 64 (1988), 276–314

Piatetski-Shapiro, I. I. see Gelbart, S. Picardello, M. A. see Casadio Tarabusi, E. Picardello, M. A. see Cohen, J. M. Picco, P. see Aaronson, J. Pierce, K. R. see Glass, A. M. W. Pierre, M., Un théorème général de génération de semi-groupes non linéaires, 23 (1976),

189–199

Pillay, A. see Hrushovski, E. Pinchuk, B., Functions of bounded boundary rotation, 10 (1971), 6–16 Pinchuk, B., Integral means of analytic functions, 17 (1974), 105–107 Pincus, D., Cardinal representatives, 18 (1974), 321–344 Pink, R. see Larsen, M. Pinkus, A., Representation theorems for Tchebycheffian polynomials with boundary

conditions and their applications, 17 (1974), 11–34

Pinsky, R., A probabilistic approach to a theorem of Gilbarg and Serrin, 74 (1991), 1–12 Pisier, G., A remarkable homogeneous Banach algebra, 34 (1979), 38–44 Pisier, G., Martingales with values in uniformly convex spaces, 20 (1975), 326–350 Pisier, G. see Milman, V. D. Pitt, L. D., Weighted L^{p} closure theorems for spaces of entire functions, 24 (1976), 94–118 Pixley, C. P., A note on the dimensions of projections of cells in E^{n}, 32 (1979), 117–123 Plant, A. T., The differentiability of nonlinear semigroups in uniformly convex spaces, 38

(1981), 257–268

Pleijel, Å., On a theorem by P. Malliavin, 1 (1963), 166–168 Plemmons, R. J. see Montague, J. S. Plotkin, B., Varieties of algebras and algebraic varieties, 96 (1996), 511–522 Plotkin, J. M., Correction to: “ZF and Boolean algebra” [Israel J. Math. 23 (1976), no.

3–4, 298–308], 24 (1976), 376–376

Plotkin, J. M., ZF and Boolean algebra, 23 (1976), 298–308 Pol, R. see Namioka, I. Polimeni, A. D. see Chartrand, G. Pollicott, M., A note on uniform distribution for primes and closed orbits, 55 (1986),

199–212

Pollicott, M., Asymptotic distribution of closed geodesics, 52 (1985), 209–224 Pollicott, M., C^{k}-rigidity for hyperbolic flows II, 69 (1990), 351–360 Pollicott, M., C^{r}-rigidity theorems for hyperbolic flows, 61 (1988), 14–28 Pollicott, M., One-dimensional maps via complex analysis in several variables, 91 (1995),

317–339

Pollicott, M., Zeta functions and analyticity of metric entropy for Anosov systems, 76

(1991), 257–263

Pollington, A. D. see Moran, W. Porcelli, P. see Dyer, J. A. Porta, H. see Berg, I. D. Posner, D. B. see Jockusch, C. G., Jr. Pouzet, M., Condition de chaîne en théorie des relations, 30 (1978), 65–84 Pouzet, M., Correction to: “Condition de chaîne en théorie des relations” [Israel J. Math.

30 (1978), no. 1-2, 65–84], 31 (1978), 395–395

Pouzet, M., Extensions complètes d’une théorie forcing complète, 16 (1973), 212–215 Power, S. C., On the Banach space isomorphism type of AF C^{∗}-algebras and their

triangular subalgebras, 94 (1996), 93–109

Prabha Arya, S. see Singal, M. K. Prat, U., On a weak sum theorem in dimension theory, 19 (1974), 277–283 Preiser, U., A characterization of PGL(2, q), q odd, 43 (1982), 177–184 Preiss, D., Phelps, R. R. and Namioka, I., Smooth Banach spaces, weak Asplund

spaces and monotone or usco mappings, 72 (1990), 257–279

Prestel, A., Algebraic number fields elementarily determined by their absolute Galois

group, 73 (1991), 199–205

Price, J. F., Local Sidon sets and uniform convergence of Fourier series, 17 (1974), 169–175 Pride, S. J. see Hill, P. Prikry, K. see Pfeffer, W. F. Procesi, C., Finite dimensional representations of algebras, 19 (1974), 169–182 Propp, J. G., A Shannon–McMillan theorem for motley names, 69 (1990), 225–234 Propp, J. G., Coding Markov chains from the past, 75 (1991), 289–328 Proppe, H., Byers, W. and Boyarsky, A., Singularity of topological conjugacies

between certain unimodal maps of the interval, 44 (1983), 277–288

Prüfer, F. see Berndt, J. Przytycki, F. see Levin, G. Puckette, E. E. see Lawler, G. F. Puczyłowski, E. R. see Ferrero, M. Purdy, G. B., Polytopes with prescribed contents of (n − 1)-facets, 30 (1978), 54–56 Purdy, G. B. see Entringer, R. C. Purves, R. A. and Sudderth, W. D., How to stay in a set or König’s lemma for random

paths, 43 (1982), 139–153

Purves, R. A. see Dubins, L. E. Pyber, L. see Lubotzky, A. Qian, N. and Zimmer, R. J., Entropy rigidity for semisimple group actions, 99 (1997),

55–67

Quas, A. N., A C^{1} expanding map of the circle which is not weak-mixing, 93 (1996),

359–372

Queffelec, M., Sur la singularité des produits de Riesz et des mesures spectrales associées

à la somme des chiffres, 34 (1979), 337–342

Quinn, D., Integral extensions of noncommutative rings, 73 (1991), 113–121 Quinto, E. T., Pompeiu transforms on geodesic spheres in real analytic manifolds, 84

(1993), 353–363

Räbiger, F. see Nagel, R. Rabin, M. O., Real time computation, 1 (1963), 203–211 Rabinowitz, P., On the convergence of interpolatory product integration rules based on

Gauss, Radau and Lobatto points, 56 (1986), 66–74

Rabinowitz, P. H., A note on topological degree theory for holomorphic maps, 16 (1973),

46–52

Radjavi, H. see Behzad, M. Radó, F., On mappings of the Galois space, 53 (1986), 217–230 Raghavan, S. see Dani, S. G. Raghunathan, M. S., A proof of Oseledec’s multiplicative ergodic theorem, 32 (1979),

356–362

Ragozin, D. L., Limits of generalized periodic D-splines, 54 (1986), 317–326 Rahe, M., Finite coding factors of Markov generators, 32 (1979), 349–355 Rahe, M. and Swanson, L., Loose block independence, 38 (1981), 100–106 Rahe, M. see Kieffer, J. C. Raianu, S. see Cohen, M. Raisonnier, J., A mathematical proof of S. Shelah’s theorem on the measure problem and

related results, 48 (1984), 48–56 Raisonnier, J. and Stern, J., The strength of measurability hypotheses, 50 (1985),

337–349

Raith, P., Continuity of the Hausdorff dimension for piecewise monotonic maps, 80 (1992),

97–133

Rajagopalan, M., A chain compact space which is not strongly scattered, 23 (1976),

117–125

Rallis, S. see Ginzburg, D. Rallis, S. see Kudla, S. S. Ram, J. see Bedi, S. S. Ramachandra Rao, A. and Rao, S. B., On the power sequence of a graph, 8 (1970),

398–402

Ramsay, A., Measurable group actions are essentially Borel actions, 51 (1985), 339–346 Ran, A., Embedding theorems for two-parameter groups of formal power series and related

problems, 9 (1971), 73–92

Rangaswamy, K. M. see Bican, L. Ransford, T. J., Norm inequalities for holomorphic semigroups, 97 (1997), 157–173 Rao, A. R., An extremal problem in graph theory, 6 (1968), 261–266 Rao, M., Capacitary inequalities for energy, 61 (1988), 179–191 Rao, M., On Green functions in R^{2}, 19 (1974), 313–328 Rao, M. and Šikić, H., Potential inequality, 83 (1993), 97–127 Rao, M. M., Extensions of the Hausdorff-Young theorem, 6 (1968), 133–149 Rao, N. V., A generalization of the Looman–Menchoff theorem, 70 (1990), 93–103 Rao, S. B. see Ramachandra Rao, A. Rao, T. S. S. R. K., Characterizations of some classes of L^{1}-preduals by the Alfsen–Effros

structure topology, 42 (1982), 20–32

Rapinchuk, A. S., Benyash-Krivetz, V. V. and Chernousov, V. I., Representation varieties of the fundamental groups of compact orientable surfaces, 93 (1996), 29–71

Ratner, M., Anosov flows with Gibbs measures are also Bernoullian, 17 (1974), 380–391 Ratner, M., Bernoulli flows over maps of the interval, 31 (1978), 298–314 Ratner, M., Horocycle flows are loosely Bernoulli, 31 (1978), 122–132 Ratner, M., Markov partitions for Anosov flows on n-dimensional manifolds, 15 (1973),

92–114

Ratner, M., Raghunathan’s conjectures for SL(2, R), 80 (1992), 1–31 Ratner, M., Some invariants of Kakutani equivalence, 38 (1981), 231–240 Ratner, M., The cartesian square of the horocycle flow is not loosely Bernoulli, 34 (1979),

72–96

Ratner, M., The central limit theorem for geodesic flows on n-dimensional manifolds of

negative curvature, 16 (1973), 181–197

Raugi, A. see Guivarc’h, Y. Rawat, R. and Sitaram, A., The injectivity of the Pompeiu transform and L^{p}-analogues

of the Wiener–Tauberian theorem, 91 (1995), 307–316

Raynaud, Y., Espaces de Banach superstables, distances stables et homéomorphismes

uniformes, 44 (1983), 33–52

Raynaud, Y., Finite representability of l_{p}(X) in Orlicz function spaces, 65 (1989), 197–213 Razar, M., Polynomial maps with constant Jacobian, 32 (1979), 97–106 Razon, A. see Jarden, M. Read, C. J., A Banach space with, up to equivalence, precisely two symmetric bases, 40

(1981), 33–53

Read, C. J., Extending an operator from a Hilbert space to a larger Hilbert space, so as

to reduce its spectrum, 57 (1987), 375–380

Read, C. J., The invariant subspace problem for a class of Banach spaces, 2: Hypercyclic

operators, 63 (1988), 1–40

Reay, J. R., Several generalizations of Tverberg’s theorem, 34 (1979), 238–244 Reay, J. R. see Bonnice, W. E. Rees, M., A point distal transformation of the torus, 32 (1979), 201–208 Rees, M., Non-conjugacy of a minimal distal diffeomorphism of the torus to a C^{1} skew-

product, 32 (1979), 193–200

Rees, M., Tangentially distal flows, 35 (1980), 9–31 Rees, R. see Vanstone, S. A. Regev, A., A note on Schur functions, 80 (1992), 155–160 Regev, A., A polynomial rate of growth for the multiplicities in cocharacters of matrices,

42 (1982), 65–74

Regev, A., Algebras satisfying a Capelli identity, 33 (1979), 149–154 Regev, A., Codimensions and trace codimensions of matrices are asymptotically equal, 47

(1984), 246–250

Regev, A., Existence of identities in A ⊗ B, 11 (1972), 131–152 Regev, A., On the height of the Kronecker product of S_{n} characters, 42 (1982), 60–64 Regev, A., On the identities of subalgebras of matrices over the Grassmann algebra, 58

(1987), 351–369

Regev, A., The T -ideal generated by the standard identity s_{3}[x_{1}, x_{2}, x_{3}], 26 (1977),

105–125

Regev, A. see Berele, A. Regev, A. see Drensky, V. Regev, A. see Klein, A. A. Regev, A. see Olsson, J. B. Reich, E., On extremal quasiconformal extensions of conformal mappings, 28 (1977), 91–97 Reich, E. and Strebel, K., On the Gerstenhaber–Rauch principle, 57 (1987), 89–100 Reich, J. I. see Blum, J. R. Reich, S. see Nevanlinna, O. Reichaw, M. see Blum, J. R. Reichaw-Reichbach, M., On compactification of metric spaces, 1 (1963), 61–74 Reichaw-Reichbach, M. see Hanani, H. Reichstein, Z. and Vonessen, N., Rational central simple algebras, 95 (1996), 253–280 Reid, K. B. and Thomassen, C., Edge sets contained in circuits, 24 (1976), 305–319 Reid, K. B. see Herzog, M. Reinhold, K. see Klemes, I. Reisner, S. see Gordon, Y. Rejtö, P. A., On reducing subspaces for one-electron Dirac operators, 9 (1971), 111–143 Rejtö, P. A., Second order concentration near the binding energy of the helium Schrodinger

operator, 6 (1968), 311–337

Rejtö, P. A., Some essentially self-adjoint one-electron Dirac operators, 9 (1971), 144–171 Rejtö, P. A. see Gustafson, K. E. Repický, M. see Judah, H. Resco, R., A dimension theorem for division rings, 35 (1980), 215–221 Resco, R., Transcendental division algebras and simple Noetherian rings, 32 (1979),

236–256

Ressayre, J.-P. see Harnik, V. Retherford, J. R., Operator characterizations of L_{p}-spaces, 13 (1972), 337–347 Retherford, J. R. see Gordon, Y. Reventós, A. see Nicolau, M. Revuz, D., Sur la définition des classes cycliques des chaînes de Harris, 33 (1979), 378–383 Revuz, D. see Brunel, A. Reznikov, A. G., Affine symplectic geometry I: Applications to geometric inequalities, 80

(1992), 207–224

Reznikov, A. G., Fundamental group of self-dual four-manifolds with positive scalar curvature, 97 (1997), 93–99 Ribe, M., Existence of separable uniformly homeomorphic nonisomorphic Banach spaces,

48 (1984), 139–147

Ribes, L., Virtually free factors of pro-p groups, 74 (1991), 337–346 Richards, F. B. and Schoenberg, I. J., Notes on spline functions IV: A cardinal spline

analogue of the theorem of the brothers Markov, 16 (1973), 94–102

Richter, L. J., On automorphisms of the degrees that preserve jumps, 32 (1979), 27–31 Rieders, E. see Ash, J. M. Riemenschneider, S. D., Convergence of interpolating cardinal splines: power growth,

23 (1976), 339–346

Rietz, R. E., A proof of the Grothendieck inequality, 19 (1974), 271–276 Rigoli, M., Salvatori, M. and Vignati, M., Subharmonic functions on graphs, 99

(1997), 1–27

Rigoli, M. see Ezin, J.-P. Riley, D. M. and Semple, J. F., Completion of restricted Lie algebras, 86 (1994), 277–299 Riley, D. M. and Semple, J. F., Power closure and the Engel condition, 97 (1997),

281–291

Riley, G. W., Approximations and the spectral properties of measure-preserving group

actions, 33 (1979), 9–31

Rinott, Y., Multivariate majorization and rearrangement inequalities with some applica-

tions to probability and statistics, 15 (1973), 60–77

del Rio, A. see Ara, P. Rips, E., An example of a non-LERF group which is a free product of LERF groups with

an amalgamated cyclic subgroup, 70 (1990), 104–110

Rips, E., Commutator equations in free groups, 39 (1981), 326–340 Rips, E., Generalized small cancellation theory and applications. I. The word problem, 41

(1982), 1–146

Rips, E., Note on primitive words, 39 (1981), 186–188 Rips, E., On a double of a free group, 96 (1996), 523–525 Rips, E., On the fourth integer dimension subgroup, 12 (1972), 342–346 Rips, E. see Gitik, R. Riskin, A., Type 3 diminimal maps on the torus, 75 (1991), 371–379 Risman, L. J., Cyclic algebras, complete fields, and crossed products, 28 (1977), 113–128 Ritter, J. see Jarden, M. Rivlin, T. J. see Micchelli, C. A. Robbins, H., Some complements to Brouwer’s fixed point theorem, 5 (1967), 225–226 Robbins, H. see Chow, Y. S. Robert, D., Sur les opérateurs linéaires qui transforment la boule unité d’un espace de

Banach en une partie latticiellement bornée d’un espace de Banach réticulé, 22 (1975), 354–360

Robert, D. see Lai, P. T. Roberts, B., The theta correspondence for similitudes, 94 (1996), 285–317 Robertson, A. G. see de la Harpe, P. Robinson, A. see Fisher, E. R. Robinson, D. J. S. see Aljadeff, E. Robinson, E. A., Jr., Transformations with highly nonhomogeneous spectrum of finite

multiplicity, 56 (1986), 75–88

Robinson, G. R. see Guralnick, R. M. Robinson, W. see Crone, L. Robson, J. C. see Agnarsson, G. Rodenhausen, H., A characterization of nonstandard liftings of measurable functions and

stochastic processes, 43 (1982), 1–22

Rodman, L. see Gohberg, I. Rodrígues-Romo, S. see Kharchenko, V. K. Rodríguez-Piazza, L. see Arias-de-Reyna, J.

Rodríguez-Salinas, B. see Hernández, F. L. Rodriguez-Palacios, A. see Mena-Jurado, J. F. Rogawski, J. D., Representations of GL(n) over a p-adic field with an Iwahori-fixed vector,

54 (1986), 242–256

Rogawski, J. see Gelbart, S. Roichman, Y., Decomposition of the conjugacy representation of the symmetric groups,

97 (1997), 305–316

Roitberg, J., Monomorphisms and epimorphisms in homotopy theory, 46 (1983), 205–211 Roitberg, J., Weak identities, phantom maps and H-spaces, 66 (1989), 319–329 Rolewicz, S., Plane sections of centrally symmetric convex bodies, 4 (1966), 135–138 Romanov, A. see Gol’dshtein, V. Romanovskii, N. S., On pro-p-groups with a single defining relator, 78 (1992), 65–73 Ron, A., Factorization theorems for univariate splines on regular grids, 70 (1990), 48–68 Ron, A. see Dyn, N. Rosa, A. see Vanstone, S. A. Rose, J. S., Universal finite group extensions and a non-splitting theorem, 15 (1973),

375–383

Rosen, J. D., ∗-Generalized polynomial identities of finite dimensional central simple

algebras, 46 (1983), 97–102

Rosenberg, J. see Howe, R. E. Rosenblatt, J. M., Equivalent invariant measures, 17 (1974), 261–270 Rosenfeld, M., Are all simple 4-polytopes Hamiltonian?, 46 (1983), 161–169 Rosenfeld, M., Independent sets in regular graphs, 2 (1964), 262–272 Rosenfeld, M., On the total coloring of certain graphs, 9 (1971), 396–402 Rosenfeld, M., Polytopes of constant weight, 21 (1975), 24–30 Rosenfeld, M. see Alspach, B. Rosenfeld, M. see Bollobás, B. Rosenfeld, M. see Dyer, J. A. Rosenmann, A., The normalized cyclomatic quotient associated with presentations of

finitely generated groups, 99 (1997), 285–313

Rosenthal, A., Strictly ergodic models and topological mixing for Z^{2}-action, 60 (1987),

31–38

Rosenthal, A., Strictly ergodic models for non-invertible transformations, 64 (1988),

57–72

Rosenthal, H. P., Correction to “On factors of C([0, 1]) with nonseparable dual” [Israel

Journal of Mathematics 13 (1972), 361–378], 21 (1975), 93–94

Rosenthal, H. P., On factors of C([0, 1]) with non-separable dual, 13 (1972), 361–378 Rosenthal, H. P., On the subspaces of L^{p} (p > 2) spanned by sequences of independent

random variables, 8 (1970), 273–303

Rosenthal, H. P. see Bourgain, J. Rosenthal, H. P. see Johnson, W. B. Rosenthal, H. P. see Lindenstrauss, J. Rosenthal, H. P. see Lotz, H. P. Rosenthal, H. P. see Odell, E. Rosenthal, P. see Feintuch, A. Rösler, U. see Kertz, R. P. Rosset, S., A new proof of the Amitsur-Levitski identity, 23 (1976), 187–188 Rosset, S., Remarks on a principle of localization, 39 (1981), 255–258 Rosset, S., The higher lower central series, 73 (1991), 257–279 Rosset, S. see Aljadeff, E. Rost, M. see Haile, D. E. Rostamian, R. see Alikakos, N. D. Rothaus, O. S., Study of the permanent conjecture and some of its generalizations, 18

(1974), 75–96 Rothblum, U. G., On orderable set functions and continuity I, 16 (1973), 375–397 Rothman, N. J., Remarks on duality and semigroups, 8 (1970), 83–89 Rothstein, A., Versik processes: first steps, 36 (1980), 205–224 Rottenberg, R. R., On finite sets of points in P^{3}, 10 (1971), 160–171 Rousseau, C. C. see Faudree, R. J. Rowen, L. H., Central simple algebras, 29 (1978), 285–301 Rowen, L. H., Correction to: “Cyclic division algebras” [Israel J. Math. 41 (1982),

213–234], 43 (1982), 277–280

Rowen, L. H., Correction to: “Monomial conditions on prime rings” [Israel J. Math. 27

(1977), no. 2, 131–149], 30 (1978), 192–192

Rowen, L. H., Cyclic division algebras, 41 (1982), 213–234 Rowen, L. H., Division algebra counterexamples of degree 8, 38 (1981), 51–57 Rowen, L. H., Examples of semiperfect rings, 65 (1989), 273–283 Rowen, L. H., Identities in algebras with involution, 20 (1975), 70–95 Rowen, L. H., Monomial conditions on prime rings, 27 (1977), 131–149 Rowen, L. H., Monomial conditions on rings, 23 (1976), 19–30 Rowen, L. H., Universal PI-algebras and algebras of generic matrices, 18 (1974), 65–74 Rowen, L. H. and Saltman, D. J., Prime to p extensions of division algebra, 78 (1992),

197–207

Rowen, L. H. and Saltman, D. J., Semidirect product division algebras, 96 (1996),

527–552

Rowen, L. H. and Tignol, J-P., On the decomposition of cyclic algebras, 96 (1996),

553–578

Rowen, L. H. see Amitsur, S. A. Rowley, P., Solubility of finite groups admitting a fixed-point-free automorphism of order

rst III, 51 (1985), 125–150

Roy, A. K., A short proof of a theorem on complex Lindenstrauss spaces, 38 (1981), 41–45 Roy, N. M., A characterization of square Banach spaces, 17 (1974), 142–148 Roy, N. M. see Cunningham, F., Jr. Rubin, B., The inversion of fractional integrals on a sphere, 79 (1992), 47–81 Rubin, J. E., Non-constructive properties of cardinal numbers, 10 (1971), 504–525 Rubin, L. R. see McCullough, D. Rubin, M., On the automorphism groups of countable Boolean algebras, 35 (1980), 151–170 Rubin, M., Theories of linear order, 17 (1974), 392–443 Rubinstein, Z. and Shaffer, D. B., On the multivalence and order of starlikeness of a

class of meromorphic functions, 17 (1974), 302–314

Rubshtein, B.-Z., Lacunary isomorphism of decreasing sequences of measurable partitions,

97 (1997), 317–345

Rudelson, M., Estimates of the weak distance between finite-dimensional Banach spaces,

89 (1995), 189–204

Rudnick, Z. see Sarnak, P. Rudolph, D. J., Counting the relatively finite factors of a Bernoulli shift, 30 (1978),

255–263

Rudolph, D. J., If a finite extension of a Bernoulli shift has no finite rotation factors, it is

Bernoulli, 30 (1978), 193–206

Rudolph, D. J., If a two-point extension of a Bernoulli shift has an ergodic square, then

it is Bernoulli, 30 (1978), 159–180

Rudolph, D. J., The second centralizer of a Bernoulli shift is just its powers, 29 (1978),

167–178

Rudolph, D. J., Two nonisomorphic K-automorphisms all of whose powers beyond one

are isomorphic, 27 (1977), 277–298

Rudolph, D. J., Two nonisomorphic K-automorphisms with isomorphic squares, 23

(1976), 274–287

Rudolph, D. J. and Schwarz, G., On attaining d̄, 24 (1976), 185–190 Rudolph, D. J. and Schwarz, G., The limits in d̄of multi-step Markov chains, 28 (1977),

103–109

Rudolph, D. J. see Assani, I. Rudolph, D. J. see del Junco, A. Rudolph, D. J. see Glasner, E. Rudolph, D. J. see Handel, M. Rudolph, D. J. see Iwanik, A. Rudolph, D. J. see Kwiatkowski, J. Ruess, W. M. and Stegall, C. P., Exposed and denting points in duals of operator

spaces, 53 (1986), 163–190

Ruess, W. M. and Summers, W. H., Weak almost periodicity and the strong ergodic

limit theorem for contraction semigroups, 64 (1988), 139–157

Rung, D. C. see Cima, J. A. Sagher, Y., On analytic families of operators, 7 (1969), 350–356 Salavessa, I. M. C. see Ezin, J.-P. Salce, L. see Bazzoni, S. Sallee, G. T., A non-continuous “Steiner point”, 10 (1971), 1–5 Sallee, G. T., Minkowski decomposition of convex sets, 12 (1972), 266–276 Sallee, G. T., Tiling convex sets by translates, 24 (1976), 368–376 Saltman, D. J., Retract rational fields and cyclic Galois extensions, 47 (1984), 165–215 Saltman, D. J. see Berele, A. Saltman, D. J. see Rowen, L. H. Salvatori, M. see Rigoli, M. de Sam Lazaro, J. see Lemańczyk, M. Samuel, E., The compound decision problem in the opponent case, 3 (1965), 117–126 Samuels, S. M. see Chow, Y. S. Sánchez-Valenzuela, O. A. see Monterde, J. Sanders, D. P. see Alon, N. Sangwine-Yager, J. R., The mean value of the area of polygons circumscribed about a

plane convex body, 37 (1980), 351–363

Sankey, A. D., Regular weights and the Johnson scheme, 97 (1997), 11–28 Sankey, A. D., Regular weights of full rank on strongly regular graphs, 95 (1996), 1–23 Santos, E. see Jara, P. Saphar, P., Applications p décomposantes et p absolument sommantes, 11 (1972), 164–179 Saphar, P., Hypothèse d’approximation á l’ordre p dans les espaces de Banach et

approximation d’applications p absolument sommantes, 13 (1972), 379–399

Saphar, P. see Feder, M. Sapir, M. V. see Margolis, S. W. Saracino, D., Existentially complete nilpotent groups, 25 (1976), 241–248 Sarkaria, K. S., A one-dimensional Whitney trick and Kuratowski’s graph planarity

criterion, 73 (1991), 79–89

Sarkaria, K. S., Tverberg’s theorem via number fields, 79 (1992), 317–320 Sárközy, A. see Erdős, P. Sarnak, P., Adams, S. and Rudnick, Z., Betti numbers of congruence groups, 88 (1994),

31–72

Sarnak, P. see Katok, S. Sastry, P. and Yekutieli, A., On residue complexes, dualizing sheaves and local

cohomology modules, 90 (1995), 325–348

Sattinger, D. H. see Payne, L. E. Saut, J. C. and Temam, R., Remarks on the Korteweg–de Vries equation, 24 (1976),

78–87

du Sautoy, M. P. F., Zeta functions of groups and rings: uniformity, 86 (1994), 1–23 Savin, G., Cusp forms, 80 (1992), 195–205 Saxl, J. see Fried, M. D.

Schacher, M. see Fein, B. Schachermayer, W., Norm attaining operators and renormings of Banach spaces, 44

(1983), 201–212

Schachermayer, W., The Banach–Saks property is not L^{2}-hereditary, 40 (1981), 340–344 Schachermayer, W., Sersouri, A. and Werner, E., Moduli of non-dentability and

the Radon–Nikodým property in Banach spaces, 65 (1989), 225–257

Schaefer, H. H., A minimax theorem for irreducible compact operators in L^{p}-spaces, 48

(1984), 196–204

Schaefer, H. H., Normed tensor products of Banach lattices, 13 (1972), 400–415 Schaefer, H. H., The zero-two law for positive contractions is valid in all Banach lattices,

59 (1987), 241–244

Schaer, J. and Wetzel, J. E., Boxes for curves of constant length, 12 (1972), 257–265 Schäffer, J. J., On the geometry of spheres in L-spaces, 10 (1971), 114–120 Schäffer, J. J., Sets with holes on the real line: A construction and an application, 4

(1966), 225–229

Schäffer, J. J., Symmetric curves, hexagons, and the girth of spheres in dimension 3, 6

(1968), 202–205

Schäffer, J. J. see James, R. C. Schappacher, W. see Desch, W. Schaps, M., Moduli of commutative and non-commutative finite covers, 58 (1987), 67–102 Schaps, M. see Fisman, E. Schechter, E., Correction to: “Perturbations of regularizing maximal monotone operators”

and a note on injectiveness, 47 (1984), 236–240

Schechter, E., Perturbations of regularizing maximal monotone operators, 43 (1982),

49–61

Schechter, M., The spectra of non-elliptic operators, 6 (1968), 384–397 Schechter, M. see Dash, A. T. Schechtman, G., A remark on unconditional basic sequences in L_{p} (1 < p < ∞), 19

(1974), 220–224

Schechtman, G., Complemented subspaces of (l_{2} ⊕ l_{2} ⊕ ⋯)_{p}(l < p < ∞) with an

unconditional basis, 20 (1975), 351–358

Schechtman, G., Examples of L_{p} spaces (1 < p ≠ 2 < ∞), 22 (1975), 138–147 Schechtman, G., On Pelczynski’s paper “Universal bases” , 22 (1975), 181–184 Schechtman, G., Random embeddings of Euclidean spaces in sequence spaces, 40 (1981),

187–192

Schechtman, G. see Johnson, W. B. Schein, B. M. see Freiman, G. A. Schellenberg, P. J. see Vanstone, S. A. Schelp, R. H. see Faudree, R. J. Schiffer, M. M. see Aharonov, D. Schiffer, M. M. see Netanyahu, E. Schild, G. see Brehm, U. Schindler, R.-D., On a Chang conjecture, 99 (1997), 221–230 Schirmer, H., Fixed point sets of homeomorphisms of compact surfaces, 10 (1971),

373–378

Schirmer, H., On fixed point sets of homeomorphisms of the n-ball, 7 (1969), 46–50 Schlumprecht, Th., An arbitrarily distortable Banach space, 76 (1991), 81–95 Schlumprecht, Th. see Odell, E. Schmeichel, E. F. see Hakimi, S. L. Schmerl, J. H., Models of Peano arithmetic and a question of Sikorski on ordered fields,

50 (1985), 145–159

Schmerl, J. H., Theories with only a finite number of existentially complete models, 28

(1977), 350–356

Schmid, J., On totally positive units of real holomorphy rings, 85 (1994), 339–350 Schmidt, K., Asymptotically invariant sequences and an action of SL(2, Z) on the 2-sphere,

37 (1980), 193–208

Schmidt, K., Hyperbolic structure preserving isomorphisms of Markov shifts, 55 (1986),

213–228

Schmidt, K., Hyperbolic structure preserving isomorphisms of Markov shifts. II, 58 (1987),

225–242

Schmidt, K., Invariant measures for certain expansive ℤ^{2}-actions, 90 (1995), 295–300 Schmidt, K., Spectra of ergodic group actions, 41 (1982), 151–153 Schmidt, K. see Lind, D. A. Schmitt, B. see Collet, P. Schmuckenschläger, M., The distribution function of the convolution square of a convex

symmetric body in ℝ^{n}, 78 (1992), 309–334

Schmutz, E., The order of a typical matrix with entries in a finite field, 91 (1995), 349–371 Schneider, H.-J., Principal homogeneous spaces for arbitrary Hopf algebras, 72 (1990),

167–195

Schneider, H.-J., Representation theory of Hopf Galois extensions, 72 (1990), 196–231 Schneider, H. see Hershkowitz, D. Schneider, R., On certain homomorphisms of restriction algebras of symmetric sets, 6

(1968), 223–232

Schneider, R., On Steiner points of convex bodies, 9 (1971), 241–249 Schneps, L., On Galois groups and their maximal 2-subgroups, 93 (1996), 125–144 Schober, G. see Baernstein, A., II Schoenberg, I. J., Norm inequalities for a certain class of C^{∞} functions, 10 (1971),

364–372

Schoenberg, I. J., Notes on spline functions III: On the convergence of the interpolating

cardinal splines as their degree tends to infinity, 16 (1973), 87–93

Schoenberg, I. J., The perfect B-splines and a time-optimal control problem, 10 (1971),

261–274

Schoenberg, I. J. see Richards, F. B. Schöneberg, R., Weak convergence of trajectories of nonexpansive semigroups in Hilbert

space, 30 (1978), 130–132

Schöneberg, R. see Kirk, W. A. Schönheim, J. and Burshtein, N., On a conjecture concerning exactly covering systems

of congruences, 8 (1970), 28–29

Schönheim, J. see Erdős, P. Schönheim, J. see Hanani, H. Schramm, O., Conformal uniformization and packings, 93 (1996), 399–428 Schramm, O., Existence and uniqueness of packings with specified combinatorics, 73

(1991), 321–341

Schramm, O., On the volume of sets having constant width, 63 (1988), 178–182 Schramm, O., Square tilings with prescribed combinatorics, 84 (1993), 97–118 Schreiber, B. M. see Borel, A. Schreiber, M., Dérivé convexe d’une loi de probabilité et lois stables. Application à un

problème d’isomorphisme, 28 (1977), 287–312

Schreiber, S., Some balanced complete block designs, 18 (1974), 31–37 Schreiber, S., Some mappings of periodic groups, 17 (1974), 117–123 Schuchert, P. see Altshuler, A. Schulz, C. see McMullen, P. Schuster, S. see Burns, D. Schuster, S. see Faudree, R. J. Schütt, C., A remark on the unconditional structure of L(E, F), 39 (1981), 16–24 Schütt, C., On the uniqueness of symmetric bases in finite dimensional Banach spaces,

40 (1981), 97–117

Schütt, C., The convex floating body and polyhedral approximation, 73 (1991), 65–77 Schütt, C., The isoperimetric quotient and some classical Banach spaces, 67 (1989), 43–61 Schütt, C., The projection constant of finite-dimensional spaces whose unconditional basis

constant is 1, 30 (1978), 207–212

Schütt, C., Unconditionality in tensor products, 31 (1978), 209–216 Schwartz, L. see Lannes, J. Schwarz, B., Bounds for the principal frequency of nonuniformly loaded strings, 1 (1963),

11–21

Schwarz, B., Carathéodory balls and norm balls of the domain H = {(z_{1}, z_{2}) ∈ C^{2}:

|z_{1}| + |z_{2}| < 1}, 84 (1993), 119–128

Schwarz, B., On the product on the distances of a point from the vertices of a polytope,

3 (1965), 29–38

Schwarz, B. see Edelstein, M. Schwarz, G., Randomizing when time is not well-ordered, 19 (1974), 241–245 Schwarz, G., Ways of randomizing and the problem of their equivalence, 17 (1974), 1–10 Schwarz, G. see Rudolph, D. J. Schweiger, F., Numbertheoretical endomorphisms with σ-finite invariant measure, 21

(1975), 308–318

Scoppola, C. M. and Shalev, A., Applications of dimension subgroups to the power

structure of p-groups, 73 (1991), 45–56

Scoppola, C. M. see Chillag, D. Scott, P. see Hass, J. Segal, D., A footnote on residually finite groups, 94 (1996), 1–5 Segal, D., Variations on polynomial subgroup growth, 94 (1996), 7–19 Segal, D. see Lubotzky, A. Segev, Y., Group actions on finite acyclic simplicial complexes, 82 (1993), 381–394 Segev, Y., On completely reducible solvable subgroups of GL(n, Δ), 51 (1985), 163–176 Segev, Y. see Aschbacher, M. Segovia, C. see Harboure, E. Seidman, T. I., An identity for normal-like operators, 7 (1969), 249–253 Seifter, N., Upper bounds for permanents of (1, −1)-matrices, 48 (1984), 69–78 Seifter, N. see Kräuter, A. R. Sela, Z., Dehn fillings that reduce Thurston norm, 69 (1990), 371–378 Sela, Z., Uniform embeddings of hyperbolic groups in Hilbert spaces, 80 (1992), 171–181 Semenov, A. M. see Agranovsky, M. L. Semple, J. F. see Riley, D. M. Series, C., Foliations of polynomial growth are hyperfinite, 34 (1979), 245–258 Series, C., Martin boundaries of random walks on Fuchsian groups, 44 (1983), 221–242 Series, C., The Rohlin tower theorem and hyperfiniteness for actions of continuous groups,

30 (1978), 99–122

Serre, J-P., Two letters on quaternions and modular forms (mod p), 95 (1996), 281–299 Sersouri, A. see Schachermayer, W. Sever, M., Hyperbolic systems of conservation laws with some special invariance properties,

75 (1991), 81–104

Sgro, J., Completeness theorems for continuous functions and product topologies, 25

(1976), 249–272

Shaffer, D. B. see Rubinstein, Z. Shafrir, I., The approximate fixed point property in Banach and hyperbolic spaces, 71

(1990), 211–223

Shalev, A., A new characterization of Frobenius complements, 87 (1994), 153–160 Shalev, A., Automorphisms of finite groups of bounded rank, 82 (1993), 395–404 Shalev, A., On associative algebras satisfying the Engel condition, 67 (1989), 287–290 Shalev, A., The nilpotency class of the unit group of a modular group algebra I, 70 (1990),

257–266

Shalev, A. see Lubotzky, A.

Shalev, A. see Mann, A. Shalev, A. see Scoppola, C. M. Shalhevet, J., On the minimal basis of a completely separating matrix, 2 (1964), 155–169 de Shalit, E., Differentials of the second kind on Mumford curves, 71 (1990), 1–16 Shamir, E., Regularization of mixed second-order elliptic problems, 6 (1968), 150–168 Shamir, E. and Upfal, E., On factors in random graphs, 39 (1981), 296–302 Shamir, E. see Bollobás, B. Shapiro, J. H., Examples of proper, closed, weakly dense subspaces in non locally convex

F -spaces, 7 (1969), 369–380

Shapiro, J. H., Subspaces of L^{p}(G) spanned by characters: 0 < p < 1, 29 (1978), 248–264 Shapiro, J. H. see Kalton, N. J. Shapiro, L. see Brezin, J. Shapiro, L. see Furstenberg, H. Sharir, M., A counterexample on extreme operators, 24 (1976), 320–328 Sharir, M., A non-nice extreme operator, 26 (1977), 306–312 Sharir, M., Characterization and properties of extreme operators into C(Y), 12 (1972),

174–183

Sharir, M. see Milman, V. D. Sharma, A., Some remarks on lacunary interpolation in the roots of unity, 2 (1964), 41–49 Shearer, M., Bifurcation in the neighbourhood of a non-isolated singular point, 30 (1978),

363–381

Shekhtman, B., On the norms of interpolating operators, 64 (1988), 39–48 Shelah, S., ℵ_{ω} may have a strong partition relation, 38 (1981), 283–288 Shelah, S., A Banach space with few operators, 30 (1978), 181–191 Shelah, S., A combinatorial principle and endomorphism rings I: On p-groups, 49 (1984),

239–257

Shelah, S., A compactness theorem for singular cardinals, free algebras, Whitehead

problem and transversals, 21 (1975), 319–349

Shelah, S., A consistent counterexample in the theory of collectionwise Hausdorff spaces,

65 (1989), 219–224

Shelah, S., A weak generalization of MA to higher cardinals, 30 (1978), 297–306 Shelah, S., Better quasi-orders for uncountable cardinals, 42 (1982), 177–226 Shelah, S., Can you take Solovay’s inaccessible away?, 48 (1984), 1–47 Shelah, S., Categoricity in ℵ_{1} of sentences in L_{ω_{1},ω}(Q), 20 (1975), 127–148 Shelah, S., Classification theory for non-elementary classes I: The number of uncountable

models of ψ ∈ L_{ω_{1},ω}. Part A, 46 (1983), 212–240

Shelah, S., Classification theory for non-elementary classes I: The number of uncountable

models of ψ ∈ L_{ω_{1},ω}. Part B, 46 (1983), 241–273

Shelah, S., Colouring without triangles and partition relation, 20 (1975), 1–12 Shelah, S., Constructions of many complicated uncountable structures and Boolean

algebras, 45 (1983), 100–146

Shelah, S., Decidability of a portion of the predicate calculus, 28 (1977), 32–44 Shelah, S., Decomposing topological spaces into two rigid homeomorphic subspaces, 63

(1988), 183–211

Shelah, S., Differentially closed fields, 16 (1973), 314–328 Shelah, S., Errata to: First order theory of permutation groups, 15 (1973), 437–441 Shelah, S., Every null-additive set is meager-additive, 89 (1995), 357–376 Shelah, S., Every two elementarily equivalent models have isomorphic ultrapowers, 10

(1971), 224–233

Shelah, S., First order theory of permutation groups, 14 (1973), 149–162 Shelah, S., For what filters is every reduced product saturated?, 12 (1972), 23–31 Shelah, S., Free limits of forcing and more on Aronszajn trees, 38 (1981), 315–334 Shelah, S., Further cardinal arithmetic, 95 (1996), 61–114 Shelah, S., “Gap 1” two-cardinal principles and the omitting types theorem for L(Q), 65

(1989), 133–152

Shelah, S., How special are Cohen and random forcings, i.e. Boolean algebras of the family

of subsets of reals modulo meagre or null, 88 (1994), 159–174

Shelah, S., Infinite abelian groups, Whitehead problem and some constructions, 18 (1974),

243–256

Shelah, S., Iterated forcing and changing cofinalities, 40 (1981), 1–32 Shelah, S., Iterated forcing and normal ideals on ω_{1}, 60 (1987), 345–380 Shelah, S., Jonsson algebras in successor cardinals, 30 (1978), 57–64 Shelah, S., Kaplansky test problem for R-modules, 74 (1991), 91–127 Shelah, S., Lifting problem of the measure algebra, 45 (1983), 90–96 Shelah, S., More on monadic logic. Part C: Monadically interpreting in stable unsuperstable

T and the monadic theory of ^{ω}λ, 70 (1990), 353–364

Shelah, S., More on monadic logic. Part D: A note on addition of theories, 68 (1989),

302–306

Shelah, S., More on powers of singular cardinals, 59 (1987), 299–326 Shelah, S., Multi-dimensionality, 74 (1991), 281–288 Shelah, S., Notes on combinatorial set theory, 14 (1973), 262–277 Shelah, S., Notes on monadic logic. Part A. Monadic theory of the real line, 63 (1988),

335–352

Shelah, S., Notes on monadic logic. Part B: Complexity of linear orders in ZFC, 69 (1990),

94–116

Shelah, S., Number of open sets for a topology with a countable basis, 83 (1993), 369–374 Shelah, S., On co-κ-Souslin relations, 47 (1984), 139–153 Shelah, S., On endo-rigid, strongly ℵ_{1}-free abelian groups in ℵ_{1}, 40 (1981), 291–295 Shelah, S., On languages with non-homogeneous strings of quantifiers, 8 (1970), 75–79 Shelah, S., On measure and category, 52 (1985), 110–114 Shelah, S., On reconstructing separable reduced p-groups with a given socle, 60 (1987),

146–166

Shelah, S., On uncountable abelian groups, 32 (1979), 311–330 Shelah, S., Products of regular cardinals and cardinal invariants of products of Boolean

algebras, 70 (1990), 129–187

Shelah, S., Refuting Ehrenfeucht conjecture on rigid models, 25 (1976), 273–286 Shelah, S., Stable theories, 7 (1969), 187–202 Shelah, S., Successors of singulars, cofinalities of reduced products of cardinals and

productivity of chain conditions, 62 (1988), 213–256

Shelah, S., The consistency of Ext(G, Z) = Q, 39 (1981), 74–82 Shelah, S., The number of non-isomorphic models of an unstable first-order theory, 9

(1971), 473–487

Shelah, S., The spectrum problem I: ℵ_{ε}-saturated models, the main gap, 43 (1982),

324–356

Shelah, S., The spectrum problem II: Totally transcendental and infinite depth, 43 (1982),

357–364

Shelah, S., The spectrum problem III: Universal theories, 55 (1986), 229–256 Shelah, S., The theorems of Beth and Craig in abstract model theory, III: Δ-logics and

infinitary logics, 69 (1990), 193–213

Shelah, S., There are just four second-order quantifiers, 15 (1973), 282–300 Shelah, S., Two cardinal compactness, 9 (1971), 193–198 Shelah, S., Uncountable constructions for B.A. e.c. groups and Banach spaces, 51 (1985),

273–297

Shelah, S., Universal graphs without instances of CH: revisited, 70 (1990), 69–81 Shelah, S., Vive la différence II. The Ax–Kochen isomorphism theorem, 85 (1994), 351–390 Shelah, S., Was Sierpinski right? I, 62 (1988), 355–380 Shelah, S., Whitehead groups may be not free, even assuming CH, I, 28 (1977), 193–204 Shelah, S., Whitehead groups may not be free even assuming CH, II, 35 (1980), 257–285 Shelah, S., Harrington, L. and Makkai, M., A proof of Vaught’s conjecture for

ω-stable theories, 49 (1984), 259–280

Shelah, S. and Hart, B., Categoricity over P for first order T or categoricity for φ ∈ L_{ω_{∞}ω}

can stop at ℵ_{k} while holding for ℵ_{0}, …, ℵ_{k−1}, 70 (1990), 219–235

Shelah, S. and Jin, R., A model in which there are Jech–Kunen trees but there are no

Kurepa trees, 84 (1993), 1–16

Shelah, S. and Soukup, L., On the number of non-isomorphic subgraphs, 86 (1994),

349–371

Shelah, S. and Stanley, L. J., Corrigendum to: “Generalized Martin’s axiom and

Souslin’s hypothesis for higher cardinals” , 53 (1986), 304–314

Shelah, S. and Stanley, L. J., Generalized Martin’s axiom and Souslin’s hypothesis for

higher cardinals, 43 (1982), 225–236

Shelah, S. and Stanley, L. J., More consistency results in partition calculus, 81 (1993),

97–110

Shelah, S. and Stanley, L. J., S-forcing, I. A “black-box” theorem for morasses, with

applications to super-Souslin trees, 43 (1982), 185–224

Shelah, S. and Stanley, L. J., S-forcing IIa: Adding diamonds and more applications:

coding sets, Arhangel’skii’s problem and L[Q_{1}^{<ω}, Q_{2}^{1}], 56 (1986), 1–65

Shelah, S. and Weiss, B., Measurable recurrence and quasi-invariant measures, 43 (1982),

154–160

Shelah, S. and Woodin, H., Large cardinals imply that every reasonably definable set of

reals is Lebesgue measurable, 70 (1990), 381–394

Shelah, S. see Abraham, U. Shelah, S. see Amit, R. Shelah, S. see Avraham, U. Shelah, S. see Ben David, S. Shelah, S. see Blass, A. Shelah, S. see Bonnet, R. Shelah, S. see Burke, M. R. Shelah, S. see Devlin, K. J. Shelah, S. see Dror Farjoun, E. Shelah, S. see Droste, M. Shelah, S. see Eklof, P. C. Shelah, S. see Erdős, P. Shelah, S. see Fremlin, D. H. Shelah, S. see Gitik, M. Shelah, S. see Grossberg, R. Shelah, S. see Gurevich, Y. Shelah, S. see Hiller, H. L. Shelah, S. see Hrushovski, E. Shelah, S. see Jech, T. J. Shelah, S. see Judah, H. Shelah, S. see Juhász, I. Shelah, S. see Kojman, M. Shelah, S. see Lachlan, A. H. Shelah, S. see Levinski, J.-P. Shelah, S. see Litman, A. Shelah, S. see Liu, K. Shelah, S. see Mekler, A. H. Shelah, S. see Moran, G. Shelah, S. see Perles, M. A. Shemer, I., How many cyclic subpolytopes can a non-cyclic polytope have?, 49 (1984),

331–342 Shemer, I., Neighborly polytopes, 43 (1982), 291–314 Shemer, I. see Altshuler, A. Shemer, I. see Bokowski, J. Shephard, G. C., A pre-Hilbert space consisting of classes of convex sets, 4 (1966), 1–10 Shephard, G. C., A theorem on cyclic polytopes, 6 (1968), 368–372 Shephard, G. C., Shadow systems of convex sets, 2 (1964), 229–236 Shephard, G. C., Spherical complexes and radial projections of polytopes, 9 (1971),

257–262

Shephard, G. C., Subpolytopes of stack polytopes, 19 (1974), 292–296 Shepp, L. A., Connectedness of certain random graphs, 67 (1989), 23–33 Shepp, L. A., Covering the circle with random arcs, 11 (1972), 328–345 Shepp, L. A. see Doyle, P. G. Sherman, S., Markov random fields and Gibbs random fields, 14 (1973), 92–103 Shibata, T., Nodal and asymptotic properties of solutions to nonlinear elliptic eigenvalue

problems on general level sets, 74 (1991), 225–240

Shibata, Y. see Murata, M. Shields, P. C. and Thouvenot, J.-P., Une remarque sur les isomorphismes de schémas

de Bernoulli qui préservent certains facteurs, 21 (1975), 173–176

Shields, P. C. see Adler, R. L. Shields, P. C. see Marton, K. Shiokawa, I. see Kano, H. Shmuely, Z., Some embeddings of infinite-dimensional spaces, 12 (1972), 5–10 Shnider, S. see Donin, J. Shock, R. C., Essentially nilpotent rings, 9 (1971), 180–185 Shock, R. C., The ring of endomorphisms of a finite dimensional module, 11 (1972),

309–314

Shore, R. A., The irregular and non-hyperregular α-r.e. degrees, 22 (1975), 28–41 Shore, R. A. see Abraham, U. Shore, R. A. see Downey, R. Shore, R. A. see Groszek, M. Shore, R. A. see Lerman, M. Shore, R. A. see Maass, W. Shpectorov, S. V. see Ivanov, A. A. Shpilrain, V., Test elements for endomorphisms of free groups and algebras, 92 (1995),

307–316

Shpilrain, V. see Moriah, Y. Shumyatsky, P., On periodic solvable groups having automorphisms with nilpotent fixed

point groups, 87 (1994), 111–116

Shustin, E., Super-high-frequency oscillations in a discontinuous dynamic system with

time delay, 90 (1995), 199–219

Sibony, M., Minimisation de fonctionnelles non differentiables, 8 (1970), 105–126 Siefkes, D., An axiom system for the weak monadic second order theory of two successors,

30 (1978), 264–284

Siejka, H., On meromorphic univalent functions omitting a disc, 54 (1986), 291–300 Sigmund, K., On the prevalence of zero entropy, 10 (1971), 281–288 Šikić, H. see Rao, M. Silva, C. E., On μ-recurrent nonsingular endomorphisms, 61 (1988), 1–13 Silver, D. S. and Williams, S. G., Augmented group systems and shifts of finite type,

95 (1996), 231–251

Simanca, S. R., A note on extremal metrics of non-constant scalar curvature, 78 (1992),

85–93

Simeonov, P. S. see Bainov, D. Simmons, H., Correction to: “Each regular number structure is biregular” [Israel J. Math.

23 (1976), no. 3-4, 347–352], 26 (1977), 95–95 Simmons, H., Each regular number structure is biregular, 23 (1976), 347–352 Simmons, H., The semiring of topologizing filters of a ring, 61 (1988), 271–284 Simmons, H. see Goldrei, D. C. Simmons, H. see Macintyre, A. Simon, P. see Balcar, B. Simon, R. S., Spież, S. and Toruńczyk, H., The existence of equilibria in certain games,

separation for families of convex functions and a theorem of Borsuk–Ulam type, 92 (1995), 1–21

Simpson, S. G. see Lerman, M. Sine, R. C., An example in the theory of stable Markov operators, 14 (1973), 283–284 Sine, R. C., Convergence theorems for weakly almost periodic Markov operators, 19 (1974),

246–255

Sine, R. C., On invariant probabilities for random rotations, 33 (1979), 384–388 Sinestrari, E. see Da Prato, G. Sinestrari, E. see Di Blasio, G. Singal, M. K. and Deb, M., On monotone and quasicompact mappings, 10 (1971),

218–223

Singal, M. K. and Mathur, A., A note on minimal topological spaces, 12 (1972), 204–206 Singal, M. K. and Prabha Arya, S., Two sum theorems for topological spaces, 8 (1970),

155–158

Singer, I., A proof of the Dvoretzky-Rogers theorem, 2 (1964), 249–250 Singer, I. see Davis, W. J. Singh, S. see Goel, V. K. Sitaram, A. see Rawat, R. Sjögren, P. and Soria, F., Weak type (1,1) estimates for some integral operators related

to rough maximal functions, 95 (1996), 211–229

Sjölin, P. and Strömberg, J.-O., Spline systems as bases in Hardy spaces, 45 (1983),

147–156

Skorobogatov, A. N., Exponential sums, the geometry of hyperplane sections, and some

Diophantine problems, 80 (1992), 359–379

Sloss, J. M., The shape of the sides of an arc, 1 (1963), 217–220 Small, L. W. and Stafford, J. T., Homological properties of generic matrix rings, 51

(1985), 27–32

Small, L. W. and Wadsworth, A. R., An example in affine PI-rings, 36 (1980), 285–286 Small, L. W. see Amitsur, S. A. Small, L. W. see Blair, W. D. Small, L. W. see Herstein, I. N. Small, L. W. see Joseph, A. Small, L. W. see Kirkman, E. E. Smilansky, Z., A non-geometric shelling of a 3-polytope, 71 (1990), 29–32 Smilansky, Z., Bi-cyclic 4-polytopes, 70 (1990), 82–92 Smilansky, Z., Convex hulls of generalized moment curves, 52 (1985), 115–128 Smith, B. see Erdős, P. Smith, J. H., Simplicial group models for Ω^{n}S^{n}X, 66 (1989), 330–350 Smith, P. F., More on the AR property and chain conditions in group rings, 35 (1980),

186–204

Smith, P. F., The AR property and chain conditions in group rings, 32 (1979), 131–144 Smith, P. F. and Snider, R. L., Group rings of hyperabelian groups, 38 (1981), 23–28 Smith, R. R. see Lewis, D. R. Smith, S. P., Central localization and Gelfand–Kirillov dimension, 46 (1983), 33–39 Smith, S. P. see Montgomery, S. Smith, W. W. see Chapman, S. T. Smits, L., Combinatorial approximation to the divergence of one-forms on surfaces, 75

(1991), 257–271 Smorodinsky, M., Block codes for Bernoulli shifts, 49 (1984), 325–330 Smorodinsky, M., Singular measures and Hausdorff measures, 7 (1969), 203–206 Smorodinsky, M. and Thouvenot, J.-P., Bernoulli factors that span a transformation,

32 (1979), 39–43

Smorodinsky, M. and Weiss, B., Normal sequences for Markov shifts and intrinsically

ergodic subshifts, 59 (1987), 225–233

Smorodinsky, M. see Burton, R. M. Smorodinsky, M. see Goldberger, J. Smorodinsky, M. see Keane, M. Smorodinsky, M. see Ornstein, D. S. Snider, R. L. see Smith, P. F. Soare, R. I. see Lerman, M. Soifer, G. A. see Abels, H. Solecki, S. see Kechris, A. S. Solel, B., Maximality of analytic operator algebras, 62 (1988), 63–89 Solel, B. see Muhly, P. S. Solomon, A., Some symmetric systems of minimal surfaces, 8 (1970), 65–74 Solomon, R. M. see Brozovic, D. P. Solomyak, M., Piecewise-polynomial approximation of functions from H^{ℓ}((0, 1)^{d}), 2ℓ = d,

and applications to the spectral theory of the Schrödinger operator, 86 (1994), 253–275

Solovay, R. M., Provability interpretations of modal logic, 25 (1976), 287–304 Solovay, R. M. see Jockusch, C. G., Jr. Solovay, R. M. see Lévy, A. Song, Y., On the relative Wall–Witt groups, 90 (1995), 189–197 Sonn, J., Brauer groups, embedding problems, and nilpotent groups as Galois groups, 85

(1994), 391–405

Sonn, J., Class groups and Brauer groups, 34 (1979), 97–105 Sonn, J., Correction to: “On equivalence of number fields” [Israel J. Math. 52 (1985), no.

3, 239–244], 71 (1990), 379–379

Sonn, J., Epimorphisms of Demushkin groups, 17 (1974), 176–190 Sonn, J., Frobenius Galois groups over quadratic fields, 31 (1978), 91–96 Sonn, J., On equivalence of number fields, 52 (1985), 239–244 Sonn, J., Rational division algebras as solvable crossed products, 37 (1980), 246–250 Sonn, J. see Chillag, D. Sonn, J. see Malle, G. Soria, F. see Sjögren, P. Sós, V. T. see Erdős, P. Soudry, D., A uniqueness theorem for representations of GSO(6) and the strong multiplicity

one theorem for generic representations of GSp(4), 58 (1987), 257–287

Soudry, D. see Gelbart, S. Soudry, D. see Ginzburg, D. Soudry, D. see Piatetski-Shapiro, I. I. Soukup, L. see Juhász, I. Soukup, L. see Shelah, S. Soulé, C. see Gillet, H. Spencer, J. H., On cliques in graphs, 9 (1971), 419–421 Spencer, J. see Alon, N. Spież, S. see Simon, R. S. Spinas, O., Iterated forcing in quadratic form theory, 79 (1992), 297–315 Spinas, O. see Mekler, A. H. Spivack, M., Derivations and nest algebras on Banach space, 50 (1985), 193–200 Srebro, U., Carathéodory balls and norm balls in H_{n} = {z ∈ ℂ^{n}: ∥z∥_{1} < 1}, 89 (1995),

61–70 Srebro, U., Conformal capacity and quasiconformal mappings in R̄^{n}, 9 (1971), 93–110 Srebro, U. see Martio, O. Srinivas Rau, S., A corollary to Kodaira–Spencer’s theorem on continuity of eigenvalues,

76 (1991), 229–235

Srinivasan, S., Two sufficient conditions for supersolvability of finite groups, 35 (1980),

210–214

Stafford, J. T., Dimensions of division rings, 45 (1983), 33–40 Stafford, J. T. see Small, L. W. Stampacchia, G. see Murthy, M. K. V. Stampfli, J. G. see Anderson, J. H. Stampfli, J. G. see Fan, P. Stanley, L. J. see Shelah, S. Starbird, T. see Benyamini, Y. Stavi, J., A model of ZF with an infinite free complete Boolean algebra, 20 (1975), 149–163 Stavi, J., On cardinal collapsing with reals, 18 (1974), 11–18 Stavi, J., On strongly and weakly defined Boolean terms, 15 (1973), 31–43 Stegall, C., A proof of the theorem of Amir and Lindenstrauss, 68 (1989), 185–192 Stegall, C., Characterizations of Banach spaces whose duals are L_{1} spaces, 11 (1972),

299–308

Stegall, C., The duality between Asplund spaces and spaces with the Radon–Nikodym

property, 29 (1978), 408–412

Stegall, C. P. see Ruess, W. M. Steif, J. E. see Burton, R. M. Stein, S. K., Large exclusive families, 6 (1968), 26–29 Stein, Y., The Jacobian problem as a system of ordinary differential equations, 89 (1995),

301–319

Stein, Y., The total reducibility order of a polynomial in two variables, 68 (1989), 109–122 Steinhorn, C. I. and King, J. H., The uniformization property for ℵ_{2}, 36 (1980), 248–256 Steitz, P. W., Upper bounds for growth in the Ryll–Nardzewski function of an ω-

categorical, ω-stable theory, 77 (1992), 335–343

Stellmacher, B., A characteristic subgroup of Σ_{4}-free groups, 94 (1996), 367–379 Steprāns, J., Martin’s axiom and the transitivity of P_{𝔠}-points, 83 (1993), 257–274 Steprāns, J. and Watson, W. S., Extending ideals, 54 (1986), 201–226 Stern, J., A Ramsey theorem for trees, with an application to Banach spaces, 29 (1978),

179–188

Stern, J., Some measure theoretic results in effective descriptive set theory, 20 (1975),

97–110

Stern, J., The problem of envelopes for Banach spaces, 24 (1976), 1–15 Stern, J. see Raisonnier, J. Sterna-Karwat, A., On existence of cone-maximal points in real topological linear spaces,

54 (1986), 33–41

Sternberg, S. see Brustein, R. Sternfeld, J., Dimension theory and superpositions of continuous functions, 20 (1975),

300–320

Sternfeld, Y., Dimension, superposition of functions and separation of points, in compact

metric spaces, 50 (1985), 13–53

Sternfeld, Y., Uniform separation of points and measures and representation by sums of

algebras, 55 (1986), 350–362

Sternfeld, Y., Uniformly separating families of functions, 29 (1978), 61–91 Sternfeld, Y. see Hansen, A. B. Sternfeld, Y. see Haydon, R. Sternfeld, Y. see Levin, M. Stewart, B. M. and Herzog, F., Semi-regular plane polygons of integral type, 11 (1972),

31–52 Stiles, W. J., On subseries convergence in F -spaces, 8 (1970), 53–56 Stinson, D. R. see Vanstone, S. A. Stob, M. see Downey, R. Stob, M. see Maass, W. Stöhr, R., Homology of free Lie powers and torsion in groups, 84 (1993), 65–87 Straus, E. G., Two comments on Dvoretzky’s sphericity theorem, 1 (1963), 221–223 Strebel, K. see Reich, E. Strömberg, J.-O. see Sjölin, P. Stroth, G., An odd characterization of J_{4}, 31 (1978), 189–192 Stroth, G., Strong quadratic modules, 79 (1992), 257–279 Struppa, D. C., Convolution equations and spaces of ultradifferentiable functions, 54

(1986), 60–70

Stuck, G., Correction to: “Low dimensional actions of semisimple groups” [Israel J. Math.

76 (1991), no. 1–2, 27–71], 80 (1992), 255–255

Stuck, G., Low dimensional actions of semisimple groups, 76 (1991), 27–71 Sturmfels, B. see Bokowski, J. Sucheston, L. see Akcoglu, M. A. Sucheston, L. see Millet, A. Sudderth, W. D. see Dubins, L. E. Sudderth, W. D. see Maitra, A. Sudderth, W. D. see Purves, R. A. Sukochev, F. A. see Chilin, V. I. Summers, W. H. see Ruess, W. M. Sun, Z.-W., On exactly m times covers, 77 (1992), 345–348 Sundaresan, K., Uniformly non-l_{n}^{(1)} Orlicz spaces, 3 (1965), 139–146 Sundberg, C. see Casazza, P. G. Suryanarayan, E. R., Intrinsic equations of rotational gas flows, 5 (1967), 118–126 Sussmann, H. J. see Nerurkar, M. G. Suzuki, R. see Mochizuki, K. Svensson, S. L., Singular differential operators and distributions, 38 (1981), 131–153 Swanson, L. see Rahe, M. Swart, J. see Jarchow, H. Swarup, G. see Maskit, B. Sweet, W. J., Jr. see Kudla, S. S. Szankowski, A., A Banach lattice without the approximation property, 24 (1976), 329–337 Szankowski, A., An example of a universal Banach space, 11 (1972), 292–296 Szankowski, A., Embedding Banach spaces with unconditional bases into spaces with

symmetric bases, 15 (1973), 53–59

Szankowski, A., On Dvoretzky’s theorem on almost spherical sections of convex bodies,

17 (1974), 325–338

Szankowski, A., On the uniform approximation property in Banach spaces, 49 (1984),

343–359

Szankowski, A., Subspaces without the approximation property, 30 (1978), 123–129 Szankowski, A. and Terenzi, P., Independent sequences in Banach spaces, 41 (1982),

147–150

Szankowski, A. and Zippin, M., A stability property of the unit vector basis of l_{p}, 20

(1975), 216–227

Szarek, S. J. see Bourgain, J. Szarek, S. J. see Dilworth, S. J. Szüsz, P., Remark on a theorem of Aharonov and Walsh, 17 (1974), 108–110 Taft, E. J. see Larson, R. G. Takahashi, S. see Dumont, J. M.

Takenaka, T. and Okazawa, N., Wellposedness of abstract Cauchy problems for second

order differential equations, 69 (1990), 257–288

Takeuchi, M., Matric bialgebras and quantum groups, 72 (1990), 232–251 Talagrand, M., A conjecture on convolution operators, and a non-Dunford–Pettis operator

on L^{1}, 68 (1989), 82–88

Talagrand, M., A new countably determined Banach space, 47 (1984), 75–80 Talagrand, M., Certaines formes linéaires pathologiques sur un espace de Banach dual,

35 (1980), 171–176

Talagrand, M., Dual Banach lattices and Banach lattices with the Radon–Nikodym

property, 38 (1981), 46–50

Talagrand, M., La propriété de Dunford–Pettis dans C(K, E) et L^{1}(E), 44 (1983),

317–321

Talagrand, M., La structure des espaces de Banach réticulés ayant la propriété de

Radon–Nikodym, 44 (1983), 213–220

Talagrand, M., Orlicz property and cotype in symmetric sequence spaces, 87 (1994),

181–192

Talagrand, M., Renormages de quelques C(K), 54 (1986), 327–334 Talagrand, M., Separabilité vague dans l’espace des mesures sur un compact, 37 (1980),

171–180

Talagrand, M., Sudakov-type minoration for Gaussian chaos processes, 79 (1992),

207–224

Talagrand, M., Sur la K-analyticité de certains espaces d’opérateurs, 32 (1979), 124–130 Talagrand, M., Sur les espaces de Banach contenant l^{1}(τ), 40 (1981), 324–330 Talagrand, M., Un nouveau C(K) qui possède la propriété de Grothendieck, 37 (1980),

181–191

Talagrand, M. see Godefroy, G. Talagrand, M. see Peck, N. T. Tamari, D. and Ginzburg, A., Representation of binary systems by families of binary

relations, 7 (1969), 21–32

Tamari, D. see Ginzburg, A. Tamburro, M. B., The evolution operator solution of the Cauchy problem for the

Hamilton–Jacobi equation, 26 (1977), 232–264

Tamir, B., On L-functions and intertwining operators for unitary groups, 73 (1991),

161–188

Tanny, D. see Madras, N. Tartar, L. see Crandall, M. G. Taskinen, J., Linearization of holomorphic mappings on C(K)-spaces, 92 (1995), 207–219 Tattersall, J. J., On the intersection of maximal m-convex subsets, 16 (1973), 300–305 Tauman, Y., A characterization of vector measure games in pNA, 43 (1982), 75–96 Taylor, A. D. see Kanamori, A. Taylor, P. D., A characterization of G-spaces, 10 (1971), 131–134 Taylor, P. D., The extension property for compact convex sets, 11 (1972), 159–163 Teicher, M., On toroidal embeddings of 3-folds, 57 (1987), 49–67 Temam, R. see Bensoussan, A. Temam, R. see Saut, J. C. Tench, R. N. see Moran, L. E. Teply, M. L., Modules semicocritical with respect to a torsion theory and their applications,

54 (1986), 181–200

Teply, M. L., Torsion-free covers II, 23 (1976), 132–136 Teply, M. L. see Golan, J. S. Teply, M. L. see Hutchinson, J. J. Terenzi, P. see Szankowski, A. Terreni, B. see Acquistapace, P.

Tervo, J., On semi-Fredholm properties of a boundary value problem in R_{+}^{n}, 63 (1988),

41–66

Tezer, C., The shift on the inverse limit of a covering projection, 59 (1987), 129–149 Thaler, M., A limit theorem for the Perron–Frobenius operator of transformations on

[0, 1] with indifferent fixed points, 91 (1995), 111–127

Thaler, M., Estimates of the invariant densities of endomorphisms with indifferent fixed

points, 37 (1980), 303–314

Thaler, M., Transformations on [0, 1] with infinite invariant measures, 46 (1983), 67–96 Thielcke, B. see Knörr , R. Thomas, A. see Blanchard, F. Thomas, J.-C. see Felix, Y. Thomas, R. M. see Campbell, C. M. Thomassen, C. see Reid, K. B. Thompson, R. C., Matrices with zero trace, 4 (1966), 33–42 Thouvenot, J.-P., On the stability of the weak Pinsker property, 27 (1977), 150–162 Thouvenot, J.-P., Quelques propriétés des systèmes dynamiques qui se décomposent en

un produit de deux systèmes dont l’un est un schéma de Bernoulli, 21 (1975), 177–207

Thouvenot, J.-P., Remarques sur les systèmes dynamiques donnés avec plusieurs facteurs,

21 (1975), 215–232

Thouvenot, J.-P., Une classe de systèmes pour lesquels la conjecture de Pinsker est vraie,

21 (1975), 208–214

Thouvenot, J.-P. see Shields, P. C. Thouvenot, J.-P. see Smorodinsky, M. Tietz, H., Über Tauber-Konstanten und eine Frage von Jakimovski, 8 (1970), 147–154 Tignol, J-P. see Haile, D. E. Tignol, J-P. see Rowen, L. H. Tignol, J-P. and Amitsur, S. A., Kummer subfields of Malcev–Neumann division

algebras, 50 (1985), 114–144

Tignol, J-P. and Amitsur, S. A., Symplectic modules, 54 (1986), 266–290 Tignol, J-P. see Amitsur, S. A. Timoney, R. M. see Dineen, S. Tintarev, K., On the geometry of P -convex sets for operators of real principal type, 64

(1988), 195–206

Tintarev, K. see Nishiwada, K. Todorcevic, S., Aronszajn trees and partitions, 52 (1985), 53–58 Tomczak-Jaegermann, N., The Banach–Mazur distance between symmetric spaces, 46

(1983), 40–66

Tomczak-Jaegermann, N., Uniform convexity of unitary ideals, 48 (1984), 249–254 Tomczak-Jaegermann, N. see Davis, W. J. Tomczak-Jaegermann, N. see Figiel, T. Tomczak-Jaegermann, N. see Garling, D. J. H. Tomczak-Jaegermann, N. see Gluskin, E. D. Tomczak-Jaegermann, N. see Komorowski, R. A. Tomczak-Jaegermann, N. see Mankiewicz, P. Tomkinson, M. J. see Menegazzo, F. Tong, A. E., On the existence of noncompact bounded linear operators between certain

Banach spaces, 10 (1971), 451–456

Torrea, J. L. see Garcia-Cuerva, J. Torrea, J. L. see Harboure, E. Torrecillas, B. see Bueso, J. L. Toruńczyk, H. see Simon, R. S. Tresser, C. see Galeeva, R. Trow, P., Decompositions of finite-to-one factor maps, 91 (1995), 129–155 Trow, P., Degrees of finite-to-one factor maps, 71 (1990), 229–238 Troyanski, S. L., On some generalizations of denting points, 88 (1994), 175–188 Troyanski, S. L. see Fabian, M. Tryba, J., A few remarks on Rowbottom cardinals, 40 (1981), 193–196 Tryba, J., On Jónsson cardinals with uncountable cofinality, 49 (1984), 315–324 Tsarpalias, A. see Negrepontis, S. Tucker, H. G. and Wright, A. L., An integral strong law of large numbers for processes

with independent increments, 38 (1981), 177–181

Tuncel, S., Conditional pressure and coding, 39 (1981), 101–112 Tuncel, S. see Cawley, E. Tuncel, S. see Kitchens, B. Turau, V., The 2-length of the Hughes subgroup, 62 (1988), 206–212 Turnidge, D. R. see Miller, R. W. Tuv, E. W. see Liskevich, V. A. Tylli, H.-O., The essential norm of an operator is not self-dual, 91 (1995), 93–110 Tzafriri, L., On Banach spaces with unconditional bases, 17 (1974), 84–93 Tzafriri, L., On multiplicity theory for Boolean algebras of projections, 4 (1966), 217–224 Tzafriri, L., On perturbation theory for spectral operators, 4 (1966), 62–64 Tzafriri, L., On the type and cotype of Banach spaces, 32 (1979), 32–38 Tzafriri, L., Remarks on contractive projections in L_{p}-spaces, 7 (1969), 9–15 Tzafriri, L., The connection between normalizable and spectral operators, 3 (1965), 75–80 Tzafriri, L. see Bourgain, J. Tzafriri, L. see Casazza, P. G. Tzafriri, L. see Gluskin, E. D. Tzafriri, L. see Johnson, W. B. Tzafriri, L. see Lindenstrauss, J. Tzimbalario, J. see Bak, J. Tzkoni, I. see Furstenberg, H. Ulbrich, K.-H., On Hopf algebras and rigid monoidal categories, 72 (1990), 252–256 Ullrich, D. C., An extension of the Kahane–Khinchine inequality in a Banach space, 62

(1988), 56–62

Upfal, E. see Shamir, E. Urbański, M. see Denker, M. Uttersrud, U., Geometrical properties of subclasses of complex L_{1}-preduals, 72 (1990),

353–371

Uttersrud, U. see Lima, A. Vaillancourt, R., Bounded pseudo-differential operators, 13 (1972), 225–231 Vaisman, I., Exotic characteristic classes of quaternionic bundles, 69 (1990), 46–58 Vaisman, I., On locally conformal almost Kähler manifolds, 24 (1976), 338–351 Valdivia, M., Banach spaces X with X^{∗∗} separable, 59 (1987), 107–111 Valdivia, M., Shrinking bases and Banach spaces Z^{∗∗}/Z, 62 (1988), 347–354 Valdivia, M., Topological direct sum decompositions of Banach spaces, 71 (1990), 289–296 Valenti, A. see Giambruno, A. Valentine, F. A., Local convexity and starshaped sets, 3 (1965), 39–42 Valette, A. see de la Harpe, P. de Valk, V., The maximal and minimal 2-correlation of a class of 1-dependent 0-1 valued

processes, 62 (1988), 181–205

Vamanamurthy, M. K. see Anderson, G. D. Van den Bergh, M., Explicit rational forms for the Poincaré series of the trace rings of

generic matrices, 73 (1991), 17–31

Van den Bergh, M., Group rings over Dedekind rings, 61 (1988), 295–300 Van den Bergh, M. and Van Geel, J., Algebraic elements in division algebras over

function fields of curves, 52 (1985), 33–45

Van Geel, J. see Van den Bergh, M.

Van Maldeghem, H. and Weiss, R., On finite Moufang polygons, 79 (1992), 321–330 Van Oystaeyen, F. and Verschoren, A., On the Brauer group of a projective curve,

42 (1982), 327–333

Van Oystaeyen, F. and Verschoren, A., On the Brauer group of a projective variety,

42 (1982), 37–59

Van Oystaeyen, F. and Zhang, Y. H., The crossed coproduct theorem and Galois

cohomology, 96 (1996), 579–607

Van Oystaeyen, F. see Nauwelaerts, E. Vanhecke, L. see Berndt, J. Vanstone, S. A., Stinson, D. R., Schellenberg, P. J., Rosa, A., Rees, R.,

Colbourn, C. J., Carter, M. W. and Carter, J. E., Hanani triple systems, 83 (1993), 305–319

Vapne, A. see Leron, U. Varela, F. see Gonzalo, J. Varma, A. K., Concerning the order of approximation of periodic continuous functions by

trigonometric interpolation polynomials, 12 (1972), 337–341

Varma, A. K., Simultaneous approximation of periodic continuous functions and their

derivatives, 6 (1968), 66–73

Varma, A. K., Some remarks on trigonometric interpolation, 7 (1969), 177–185 Varma, A. K. see Mills, T. M. Vaserstein, L. see Grunewald, F. Vázquez, J. L., Monotone perturbations of the Laplacian in L^{1}(R^{N}), 43 (1982), 255–272 Veech, W. A., Finite group extensions of irrational rotations, 21 (1975), 240–259 Veech, W. A., Minimality of horospherical flows, 21 (1975), 233–239 Veech, W. A., The tail of a positivity-preserving semigroup, 18 (1974), 167–206 Velázquez, J. J. L. see Herrero, M. A. Vella, A. D. see Hill, P. Venkataramana, T. N., On some rigid subgroups of semisimple Lie groups, 89 (1995),

227–236

Vera López, A., Arithmetical conditions on the conjugacy vector of a finite group, 56

(1986), 179–187

Vera López, A., Conjugacy classes in finite solvable groups, 47 (1984), 241–245 Vera López, A. and Fernández-Alcober, G. A., The conjugacy vector of a p-group of

maximal class, 86 (1994), 233–252

Vera López, A. and Larrea, M^{a} C., On the number of conjugacy classes in a finite

group II, 64 (1988), 87–127

Vera López, A. and Vera López, J., Classification of finite groups according to the

number of conjugacy classes, 51 (1985), 305–338

Vera López, A. and Vera López, J., Classification of finite groups according to the

number of conjugacy classes II, 56 (1986), 188–221

Vera López, J. see Vera López, A. Verschoren, A., The affine scheme associated to a nonnoetherian prime PI algebra, 53

(1986), 121–128

Verschoren, A. see Van Oystaeyen, F. Verstraelen, L. see Dillen, F. Vespri, V. see Cannarsa, P. Vesterstrom, J., Positive linear extension operators for spaces of affine functions, 16

(1973), 203–211

Vidossich, G., A theorem on uniformly continuous extensions of mappings defined in

finite-dimensional spaces, 7 (1969), 207–210

Vidossich, G., On a theorem of Corson and Lindenstrauss on Lindelöf function spaces, 9

(1971), 271–278

Vignati, M. see Rigoli, M. Villamayor, O. E. see Evyatar, A. A.

Vinner, S., A generalization of Ehrenfeucht’s game and some applications, 12 (1972),

279–298

Vitale, R. A., The Steiner point in infinite dimensions, 52 (1985), 245–250 Völklein, H., Braid group action via GL_{n}(q) and U_{n}(q), and Galois realizations, 82

(1993), 405–427

Völklein, H., Moduli spaces for covers of the Riemann sphere, 85 (1994), 407–430 Völklein, H. see Fried, M. D. Völklein, H. see Haran, D. Volný, D. see Liardet, P. Vonessen, N. see Reichstein, Z. Vrabie, I. I., The nonlinear version of Pazy’s local existence theorem, 32 (1979), 221–235 Vrećica, S. T. and Živaljević, R. T., The ham sandwich theorem revisited, 78 (1992),

21–32

Vuorinen, M. see Anderson, G. D. Wadsworth, A. R. see Jacob, B. Wadsworth, A. R. see Morandi, P. J. Wadsworth, A. R. see Small, L. W. Wajnryb, B., A braidlike presentation of Sp(n, p), 76 (1991), 265–288 Wajnryb, B., A simple presentation for the mapping class group of an orientable surface,

45 (1983), 157–174

Wajnryb, B., Finiteness conditions in Krull subrings of a ring of polynomials, 43 (1982),

169–176

Wajnryb, B. and Birman, J. S., Errata: Presentations of the mapping class group, 88

(1994), 425–427

Wajnryb, B. see Birman, J. S. Waksman, Z. and Wasilewsky, J., A theorem on level lines of continuous functions, 27

(1977), 247–251

Wales, D. B. see Cohen, A. M. Wall, C. E. see Chartrand, G. Wallis, J. see Wallis, W. D. Wallis, W. D., Integral equivalence of Hadamard matrices, 10 (1971), 349–358 Wallis, W. D., Some notes on integral equivalence of combinatorial matrices, 10 (1971),

457–464

Wallis, W. D. and Wallis, J., Equivalence of Hadamard matrices, 7 (1969), 122–128 Walls, G. L., Groups with maximal subgroups of Sylow subgroups normal, 43 (1982),

166–168

Walther, H., Longest cycles in polyhedral graphs, 83 (1993), 203–212 Wamsley, J. W., Minimal presentations for certain group extensions, 9 (1971), 459–463 Wamsley, J. W., On a class of groups of prime-power order, 11 (1972), 297–298 Wamsley, J. W. see Johnson, D. L. Wang, A. T. and Chen, C. S., Functions of an n-dimensional Brownian motion that are

Markovian, 34 (1979), 343–352

Wang, E. T., On permanents of (1, −1)-matrices, 18 (1974), 353–361 Wang, L., Brauer–Manin obstruction to weak approximation on abelian varieties, 94

(1996), 189–200

Wang, L.-Y. see Liao, G.-F. Wang, S. see deLaubenfels, R. Wang, Y. see deLaubenfels, R. Ward, T. B., Almost block independence for the three dot ℤ^{2} dynamical system, 76

(1991), 237–256

Ward, T. B., The Bernoulli property for expansive ℤ^{2} actions on compact groups, 79

(1992), 225–249

Ward, T. see Chothi, V. Warren, P. see Beck, A.

Wasilewsky, J. see Waksman, Z. Watanabe, T., The local theta correspondence for unramified unitary groups, 92 (1995),

397–410

Watson, P. see Cooper, S. B. Watson, W. S. see Steprāns, J. Webb, G. F., Nonlinear perturbations of linear accretive operators in Banach spaces, 12

(1972), 237–248

Wegner, G., Note on a paper of B. Grünbaum on acyclic colorings, 14 (1973), 409–412 Wegner, G., Über eine kombinatorisch-geometrische Frage von Hadwiger und Debrunner,

3 (1965), 187–198

Wegner, G. see Barnette, D. W. Wehrfritz, B. A. F., On division rings generated by polycyclic groups, 47 (1984), 154–164 Wehrfritz, B. A. F., Subgroups of prescribed finite index in linear groups, 58 (1987),

125–128

Weichsel, P. M., A note on asymmetric graphs, 10 (1971), 234–243 Weichsel, P. M., Just-irregular p-groups, 10 (1971), 359–363 Weigel, T. S., Residual properties of free groups, III, 77 (1992), 65–81 Weil, W., Centrally symmetric convex bodies and distributions, 24 (1976), 352–367 Weil, W., Centrally symmetric convex bodies and distributions II, 32 (1979), 173–182 Weinberger, S., On fibering four- and five-manifolds, 59 (1987), 1–7 Weinberger, S., Semifree locally linear PL actions on the sphere, 66 (1989), 351–363 Weinberger, S. see Manevitz, L. M. Weiss, B., A combinatorial proof of the Borsuk–Ulam antipodal point theorem, 66 (1989),

364–368

Weiss, B. see Aaronson, J. Weiss, B. see Adler, R. L. Weiss, B. see Arnoux, P. Weiss, B. see Connes, A. Weiss, B. see Furstenberg, H. Weiss, B. see Glasner, S. Weiss, B. see Kalikow, S. Weiss, B. see Kamae, T. Weiss, B. see Katznelson, Y. Weiss, B. see Ornstein, D. S. Weiss, B. see Shelah, S. Weiss, B. see Smorodinsky, M. Weiss, E.-A., Jr. see Gilat, D. Weiss, G., Admissible observation operators for linear semigroups, 65 (1989), 17–43 Weiss, G. H. see Kiefer, J. E. Weiss, R. see Van Maldeghem, H. Weiss, W. see Dow, A. Weissler, F. B., Construction of non-linear semi-groups using product formulas, 29

(1978), 265–275

Weissler, F. B., Existence and non-existence of global solutions for a semilinear heat

equation, 38 (1981), 29–40

Weit, Y., On the one-sided Wiener’s theorem for the motion group on R^{N}, 55 (1986),

111–120

Weit, Y., Translation-invariant subspaces of functions on the group of linear transformations

on the real line, 37 (1980), 256–272

Wenzel, J., Vector-valued Walsh–Paley martingales and geometry of Banach spaces, 97

(1997), 29–49

Werner, E. see Schachermayer, W. Wesley, E., Extensions of the measurable choice theorem by means of forcing, 14 (1973),

104–114 Weston, A., Some non-uniformly homeomorphic spaces, 83 (1993), 375–380 Westreich, D., Periodic solutions of second order non-Lagrangian systems, 16 (1973),

279–286

Westreich, S. see Cohen, M. Wetzel, J. E. see Schaer, J. Wheeler, W. H., Model-companions and definability in existentially complete structures,

25 (1976), 305–330

Whitfield, J. H. M. see Fabian, M. Wierdl, M., Pointwise ergodic theorem along the prime numbers, 64 (1988), 315–336 Williams, F. see Lin, J. P. Williams, K. S. see Hudson, R. H. Williams, S. G. see Silver, D. S. Willis, G. A. see Grønbæk, N. Wills, J. M. see Kleinschmidt, P. Wills, J. M. see McMullen, P. Wimmers, E. L., The Shelah P -point independence theorem, 43 (1982), 28–48 Winkler, P. M. see Duke, R. A. Witsenhausen, H. S. see Gordon, Y. Wittmann, R., Analogues of the “zero-two” law for positive linear contractions in L^{p} and

C(X), 59 (1987), 8–28

Wittmann, R. see Krengel, U. Wittmann, R. see Lin, M. Witzany, J. see Gitik, M. Woess, W., Boundaries of random walks on graphs and groups with infinitely many ends,

68 (1989), 271–301

Wojtaszczyk, P., Uniqueness of unconditional bases in quasi-Banach spaces with appli-

cations to Hardy spaces, II, 97 (1997), 253–280

Wojtaszczyk, P. and Woźniakowski, K., Orthonormal polynomial bases in function

spaces, 75 (1991), 167–191

Wojtaszczyk, P. see Kalton, N. J. Wolf, J. A. see Keene, F. W. Wolfe, J., Injective Banach spaces of type C(T), 18 (1974), 133–140 Wolff, M. P. H., Skew-product extensions of Markov operators and products of dependent

random variables, 34 (1979), 161–175

Wolff, M. P. H., Spectral theory of group representations and their nonstandard hull,

48 (1984), 205–224

Wolfson, H. see Milman, V. D. Woo, J. Y. T., On a class of universal modular sequence spaces, 20 (1975), 193–215 Wood, C., The model theory of differential fields revisited, 25 (1976), 331–352 Wood, G. R., Divisible points of compact convex sets, 54 (1986), 351–365 Woodin, H. see Shelah, S. Woods, R. G., Some ℵ_{0}-bounded subsets of Stone-Čech compactifications, 9 (1971),

250–256

Woon, E. Y., n-Connectedness in pure 2-complexes, 52 (1985), 177–192 Woźniakowski, K. see Wojtaszczyk, P. Wright, A. L. see Tucker, H. G. Wschebor, M., Sur le recouvrement du cercle par des ensembles placés au hasard, 15

(1973), 1–11

Wu, Q., Gelfand–Kirillov dimension under base field extension, 73 (1991), 289–296 Xiao, P. A. see Zhou, J. Xu, K., Nilpotent subgroups of self equivalences of torsion spaces, 93 (1996), 189–194 Xu, Q., Some results related to interpolation on Hardy spaces of regular martingales, 91

(1995), 173–187

Yahav, J. A. see Bickel, P. J.

Yakir, A. see Frumkin, A. Yamashita, S., An extremal problem for Dirichlet-finite holomorphic functions on Riemann

surfaces, 50 (1985), 201–206

Yamashita, S., Functions with slowly-growing area and harmonic majorants, 39 (1981),

259–264

Yang, S., Estimates for coefficients of univalent functions from integral means and Grunsky

inequalities, 87 (1994), 129–142

Yanushka, A., Generalized hexagons of order (t, t), 23 (1976), 309–324 Yekutieli, A. see Sastry, P. Yomdin, Y., C^{k}-resolution of semialgebraic mappings. Addendum to: “Volume growth

and entropy”, 57 (1987), 301–317

Yomdin, Y., Some estimates for a decomposition of hypersurface singularities into the

simplest ones, 36 (1980), 126–132

Yomdin, Y., Volume growth and entropy, 57 (1987), 285–300 Yomdin, Y. see Briskin, M. Yorozu, S., A_{ν}-operator on complete foliated Riemannian manifolds, 56 (1986), 349–354 Yoshida, H., On subharmonic functions dominated by certain functions, 54 (1986), 366–380 Yoshino, T., A counter-example to a conjecture of Friedland, 75 (1991), 273–275 Yue, C. B., Foliations with leaves of nonpositive curvature, 97 (1997), 113–123 Yue, C. B. see Kononenko, A. Zabrodsky, A., A note on cohomology with limited torsion, 8 (1970), 159–164 Zabrodsky, A., Correction to: “On the homotopy type of principal classical group bundles

over spheres”, 12 (1972), 331–331

Zabrodsky, A., On phantom maps and a theorem of H. Miller, 58 (1987), 129–143 Zabrodsky, A., On spaces of functions between classifying spaces, 76 (1991), 1–26 Zabrodsky, A., On the homotopy type of principal classical group bundles over spheres,

11 (1972), 315–325

Zabrodsky, A., Some relations in the mod 3 cohomology of H-spaces, 33 (1979), 59–72 Zachariades, T. see Argyros, S. Zafran, M., On the symbolic calculus in homogeneous Banach algebras, 32 (1979), 183–192 Zaharopol, R., Quelques observations sur les L^{2}-resolvantes, 33 (1979), 52–58 Zakai, M., Some moment inequalities for stochastic integrals and for solutions of stochastic

differential equations, 5 (1967), 170–176

Zakrzewski, P., The existence of invariant σ-finite measures for a group of transformations,

83 (1993), 275–287

Zakrzewski, P., The existence of invariant probability measures for a group of transfor-

mations, 83 (1993), 343–352

Zaks, A., Dedekind subrings of k[x_{1}, …, x_{n}] are rings of polynomials, 9 (1971), 285–289 Zaks, A., Half-factorial-domains, 37 (1980), 281–302 Zaks, A., Restricted left principal ideal rings, 11 (1972), 190–215 Zaks, A., Semi-primary hereditary rings, 6 (1968), 359–362 Zaks, A., Some rings are hereditary rings, 10 (1971), 442–450 Zaks, J., 6-valent analogues of Eberhard’s theorem, 18 (1974), 19–29 Zaks, J., Extending Kotzig’s theorem, 45 (1983), 281–296 Zaks, J., On realizing symmetric 3-polytopes, 10 (1971), 244–251 Zaks, J., The analogue of Eberhard’s theorem for 4-valent graphs on the torus, 9 (1971),

299–305

Zalcman, L., Mean values and differential equations, 14 (1973), 339–352 Zalcman, L. see Aharonov, D. Zalesskiĭ, A. E., Modular group rings of the finitary symmetric group, 96 (1996), 609–621 Zalesskiĭ, A. E. see Hartley, B. Zamfiresco, T., Les courbes fermées doubles sans points triples associées a une famille

continue, 7 (1969), 69–89

Zamir, S. see Mertens, J.-F.

Zanco, C. see Klee, V. Zapletal, J., Strongly almost disjoint functions, 97 (1997), 101–111 Zarantonello, E. H., Dense single-valuedness of monotone operators, 15 (1973), 158–166 Zedek, M., Cayley’s decomposition and Polya’s W -property of ordinary linear differential

equations, 3 (1965), 81–86

Zeitz, P., The centralizer of a rank-one flow, 84 (1993), 129–145 Zelinsky, D., Berkson’s theorem, 2 (1964), 205–209 Zelmanov, E. I., On periodic compact groups, 77 (1992), 83–95 Zelmanov, E. I. see Martinez, C. Zeman, M., The Levi–Lax condition for partial differential equations with real characteristics

of constant multiplicity, 31 (1978), 57–77

Zeman, M. see Grimmer, R. Zhang, Q., An example of compact extensions of Kronecker factors and skew products of

irrational rotations with finite groups, 92 (1995), 249–261

Zhang, Y. H. see Van Oystaeyen, F. Zhitomirskii, M. see Mormul, P. Zhou, J. and Xiao, P. A., A note on the valence of certain means, 83 (1993), 289–294 Zhou, S. P., A counterexample in copositive approximation, 78 (1992), 75–83 Zhu, C.-B. see Lee, S. T. Zhu, S. see Caenepeel, S. Zhu, S. see Cohen, M. Ziegler, G. M., Shellability of chessboard complexes, 87 (1994), 97–110 Ziegler, Z. see Amir, D. Ziegler, Z. see Karlin, S. Zimmer, R. J., Cocycles and the structure of ergodic group actions, 26 (1977), 214–220 Zimmer, R. J., Groups generating transversals to semisimple Lie group actions, 73 (1991),

151–159

Zimmer, R. J., On the cohomology of ergodic group actions, 35 (1980), 289–300 Zimmer, R. J., On the Mostow rigidity theorem and measurable foliations by hyperbolic

space, 43 (1982), 281–290

Zimmer, R. J., Random walks on compact groups and the existence of cocycles, 26 (1977),

84–90

Zimmer, R. J., Spectrum, entropy, and geometric structures for smooth actions of Kazhdan

groups, 75 (1991), 65–80

Zimmer, R. J., Superrigidity, Ratner’s theorem, and fundamental groups, 74 (1991),

199–207

Zimmer, R. J. see Lubotzky, A. Zimmer, R. J. see Pansu, P. Zimmer, R. J. see Qian, N. Zippin, M., A remark on bases and reflexivity in Banach spaces, 6 (1968), 74–79 Zippin, M., Correction to: “The finite-dimensional P_{λ} spaces with small λ”, 48 (1984),

255–256

Zippin, M., L_{1} subspaces of l_{1}, 22 (1975), 110–117 Zippin, M., On a certain basis in c_{0}, 4 (1966), 199–204 Zippin, M., On perfectly homogeneous bases in Banach spaces, 4 (1966), 265–272 Zippin, M., The finite-dimensional P_{λ} spaces with small λ, 39 (1981), 359–364 Zippin, M., The range of a projection of small norm in l_{1}^{n}, 39 (1981), 349–358 Zippin, M., The separable extension problem, 26 (1977), 372–387 Zippin, M. see Johnson, W. B. Zippin, M. see Lazar, A. J. Zippin, M. see Szankowski, A. Zisser, I., Irreducible products of characters in A_{n}, 84 (1993), 147–151 Zisser, I., The covering numbers of the sporadic simple groups, 67 (1989), 217–224 Živaljević, R. T. see Vrećica, S. T.

Zizler, V. see Deville, R. Zizler, V. see Fabian, M. Zizler, V. see John, K. Zlobec, S. and Ben-Israel, A., On explicit solutions of interval linear programs, 8

(1970), 12–22

Zuazua, E. see Escobedo, M. Zwonek, W., Carathéodory balls and norm balls of the domains H_{n} = {z ∈ ℂ^{n}:

|z_{1}| + ⋯ + |z_{n}| < 1}, 89 (1995), 71–76